### v3

Abstract
We propose a nonparametric, kernel-based joint estimator for conditional mean and covariance matrices in large unbalanced panels. Our estimator, with proven consistency and finite-sample guarantees, is applied to a comprehensive panel of monthly US stock excess returns from 1962 to 2021, conditioned on macroeconomic and firm-specific covariates. The estimator captures time-varying cross-sectional dependencies effectively, demonstrating robust statistical performance. In asset pricing, it generates conditional mean-variance efficient portfolios with out-of-sample Sharpe ratios that substantially exceed those of equal-weighted benchmarks.
Keywords: nonparametric estimation, conditional mean, conditional covariance matrix, unbalanced panels, panel data econometrics, mean-variance efficient portfolio
JEL classification: C14, C58, G11
1 Introduction
The relationship between conditional expected returns, conditional risk, and asset characteristics has been a central topic in financial economics for decades. Yet, inference in this domain remains constrained by the unbalanced and high-dimensional nature of real-world data. In this paper, we address these challenges by introducing a nonparametric, kernel-based framework for the joint estimation of conditional mean and covariance matrices, providing a powerful and tractable solution to the econometric inference problem highlighted by Cochrane (2011). Our framework is specifically designed to deliver positive semidefinite second moment and covariance matrices across any state and for cross-sections of varying sizes, filling a significant gap in the literature.
Since Fama and MacBeth (1973), empirical researchers studying unbalanced panels have primarily relied on tools such as portfolio sorts (Fama and French, 1993, 2019; Kozak et al., 2020), models for expected returns applicable to both balanced and unbalanced panels (Connor et al., 2012; Fan et al., 2016; Freyberger et al., 2020; Gu et al., 2020b; Kelly et al., 2019; Kozak and Nagel, 2024), and econometric inference methods for linear factor models (Zaffaroni, 2019; Fortin et al., 2023a, b). While recent econometric literature has introduced conditional covariance estimators tailored for high-dimensional settings (Engle et al., 2019), there remains a critical gap: a scalable, nonparametric framework capable of jointly and consistently estimating conditional means and covariances in a large-scale, unbalanced context.
We address this gap by proposing a novel, nonparametric, kernel-based model for jointly estimating conditional first and second moments for unbalanced panels of arbitrary size, requiring only that these conditional moments can be represented within a large and flexible hypothesis space. Our model uniquely ensures that, at any point in time and across any cross-sectional dimension, conditional return covariances incorporate both systematic and idiosyncratic components, and remain symmetric and positive semidefinite, despite their nonparametric construction. We refer to this model as the joint conditional mean and covariance (COCO) estimator. Our approach tackles a significant limitation in the literature, which typically focuses on either covariances or first moments independently. Moreover, our model’s functional form is optimal with respect to the mean squared loss and, while it is broadly applicable, aligns precisely with the characterization of economies that can be spanned by factor portfolios, as discussed in Kozak and Nagel (2024).
The COCO estimator is computationally efficient and scalable, capable of handling large datasets on standard desktop computing hardware. Its parsimonious structure means that finite-dimensional specifications may not even require validation, enhancing practical applicability. Furthermore, the model provides a natural low-rank representation with controlled approximation error, leading to a Chamberlain and Rothschild (1983)-type conditional factor structure, where rank corresponds to the number of systematic factors. Crucially, the estimator emerges from a convex optimization problem, ensuring reproducibility—a distinct advantage over non-convex models prevalent in deep learning and other econometric frameworks.
We empirically test the COCO estimator on an extensive unbalanced panel of monthly US stock returns from 1962 to 2021, incorporating both firm-specific characteristics and macroeconomic covariates. Our results indicate that the COCO estimator offers moderate predictability for realized excess returns, with stronger and more reliable predictability for their squares and mixed products, which correspond to conditional second moments, especially in the early sample period. When jointly assessing both moments, the COCO estimator significantly outperforms a baseline model that accounts only for idiosyncratic risk. The conditional mean-variance efficient (cMVE) portfolio constructed from the COCO estimates achieves substantial annualized out-of-sample Sharpe ratios, markedly outperforming equal-weighted portfolios over the entire sample period. Furthermore, cMVE returns exhibit weak correlations with the Fama–French five factors (Fama and French, 2015). As the number of systematic factors in our model increases, the connection to the Fama–French factors diminishes, ultimately rendering the variation in cMVE portfolio returns largely unrelated to the traditional five-factor model.
The remainder of this paper is organized as follows. Section 2 introduces the nonparametric model for conditional moments and connects it to generative linear factor models. Section 3 defines the joint estimator, deriving a representation theorem for the optimal conditional moment function and a corresponding low-rank approximation. Section 4 establishes consistency and provides finite-sample guarantees. Section 5 presents a large-scale application of the model on a panel of US stock returns. Section 6 concludes. The appendix contains all proofs.
2 Conditional mean and covariance model
We first introduce the basic econometric framework established in this paper, along with the necessary notation. We consider discrete time periods $(t,t+1]$, $t=0,1,2,\dots$, e.g., months. For each period $(t,t+1]$, there are $N_{t}$ assets $i=1,\dots,N_{t}$ with observable covariates $z_{t,i}$ at $t$, taking values in some covariate space ${\mathcal{Z}}$, which yield returns $x_{t+1,i}$ over $(t,t+1]$. We remain agnostic about the type of “return” that could be gross, simple, logarithmic, excess, or forward gross. In the empirical study, we will work with simple excess returns, as is customary in the literature and convenient for asset pricing. Following the introduction, we denote by $\bm{x}_{t+1}\coloneqq[x_{t+1,i}:1\leq i\leq N_{t}]\in{\mathbb{R}}^{N_{t}}$ and $\bm{z}_{t}\coloneqq[z_{t,i}:1\leq i\leq N_{t}]\in{\mathcal{Z}}^{N_{t}}$ the corresponding return and covariate arrays.
Our goal is to define a model for conditional first and second moments, ${\mathbb{E}}_{t}[x_{t+1,i}]$ and ${\mathbb{E}}_{t}[x_{t+1,i}x_{t+1,j}]$, of the returns, given the information set at time $t$. To this end, we assume that these conditional moments are given by functions $\mu:{\mathcal{Z}}\to{\mathbb{R}}$ and $q:{\mathcal{Z}}\times{\mathcal{Z}}\to{\mathbb{R}}$ of the respective covariates such that
${\mathbb{E}}_{t}[x_{t+1,i}]$ $=\mu(z_{t,i}),$
${\mathbb{E}}_{t}[x_{t+1,i}x_{t+1,j}]$ $=q(z_{t,i},z_{t,j}),$
which implies the conditional covariance $\operatorname{Cov}_{t}[x_{t+1,i},x_{t+1,j}]=q(z_{t,i},z_{t,j})-\mu(z_{t,i})\mu(z_{t,j})$. We denote by $\mu(\bm{z}_{t})\coloneqq[\mu(z_{t,i}):1\leq i\leq N_{t}]$ and $q(\bm{z}_{t},\bm{z}_{t}^{\top})\coloneqq[q(z_{t,i},z_{t,j}):1\leq i,j\leq N_{t}]$ the respective arrays of values. The challenge is to find functions $\mu$ and $q$ such that, for any $t$, the $N_{t}\times N_{t}$-conditional second moment, and covariance, matrix
  1. (i)
    $q(\bm{z}_{t},\bm{z}_{t}^{\top})$ is symmetric and positive semidefinite, and
  2. (ii)
    $q(\bm{z}_{t},\bm{z}_{t}^{\top})-\mu(\bm{z}_{t})\mu(\bm{z}_{t})^{\top}$ is symmetric and positive semidefinite,
respectively. Conditions (i) and (ii) ensure that all conditional moments are mutually consistent, as if a conditional probability measure had generated these moments. Property (i) coincides with the defining property of a real-valued kernel function, see, e.g., Paulsen and Raghupathi (2016, Section 2.2). We thus assume that $q$ is a kernel function on ${\mathcal{Z}}\times{\mathcal{Z}}$. To assert property (ii), we extend the covariate space ${\mathcal{Z}}_{\Delta}\coloneqq{\mathcal{Z}}\cup\{\Delta\}$, for some auxiliary point ${\Delta}\notin{\mathcal{Z}}$. We then extend $q$ to be a kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ such that
$q({\Delta},{\Delta})=1,$ (1)
and set $\mu(z)\coloneqq q(z,{\Delta})$. This implies that the implied covariance function $c(z,z^{\prime})\coloneq q(z,z^{\prime})-\mu(z)\mu(z^{\prime})=q(z,z^{\prime})-q(z,{\Delta})q(z^{\prime},{\Delta})$ is the Schur complement of $q$ with respect to ${\Delta}$. It is therefore itself a kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ (see Paulsen and Raghupathi, 2016, Theorem 4.5), and thus property (ii) holds.
It remains to specify an appropriate kernel function $q$ such that (1) holds. Finding such extensions to ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ of a given kernel function $q$ on ${\mathcal{Z}}\times{\mathcal{Z}}$, other than trivially setting $q({\Delta},z)=q(z,{\Delta})\coloneq 1_{z=\Delta}$, is rather difficult in general. Instead, we propose here a novel nonparametric approach to directly learning a kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ that reflects principles from financial econometrics and also suits the data.
Specifically, we follow the standard assumption in asset pricing that the conditional covariance can be decomposed into a systematic and an idiosyncratic component. The former captures the conditional dependence between returns explained by common underlying risk factors. The latter captures the conditional uncorrelated individual return risks, which asymptotically have a conditional mean of zero under the absence of arbitrage in large cross sections (see Ross, 1976; Chamberlain, 1983; Chamberlain and Rothschild, 1983; Reisman, 1988). We take this into account and decompose $q(z,z^{\prime})=q^{\rm sy}(z,z^{\prime})+q^{\rm id}(z,z^{\prime})$ into the sum of two corresponding kernel functions, where the idiosyncratic component $q^{\rm id}(z,z^{\prime})=q^{\rm id}(z,z^{\prime})1_{z=z^{\prime}}$ is supported on the diagonal of the product space ${\mathcal{Z}}\times{\mathcal{Z}}$. Accordingly, we set $q^{\rm sy}(\Delta,\Delta)=1$ and $q^{\rm id}(\Delta,\Delta)=0$ so that the systematic component captures the structural condition (1).
Our formal framework relies on an auxiliary separable Hilbert space ${\mathcal{C}}$, assumed to be sufficiently large to accommodate the setup described. [^4] See Lemma 3.3. We fix an arbitrary unit vector $p\in{\mathcal{C}}$ such that $\langle p,p\rangle_{\mathcal{C}}=1$. For any pair of feature maps $h=(h^{{\rm sy}},h^{{\rm id}})$ with $h^{\tau}:{\mathcal{Z}}\to{\mathcal{C}}$, we extend them to ${\mathcal{Z}}_{\Delta}$ by setting their values at $\Delta$ to be the zero element in ${\mathcal{C}}$:
$h^{\tau}(\Delta)\coloneqq 0,\quad\text{for }\tau\in\{{\rm sy},{\rm id}\}.$ (2)
This allows us to define a moment kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ as follows: [^5] We use that inner products are kernel functions, and so are sums, products and pull-backs of kernel functions, see Paulsen and Raghupathi (2016, Section 2.3.4 and Chapter 5).
$q_{h}(z,z^{\prime})\coloneqq\underbrace{\langle h^{\rm sy}(z)+p1_{z=\Delta},h^{\rm sy}(z^{\prime})+p1_{z^{\prime}=\Delta}\rangle_{\mathcal{C}}}_{\text{systematic component }q^{{\rm sy}}_{h}(z,z^{\prime})}+\underbrace{\|h^{\rm id}(z)\|_{\mathcal{C}}^{2}\,1_{z=z^{\prime}}}_{\text{idiosyncratic component }q^{{\rm id}}_{h}(z,z^{\prime})}.$ (3)
From (2), it follows that (1) holds. This implies the conditional mean and covariance functions
$\mu_{h}(z)$ $=\langle{h^{\rm sy}}(z),p\rangle_{\mathcal{C}},$ (4)
[…]
We henceforth assume that $z_{t,i}=z_{t,j}$ if and only if $i=j$, for each cross section $t$. This is without loss of generality, as otherwise we could simply assume that the index $i$ is part of the characteristics $z_{t,i}$. In turn, we obtain a diagonal idiosyncratic matrix component in the expressions below.
Our framework (3) for the moment kernel function is universal and covers any generative conditional factor model of the form
$x_{t+1,i}=\underbrace{\alpha(z_{t,i})}_{\text{intercept}}+\underbrace{\langle\beta(z_{t,i}),g_{t+1}\rangle_{\mathcal{C}}}_{\text{systematic risk}}+\underbrace{\gamma(z_{t,i})\xi_{t+1}(z_{t,i})}_{\text{idiosyncratic risk}},$ (5)
for some conditional intercept function $\alpha:{\mathcal{Z}}\to{\mathbb{R}}$, factor loadings map $\beta:{\mathcal{Z}}\to{\mathcal{C}}$ and idiosyncratic volatility function $\gamma:{\mathcal{Z}}\to[0,\infty)$. Further, $g_{t+1}$ is a ${\mathcal{C}}$-valued stationary risk factor process, with constant conditional mean $b\coloneqq{\mathbb{E}}_{t}[g_{t+1}]$ and covariance operator $Q\coloneqq\operatorname{Cov}_{t}[g_{t+1}]$. In addition, $\{\xi_{t+1}(z):z\in{\mathcal{Z}}\}$ is a white noise process, such that ${\mathbb{E}}_{t}[\xi_{t+1}(z)]=0$ and ${\mathbb{E}}_{t}[\xi_{t+1}(z)\xi_{t+1}(z^{\prime})]=1_{z=z^{\prime}}$, which is conditionally uncorrelated with $g_{t+1}$. We also assume that ${\mathcal{C}}$ is large enough such that $g_{t+1}$ and $\beta(z)$ take values in a subspace of ${\mathcal{C}}$ of codimension 1.666This is in line with Lemma 3.3 (i) and without loss of generality, as otherwise we simply extend ${\mathcal{C}}$ to ${\mathbb{R}}\oplus{\mathcal{C}}$. Note that the representation (5) of $x_{t+1,i}$ is not unique. For instance, we can demean the factors, replacing $g_{t+1}$ by $g_{t+1}-b$ and $\alpha(z)$ by $\alpha(z)+\langle\beta(z),b\rangle_{\mathcal{C}}$. Further, we can incorporate the intercept in the systematic component, replacing $\beta(z)$ by $\alpha(z)q+\beta(z)$ and $g_{t+1}$ by $g_{t+1}+q$, for some unit vector $q\in{\mathcal{C}}$ that is orthogonal to $g_{t+1}$ and $\beta(z)$, which by assumption exists. Moreover, we can rotate the factors, replacing $g_{t+1}$ by $Ag_{t+1}$ and $\beta(z)$ by $B\beta(z)$ for any linear operators $A,B$ on ${\mathcal{C}}$ such that $B^{\ast}A=I_{\mathcal{C}}$.
The following theorem formalizes our claim. The third part gives a representation of the generative conditional factor model (5) in terms of factors that are linear in $\bm{x}_{t+1}$ and therefore observable, in contrast to $g_{t+1}$, which may be latent. These observable factors can be interpreted as portfolio returns, with significant implications for asset pricing, as examined in detail in Filipović and Schneider (2024). For any linear operator $A:{\mathcal{C}}\to{\mathbb{R}}^{N_{t}}$, we denote by $A^{+}$ its pseudoinverse, which is pointwise defined in terms of its adjoint $A^{\ast}$ by $A^{+}\bm{v}\coloneqq\lim_{\lambda\downarrow 0}(A^{\ast}A+\lambda I_{\mathcal{C}})^{-1}A^{\ast}\bm{v}$.
Theorem 2.1.
  1. (i)
    Every generative conditional factor model (5) has conditional mean and covariance functions of the form (4).
  2. (ii)
    Conversely, for every moment kernel function (3) there exists a generative conditional factor model of the form (5) with conditional mean and covariance functions given by (4).
  3. (iii)
    If $\alpha(z)=0$, the generative conditional factor model (5) can be represented as
    $x_{t+1,i}=\langle\beta(z_{t,i}),f_{t+1}\rangle_{\mathcal{C}}+\epsilon_{t+1,i}$ (6)
    in terms of the linear ${\mathcal{C}}$-valued factors $f_{t+1}\coloneqq(\bm{S}_{t}\beta(\bm{z}_{t}))^{+}\bm{S}_{t}\bm{x}_{t+1}$, where $\bm{S}_{t}$ is the $N_{t}\times N_{t}$-diagonal matrix with diagonal elements $S_{t,ii}\coloneqq\gamma(z_{t,i})^{-1}$ if $\gamma(z_{t,i})>0$ and $S_{t,ii}\coloneqq 1$ otherwise. The residuals given by $\epsilon_{t+1,i}\coloneqq x_{t+1,i}-\langle\beta(z_{t,i}),f_{t+1}\rangle_{\mathcal{C}}$ have zero conditional mean ${\mathbb{E}}_{t}[\epsilon_{t+1,i}]=0$ and are conditionally uncorrelated with $f_{t+1}$.
3 Joint estimation
To develop an objective function suitable for estimating $h=(h^{\rm sy},h^{\rm id})$, we leverage the law of iterated expectations for conditional moments, formulating the estimation problem as a matrix-valued regression:
$\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}=\begin{bmatrix}1&\langle p,h^{\rm sy}(\bm{z}_{t})\rangle_{\mathcal{C}}\\ \langle h^{\rm sy}(\bm{z}_{t}),p\rangle_{\mathcal{C}}&\langle h^{\rm sy}(\bm{z}_{t}),h^{\rm sy}(\bm{z}_{t})\rangle_{\mathcal{C}}+\operatorname{diag}(\|h^{\rm id}(\bm{z}_{t})\|_{\mathcal{C}}^{2})\end{bmatrix}+\bm{E}_{t+1},$
where $\bm{E}_{t+1}$ is a matrix of errors assumed to satisfy ${\mathbb{E}}_{t}[\bm{E}_{t+1}]=\bm{0}$. It is convenient to introduce the notion of a data point $\xi_{t}\coloneq(N_{t},\bm{x}_{t+1},\bm{z}_{t})$, which summarizes the cross section. We can then define the loss function that is natural for estimating conditional first and second moments, [^7] We can easily generalize the weighting in the loss function (7) by any exogenous weights $\nu_{t,i}\in(0,1)$, $0\leq i\leq N_{t}$, such that $\sum_{i}\nu_{t,i}=1$ and set ${\mathcal{L}}(h,\xi_{t})=(N_{t}+1)^{-2}\sum_{0\leq i,j\leq N_{t}}\nu_{t,i}\nu_{t,j}(x_{t+1,i}x_{t+1,j}-q_{h}(z_{t,i},z_{t,j}))^{2}.$ This is captured by (7) simply by replacing the data $x_{t,i}$ by $\nu_{t,i}^{1/2}x_{t,i}$ and $q_{h}(z_{t,i},z_{t,j})$ by $\nu_{t,i}^{1/2}q_{h}(z_{t,i},z_{t,j})\nu_{t,j}^{1/2}$. For example, choosing $\nu_{t,0}\in(0,1)$ and setting $\nu_{t,i}=(1-\nu_{t,0})/N_{t}$ for all $i\geq 1$, allows to balance the weights given to the first and second moment error terms in (7).
${\mathcal{L}}(h,\xi_{t})$ :-$\coloneq(N_{t}+1)^{-2}\left\|\bm{E}_{t+1}\right\|_{F}^{2}:$ (7)
$=2\underbrace{(N_{t}+1)^{-2}\left\|\bm{x}_{t+1}-\langle h^{\rm sy}(\bm{z}_{t}),p\rangle_{\mathcal{C}}\right\|_{2}^{2}}_{\text{first moment error}}$
$\quad+\underbrace{(N_{t}+1)^{-2}\left\|\bm{x}_{t+1}\bm{x}_{t+1}^{\top}-\langle h^{\rm sy}(\bm{z}_{t}),h^{\rm sy}(\bm{z}_{t})^{\top}\rangle_{\mathcal{C}}-\operatorname{diag}(\|h^{\rm id}(\bm{z}_{t})\|_{\mathcal{C}}^{2})\right\|_{F}^{2}}_{\text{second moment error}},$
where $\|\cdot\|_{F}$ and $\|\cdot\|_{2}$ denote the Frobenius and Euclidean norm, respectively.
The flexibility and empirical success of our approach is based on the specification of the feature map $h=(h^{\rm sy},h^{\rm id})$ as an element in a potentially infinite-dimensional hypothesis space ${\mathcal{H}}$. Specifically, we assume that ${\mathcal{H}}={\mathcal{H}}^{\rm sy}\times{\mathcal{H}}^{\rm id}$ is the product space of separable ${\mathcal{C}}$-valued reproducing kernel Hilbert spaces (RKHS) ${\mathcal{H}}^{\rm sy}$, ${\mathcal{H}}^{\rm id}$, consisting of functions $h^{\rm sy},h^{\rm id}:{\mathcal{Z}}\to{\mathcal{C}}$, and with operator-valued reproducing kernels $K^{\rm sy},K^{\rm id}$ on ${\mathcal{Z}}$. For tractability we further assume that the kernels are separable, $K^{\rm sy}(z,z^{\prime})=k^{\rm sy}(z,z^{\prime})I_{\mathcal{C}}$, $K^{\rm id}(z,z^{\prime})=k^{\rm id}(z,z^{\prime})I_{\mathcal{C}}$, for some given scalar reproducing kernels $k^{\rm sy},k^{\rm id}$ of separable RKHS ${\mathcal{G}}^{\rm sy},{\mathcal{G}}^{\rm id}$ on ${\mathcal{Z}}$, so that ${\mathcal{H}}^{\rm sy}\cong{\mathcal{G}}^{\rm sy}\otimes{\mathcal{C}}$, ${\mathcal{H}}^{\rm id}\cong{\mathcal{G}}^{\rm id}\otimes{\mathcal{C}}$ can be identified with tensor product spaces (see Paulsen and Raghupathi, 2016, Chapter 6). Adding penalty terms with $\lambda^{\rm sy},\lambda^{\rm id}\geq 0$ to (7), we obtain the regularized loss function
${\mathcal{R}}(h,\xi_{t})\coloneqq{\mathcal{L}}(h,\xi_{t})+\underbrace{\lambda^{\rm sy}\|h^{\rm sy}\|_{{\mathcal{H}}^{\rm sy}}^{2}+\lambda^{\rm id}\|h^{\rm id}\|_{{\mathcal{H}}^{\rm id}}^{2}}_{\text{regularization}}.$
Finally, taking the sample average, we arrive at the non-standard kernel ridge regression problem,
minimize $\underset{h\in{\mathcal{H}}}{\text{minimize }}\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{R}}(h,\xi_{t}).$ (8)
Notably, problem (8) is not convex in $h\in{\mathcal{H}}$, due to the inner product appearing in the loss function (7). [^8] In fact, for any given $z,z^{\prime}\in{\mathcal{Z}}$, the function $Q:{\mathcal{H}}^{\rm sy}\to{\mathbb{R}}$, $h^{\rm sy}\mapsto Q(h^{\rm sy})=\langle h^{\rm sy}(z),h^{\rm sy}(z^{\prime})\rangle_{\mathcal{C}}$ is neither convex nor concave in $h^{\rm sy}$ in general. We see this by means of the following example. Let $h^{\rm sy}_{1},h^{\rm sy}_{2}\in{\mathcal{H}}^{\rm sy}$ such that $h^{\rm sy}_{1}(z)=0$ and $h^{\rm sy}_{2}(z^{\prime})=0$. Then $Q(h^{{\rm sy}}_{1})=Q(h^{\rm sy}_{2})=0$. On the other hand, for any $s\in(0,1)$, $Q(sh^{\rm sy}_{1}+(1-s)h^{\rm sy}_{2})=(1-s)s\big{\langle}h^{\rm sy}_{2}(z),h^{\rm sy}_{1}(z^{\prime})\big{\rangle}_{\mathcal{C}}$, which could be either positive or negative. It can therefore neither be bounded below nor above by $sQ(h^{\rm sy}_{1})+(1-s)Q(h^{\rm sy}_{2})=0$. It follows that, in general, there are infinitely many solutions $h$ of (8), although they all imply the same optimal moment kernel function $q_{h}$, as we will see in Lemma 3.3.
As a first step towards solving (8), we establish a representer theorem for our non-standard problem, which generalizes Micchelli and Pontil (2005, Theorem 4.1). For further use, we denote the total sample size by $N_{\rm tot}\coloneqq\sum_{t=0}^{T-1}N_{t}$.
Theorem 3.1 (Representer Theorem).
Any minimizer $h=(h^{\rm sy},h^{\rm id})$ of (8) with minimal ${\mathcal{H}}^{\tau}$-norms is of the form
$h^{\tau}(\cdot)=\sum_{t=0}^{T-1}\sum_{i=1}^{N_{t}}k^{\tau}(\cdot,z_{t,i})\gamma^{\tau}_{t,i},\quad\text{for some $\gamma^{\tau}_{t,i}\in{\mathcal{C}}$,}$ (9)
for both components $\tau\in\{{\rm sy},{\rm id}\}$.
Inserting the optimal functional form (9), problem (8) can be equivalently expressed in terms of $N_{\rm tot}$ pairs of coefficients $(\gamma^{{\rm sy}}_{t,i},\gamma^{{\rm id}}_{t,i})\in{\mathcal{C}}\times{\mathcal{C}}$. Although the optimal form (9) is a considerable simplification of the full infinite-dimensional problem, it is generally still computationally infeasible for large $N_{\rm tot}$. In the following we therefore propose a low-rank approximation, along with a reparametrization, of problem (8). This will result in a low-dimensional convex optimization problem, which approximates the original problem.
To this end, we consider the Nyström method (Drineas and Mahoney, 2005), and denote by $\bm{Z}\coloneq[\bm{z}_{t}:0\leq t\leq T-1]\in{\mathcal{Z}}^{N_{\rm tot}}$ the full sample array of characteristics. For each component $\tau\in\{{\rm sy},{\rm id}\}$, we consider a subsample $\Pi^{\tau}\subset\{1,\dots,N_{\rm tot}\}$ of size $m^{\tau}\leq N_{\rm tot}$ that approximates the full kernel matrix such that the trace error
$\epsilon^{\tau}_{\rm approx}\coloneqq\operatorname{tr}\left(k^{\tau}(\bm{Z},\bm{Z}^{\top})-k^{\tau}(\bm{Z},\bm{Z}_{\Pi^{\tau}}^{\top})k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}_{\Pi^{\tau}}^{\top})^{-1}k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}^{\top})\right)$ (10)
is small. This subsample selection is facilitated by a pivoted Cholesky decomposition, see Chen et al. (2023). This yields $m^{\tau}$ linearly independent functions $\phi^{\tau}_{i}(\cdot)$ in ${\mathcal{G}}^{\tau}$, forming an ${\mathbb{R}}^{m^{\tau}}$-valued feature map defined as $\bm{\phi}^{\tau}(\cdot)\coloneqq[\phi^{\tau}_{1}(\cdot),\dots,\phi^{\tau}_{m^{\tau}}(\cdot)]\coloneq k^{\tau}(\cdot,\bm{Z}_{\Pi^{\tau}}^{\top})\bm{B}^{\tau}$, where $\bm{B}^{\tau}$ is an arbitrarily chosen invertible square matrix. [^9] The functions $\bm{\phi}^{\tau}$ are orthonormal in ${\mathcal{G}}^{\tau}$ if and only if $\bm{B}^{\tau}{\bm{B}^{\tau}}^{\top}=k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}_{\Pi^{\tau}}^{\top})^{-1}$. However, this assumption is not imposed here, allowing for the use of flexible, user-defined feature maps and thereby enhancing the modularity of our framework. We restrict problem (8) to the subspace ${\mathcal{H}}_{0}={\mathcal{H}}^{\rm sy}_{0}\times{\mathcal{H}}^{\rm id}_{0}$ of ${\mathcal{H}}$ consisting of functions of the form
$h^{\tau}_{0}(\cdot)=\sum_{i=1}^{m^{\tau}}\phi^{\tau}_{i}(\cdot)\gamma^{\tau}_{i}=\bm{\phi}^{\tau}(\cdot)\bm{\gamma}^{\tau},$ (11)
for coefficients $\bm{\gamma}^{\tau}\coloneqq[\gamma^{\tau}_{1},\dots,\gamma^{\tau}_{m^{\tau}}]^{\top}\in{\mathcal{C}}^{m^{\tau}}$. The following proposition provides a heuristic for assessing the quality of this low-rank approximation. [^10] However, note that the optimizer of problem (8) restricted to $h_{0}\in{\mathcal{H}}_{0}={\mathcal{H}}^{\rm sy}_{0}\times{\mathcal{H}}^{\rm id}_{0}$ is generally not given as orthogonal projection on ${\mathcal{H}}_{0}$ of any optimizer of the unrestricted problem.
Proposition 3.2.
Let $h^{\tau}\in{\mathcal{H}}^{\tau}$ be an arbitrary candidate function of the form (9), and denote by $h^{\tau}_{0}$ its projection on ${\mathcal{H}}^{\tau}_{0}$, which is given by the expression on the right hand side of (9) with the kernel function $k^{\tau}(z,z^{\prime})$ replaced by its projection $k^{\tau}_{0}(z,z^{\prime})\coloneqq k^{\tau}(z,\bm{Z}_{\Pi^{\tau}}^{\top})k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}_{\Pi^{\tau}}^{\top})^{-1}k^{\tau}(\bm{Z}_{\Pi^{\tau}},z^{\prime})=\bm{\phi}^{\tau}(z)\langle{\bm{\phi}^{\tau}}^{\top},\bm{\phi}^{\tau}\rangle_{{\mathcal{G}}^{\tau}}^{-1}\bm{\phi}^{\tau}(z^{\prime})^{\top}$. Then the difference $q_{h}(z,z^{\prime})-q_{h_{0}}(z,z^{\prime})$ is a kernel function, and the aggregated cross-sectional approximation error of the implied conditional moment matrices is bounded by
$\sum_{t=0}^{T-1}\big{\|}q_{h}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})-q_{h_{0}}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})\big{\|}_{F}\leq\sum_{\tau\in\{{\rm sy},{\rm id}\}}\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}^{2}\epsilon_{\rm approx}^{\tau},$ (12)
where we denote the extended covariate array $\bar{\bm{z}_{t}}^{\top}\coloneqq\begin{bmatrix}\Delta&\bm{z}_{t}^{\top}\end{bmatrix}\in\{\Delta\}\times{\mathcal{Z}}^{N_{t}}$.
The regularized loss function ${\mathcal{R}}(h_{0},\xi_{t})$ restricted to feature maps $h_{0}(\cdot)=(h^{\rm sy}_{0}(\cdot),h^{\rm id}_{0}(\cdot))$ of the form (11) can equivalently be expressed in terms of the coefficients $\bm{\gamma}=(\bm{\gamma}^{\rm sy},\bm{\gamma}^{\rm id})\in{\mathcal{C}}^{m^{\rm sy}}\times{\mathcal{C}}^{m^{\rm id}}$ as [^11] We use that the norm of $h^{\tau}$ becomes $\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}^{2}=\sum_{i,j=1}^{m^{\tau}}\langle\phi^{\tau}_{i},\phi^{\tau}_{j}\rangle_{{\mathcal{G}}^{\tau}}\langle\gamma^{\tau}_{i},\gamma^{\tau}_{j}\rangle_{\mathcal{C}}=\operatorname{tr}(\bm{G}^{\tau}\bm{U}^{\tau}(\bm{\gamma}^{\tau}))$.
${\mathcal{R}}(\bm{\gamma},\xi_{t})\coloneqq(N_{t}+1)^{-2}\bigg{\|}\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}-\bm{\Psi}^{\rm sy}(\bm{z}_{t})\bm{U}^{\rm sy}(\bm{\gamma}^{\rm sy})\bm{\Psi}^{\rm sy}(\bm{z}_{t})^{\top}\\ -\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}(\bm{\gamma}^{\rm id})\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})\bigg{\|}_{2}^{2}\\ +\lambda^{\rm sy}\operatorname{tr}(\bm{G}^{\rm sy}\bm{U}^{\rm sy}(\bm{\gamma}^{\rm sy}))+\lambda^{\rm id}\operatorname{tr}(\bm{G}^{\rm id}\bm{U}^{\rm id}(\bm{\gamma}^{\rm id})),$
for the matrices given by
${\bm{U}}^{\rm sy}(\bm{\gamma}^{\rm sy})\coloneqq\begin{bmatrix}1&\langle p,{\bm{\gamma}^{\rm sy}}^{\top}\rangle_{\mathcal{C}}\\ \langle p,{\bm{\gamma}^{\rm sy}}\rangle_{\mathcal{C}}&\langle{\bm{\gamma}^{\rm sy}},{\bm{\gamma}^{\rm sy}}^{\top}\rangle_{\mathcal{C}}\end{bmatrix}\in{\mathbb{S}}^{m^{\rm sy}+1}_{+},\quad{\bm{U}}^{\rm id}(\bm{\gamma}^{\rm id})\coloneqq\langle{\bm{\gamma}^{\rm id}},{\bm{\gamma}^{\rm id}}^{\top}\rangle_{\mathcal{C}}\in{\mathbb{S}}^{m^{\rm id}}_{+},$
$\bm{\Psi}^{\rm sy}(\bm{z}_{t})\coloneqq\begin{bmatrix}1&\bm{0}^{\top}\\ \bm{0}&\bm{\phi}^{\rm sy}(\bm{z}_{t})\end{bmatrix}\in{\mathbb{R}}^{(N_{t}+1)\times(m^{\rm sy}+1)},\quad\bm{\Psi}^{\rm id}(\bm{z}_{t})\coloneqq\begin{bmatrix}\bm{0}^{\top}\\ \bm{\phi}^{\rm id}(\bm{z}_{t})\end{bmatrix}\in{\mathbb{R}}^{(N_{t}+1)\times m^{\rm id}},$
$\bm{G}^{\rm sy}\coloneqq\begin{bmatrix}0&\bm{0}^{\top}\\ \bm{0}&\langle{\bm{\phi}^{\rm sy}}^{\top},\bm{\phi}^{\rm sy}\rangle_{{\mathcal{G}}^{\rm sy}}\end{bmatrix}\in{\mathbb{S}}_{+}^{m^{\rm sy}+1},\quad\bm{G}^{\rm id}\coloneqq\langle{\bm{\phi}^{\rm id}}^{\top},\bm{\phi}^{\rm id}\rangle_{{\mathcal{G}}^{\rm id}}\in{\mathbb{S}}_{+}^{m^{\rm id}}.$
Here $\operatorname{Diag}(\bm{A})\coloneqq\operatorname{diag}(\operatorname{diag}(\bm{A}))$ denotes the matrix-to-diagonal matrix operator, which extracts the diagonal of a square matrix $\bm{A}$ and converts that vector to a conformal diagonal matrix. [^12] We follow the convention of overloading the $\operatorname{diag}(\cdot)$ operator, such that $\operatorname{diag}(\bm{v})$ returns a square diagonal matrix with the elements of vector $\bm{v}$ on the main diagonal, and $\operatorname{diag}(\bm{A})$ returns a column vector of the main diagonal elements of a square matrix $\bm{A}$.
Next, we define the convex feasible set ${\mathcal{D}}\coloneq{\mathcal{D}}^{\rm sy}\times{\mathbb{S}}^{m^{\rm id}}_{+}$ of pairs of matrices $\bm{U}=({\bm{U}^{\rm sy}},{\bm{U}^{\rm id}})$ where
${\mathcal{D}}^{\rm sy}\coloneq\left\{{\bm{U}^{\rm sy}}\in{\mathbb{S}}^{m^{\rm sy}+1}_{+}:\bm{U}^{\rm sy}_{11}=1\right\}.$
The following lemma justifies the subsequent reparametrization.
Lemma 3.3.
  1. (i)
    The mapping ${\bm{U}}^{\rm sy}:{\mathcal{C}}^{m^{\rm sy}}\to{\mathcal{D}}^{\rm sy}$ is surjective if and only if $\dim{\mathcal{C}}\geq m^{\rm sy}+1$. If $\dim{\mathcal{C}}\geq 3$ then for any $\bm{\gamma}^{\rm sy}\in{\mathcal{C}}^{m^{\rm sy}}$ there exists infinitely many $\tilde{\bm{\gamma}}^{\rm sy}\neq\bm{\gamma}^{\rm sy}$ in ${\mathcal{C}}^{m^{\rm sy}}$ such that $\bm{U}^{\rm sy}(\tilde{\bm{\gamma}}^{\rm sy})=\bm{U}^{\rm sy}(\bm{\gamma}^{\rm sy})$.
  2. (ii)
    The mapping ${\bm{U}}^{\rm id}:{\mathcal{C}}^{m^{\rm id}}\to{\mathbb{S}}^{m^{\rm id}}_{+}$ is surjective if and only if $\dim{\mathcal{C}}\geq m^{\rm id}$. If $\dim{\mathcal{C}}\geq 2$ then for any $\bm{\gamma}^{\rm id}\in{\mathcal{C}}^{m^{\rm id}}$ there exists infinitely many $\tilde{\bm{\gamma}}^{\rm id}\neq\bm{\gamma}^{\rm id}$ in ${\mathcal{C}}^{m^{\rm id}}$ such that $\bm{U}^{\rm id}(\tilde{\bm{\gamma}}^{\rm id})=\bm{U}^{\rm id}(\bm{\gamma}^{\rm id})$.
We henceforth assume that $\dim{\mathcal{C}}\geq\max\{m^{\rm sy}+1,m^{\rm id}\}$. It then follows from Lemma 3.3 that we can equivalently reparametrize problem (8) directly in terms of $\bm{U}=(\bm{U}^{\rm sy},\bm{U}^{\rm id})\in{\mathcal{D}}$, such that the regularized loss function
${\mathcal{R}}(\bm{U},\xi_{t})\coloneqq(N_{t}+1)^{-2}\bigg{\|}\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}-\bm{\Psi}^{\rm sy}(\bm{z}_{t})\bm{U}^{\rm sy}\bm{\Psi}^{\rm sy}(\bm{z}_{t})^{\top}\\ -\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})\bigg{\|}_{2}^{2}\\ +\lambda^{\rm sy}\operatorname{tr}(\bm{G}^{\rm sy}\bm{U}^{\rm sy})+\lambda^{\rm id}\operatorname{tr}(\bm{G}^{\rm id}\bm{U}^{\rm id}),$
becomes linear-quadratic and convex in $\bm{U}$. As a result, our estimator of the moment kernel function (3) is determined by ${\bm{U}}=({\bm{U}^{\rm sy}},{\bm{U}^{\rm id}})\in{\mathcal{D}}$ and given by
$q_{\bm{U}}(z,z^{\prime})=\begin{bmatrix}1_{z=\Delta}&\bm{\phi}^{\rm sy}(z)\end{bmatrix}\bm{U}^{\rm sy}\begin{bmatrix}1_{z^{\prime}=\Delta}&\bm{\phi}^{\rm sy}(z^{\prime})\end{bmatrix}^{\top}+\bm{\phi}^{\rm id}(z)\bm{U}^{\rm id}\bm{\phi}^{\rm id}(z^{\prime})^{\top}1_{z=z^{\prime}}.$ (13)
It implies the conditional mean and covariance (COCO) functions (4) given by
$\mu_{\bm{U}}(z)$ $=\bm{\phi}^{\rm sy}(z){\bm{b}},$ (14)
$c_{\bm{U}}(z,z^{\prime})$ $=\bm{\phi}^{\rm sy}(z)\big{(}\bm{V}-{\bm{b}}{\bm{b}}^{\top}\big{)}\bm{\phi}^{\rm sy}(z^{\prime})^{\top}$
$\quad+\bm{\phi}^{\rm id}(z)\bm{U}^{\rm id}\bm{\phi}^{\rm id}(z^{\prime})^{\top}1_{z=z^{\prime}},\qquad\text{for $\begin{bmatrix}1&{\bm{b}}^{\top}\\ {\bm{b}}&\bm{V}\end{bmatrix}\coloneqq{\bm{U}^{\rm sy}}$.}$
For a cross section with data point $\xi_{t}=(N_{t},\bm{x}_{t+1},\bm{z}_{t})$ we thus obtain the COCO estimates
$\bm{\mu}_{t}$ $=\bm{\phi}^{\rm sy}(\bm{z}_{t}){\bm{b}},$ (15)
$\bm{\Sigma}_{t}$ $=\underbrace{\bm{\phi}^{\rm sy}(\bm{z}_{t})\big{(}\bm{V}-{\bm{b}}{\bm{b}}^{\top}\big{)}\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}}_{\eqqcolon\bm{\Sigma}^{\rm sy}_{t}}+\underbrace{\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})}_{\eqqcolon\bm{\Sigma}^{\rm id}_{t}}.$
with systematic and idiosyncratic components $\bm{\Sigma}_{t}^{\rm sy}$ and $\bm{\Sigma}_{t}^{\rm id}$.
It follows from (15) that $\bm{\mu}_{t}\in\operatorname{Im}(\bm{\Sigma}_{t})$ if either $\bm{V}-\bm{b}\bm{b}^{\top}$ or $\bm{\Sigma}_{t}^{\rm id}$ is invertible, an assumption we adopt henceforth. [^13] It is always satisfied in our empirical study below. This implies that the conditional mean-variance efficient portfolio, with weights $\bm{w}_{t}=\bm{\Sigma}_{t}^{+}\bm{\mu}_{t}$, is well-defined and attains the maximum Sharpe ratio, which in annualized terms is given by $\sqrt{12}\times\sqrt{\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{\mu}_{t}}$. [^14] Filipović and Schneider (2024) demonstrate that the conditional mean-variance efficient portfolio can be replicated by trading exclusively in the $m^{\rm sy}$ factor portfolios $\bm{f}_{t+1}$ defined in Theorem 2.1 (iii).
From the COCO estimates (15), we can also deduce the linear factor representation $\bm{x}_{t+1}=\bm{\phi}^{\rm sy}(\bm{z}_{t})\bm{g}_{t+1}+(\bm{\Sigma}_{t}^{\rm id})^{1/2}\xi_{t+1}(\bm{z}_{t})$, with equality in distribution. Here, $\bm{g}_{t+1}$ represents an $m^{\rm sy}$-dimensional systematic risk factor process with constant conditional mean ${\mathbb{E}}_{t}[\bm{g}_{t+1}]=\bm{b}$ and covariance matrix $\operatorname{Cov}_{t}[\bm{g}_{t+1}]=\bm{V}-\bm{b}\bm{b}^{\top}$, and $\xi_{t+1}(\bm{z}_{t})$ is a conditionally uncorrelated white noise process, as specified after (5). This result aligns with and constitutes a special case of Theorem 2.1 (ii).
In the empirical study below we specify the idiosyncratic component as follows.
Example 3.4.
Arguably, the simplest idiosyncratic specification is in dimension $m^{\rm id}=1$, with constant feature map $\bm{\phi}^{\rm id}(\cdot)=\phi^{\rm id}_{1}(\cdot)\coloneqq 1$, and $\bm{U}^{\rm id}=u^{\rm id}\in[0,\infty)$. The idiosyncratic component of the covariance function in (14) becomes $u^{\rm id}1_{z=z^{\prime}}$, and the estimate in (15) reads $\bm{\Sigma}_{t}^{\rm id}=u^{\rm id}\bm{I}_{N_{t}}$.
4 Theoretical properties of the COCO estimator
In this section we prove asymptotic uniqueness, consistency, and finite-sample guarantees for our COCO estimator. For this analysis and further implementation we first vectorize the regularized loss function. We then state our theoretical results in terms of these vectorized parameters.
4.1 Vectorization of the loss function
We use the (half-)vectorization of (symmetric) matrices $\bm{A}\in{\mathbb{R}}^{n\times n}$ defined as
$\operatorname{vec}(\bm{A})$ $\coloneqq[A_{11},A_{21},\dots,A_{n1},A_{21},\dots,A_{nn}]^{\top}\in{\mathbb{R}}^{n^{2}},$
$\operatorname{vech}(\bm{A})$ $\coloneqq[A_{11},A_{21},\dots,A_{n1},A_{22},A_{23},\dots,A_{nn}]^{\top}\in{\mathbb{R}}^{n(n+1)/2},$
as well as the duplication matrix $\bm{D}_{n}\in{\mathbb{R}}^{n^{2}\times n(n+1)/2}$, defined such that $\operatorname{vec}(\bm{A})=\bm{D}_{n}\operatorname{vech}(\bm{A})$ for all $\bm{A}\in{\mathbb{S}}^{n}$. The composition of $\operatorname{vec}$ and $\operatorname{diag}$ can be expressed as $\operatorname{vec}\operatorname{diag}(\bm{v})=\bm{R}_{n}\bm{v}$ for the $n^{2}\times n$-matrix $\bm{R}_{n}$ whose $i$th column is the standard basis vector $\bm{e}_{(i-1)n+i}$ in ${\mathbb{R}}^{n^{2}}$. In turn, $\bm{R}_{n}^{\top}(\bm{A}\otimes\bm{A})$ is the $n\times m^{2}$-matrix whose $i$th row is $\bm{A}_{i,\cdot}\otimes\bm{A}_{i,\cdot}$, for a $n\times m$-matrix $\bm{A}$. Note that $\bm{R}_{n}^{\top}\bm{R}_{n}=\bm{I}_{n}$ and $\bm{R}_{n}\bm{R}_{n}^{\top}$ is the orthogonal projection in ${\mathbb{R}}^{n^{2}}$ on the $n$-dimensional subspace spanned by $\bm{e}_{(i-1)n+i}$, $i=1,\dots,n$.
Using this terminology, we define
$M$ $\coloneqq(m^{\rm sy}+1)(m^{\rm sy}+2)/2+m^{\rm id}(m^{\rm id}+1)/2,$
${\mathcal{U}}$ $\coloneqq\operatorname{vech}({\mathcal{D}})={\mathcal{U}}^{\rm sy}\times{\mathcal{U}}^{\rm id}\subset{\mathbb{R}}^{M},$
$\bm{y}(\bm{x}_{t+1})$ $\coloneqq\operatorname{vec}\bigg{(}\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}\bigg{)}\in{\mathbb{R}}^{(N_{t}+1)^{2}},$
and
${\mathcal{U}}^{\rm sy}$ $\coloneqq\operatorname{vech}({\mathcal{D}}^{\rm sy})\subset{\mathbb{R}}^{(m^{\rm sy}+1)(m^{\rm sy}+2)/2},$ ${\mathcal{U}}^{\rm id}$ $\coloneqq\operatorname{vech}({\mathbb{S}}_{+}^{m^{\rm id}})\subset{\mathbb{R}}^{m^{\rm id}(m^{\rm id}+1)/2},$
$\bm{u}^{\rm sy}$ $\coloneqq\operatorname{vech}(\bm{U}^{\rm sy}),$ $\bm{u}^{\rm id}$ $\coloneqq\operatorname{vech}(\bm{U}^{\rm id}),$
$\bm{g}^{\rm sy}$ $\coloneqq\operatorname{vec}(\bm{G}^{\rm sy})\in{\mathbb{R}}^{(m^{\rm sy}+1)^{2}},$ $\bm{g}^{\rm id}$ $\coloneqq\operatorname{vec}(\bm{G}^{\rm id})\in{\mathbb{R}}^{(m^{\rm id})^{2}}.$
We can then express the regularized loss function ${\mathcal{R}}(\bm{U},\xi_{t})$ in terms of the vectorized parameter $\bm{u}=\begin{bmatrix}\bm{u}^{\rm sy}\\ \bm{u}^{\rm id}\end{bmatrix}$ as
${\mathcal{R}}(\bm{u},\xi_{t})\coloneqq(N_{t}+1)^{-2}\Big{\|}\bm{y}(\bm{x}_{t+1})-(\bm{\Psi}^{\rm sy}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm sy}(\bm{z}_{t}))\bm{D}_{m^{\rm sy}+1}\bm{u}^{\rm sy}\\ -\bm{R}_{N_{t}+1}\bm{R}_{N_{t}+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm id}(\bm{z}_{t}))\bm{D}_{m^{\rm id}}\bm{u}^{\rm id}\Big{\|}_{2}^{2}\\ +\lambda^{\rm sy}{\bm{g}^{\rm sy}}^{\top}\bm{D}_{m^{\rm sy}+1}\bm{u}^{\rm sy}+\lambda^{\rm id}{\bm{g}^{\rm id}}^{\top}\bm{D}_{m^{\rm id}}\bm{u}^{\rm id},$
where we used that $\operatorname{vec}(\bm{\Psi}^{\rm sy}(\bm{z}_{t})\bm{U}^{\rm sy}\bm{\Psi}^{\rm sy}(\bm{z}_{t})^{\top})=(\bm{\Psi}^{\rm sy}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm sy}(\bm{z}_{t}))\operatorname{vec}(\bm{U}^{\rm sy})$ and
$\operatorname{vec}(\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top}))=\bm{R}_{N_{t}+1}\bm{R}_{N_{t}+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm id}(\bm{z}_{t}))\bm{D}_{m^{\rm id}}\bm{u}^{\rm id},$
given the $i$th diagonal element $(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})_{ii}=(\bm{\Psi}_{i,\cdot}^{\rm id}(\bm{z}_{t})\otimes\bm{\Psi}_{i,\cdot}^{\rm id}(\bm{z}_{t}))\operatorname{vec}(\bm{U}^{\rm id})$.
For the simple idiosyncratic specification in dimension $m^{\rm id}=1$ in Example 3.4, the above expression simplifies to
$\operatorname{vec}(\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top}))=\underbrace{\bm{R}_{N_{t}+1}}_{(N_{t}+1)^{2}\times(N_{t}+1)}\underbrace{\begin{bmatrix}0\\ \bm{1}\end{bmatrix}}_{(N_{t}+1)\times 1}u^{\rm id},$
and the regularization penalty term reads $\lambda^{\rm id}{\bm{g}^{\rm id}}^{\top}\bm{D}_{m^{\rm id}}\bm{u}^{\rm id}=\lambda^{\rm id}u^{\rm id}$.
4.2 Uniqueness, consistency, and finite-sample guarantees
Our data points $\xi_{t}=(N_{t},\bm{x}_{t+1},\bm{z}_{t})$ take values in $\Xi\coloneqq\bigcup_{n\in{\mathbb{N}}}\{\{n\}\times{\mathbb{R}}^{n}\times{\mathcal{Z}}^{n}\}$, the union of ranges of cross sections of any size. We assume that our data points $\xi_{t}$, $t=0,\dots,T-1$, are i.i.d. drawn from a distribution ${\mathbb{P}}$ with support in $\Xi$.
We write $\xi=(N,\bm{x},\bm{z})$ for a generic point in $\Xi$. The regularized loss function ${\mathcal{R}}(\bm{u},\xi)$ is linear-quadratic in $\bm{u}\in{\mathbb{R}}^{M}$, and can be expanded as
${\mathcal{R}}(\bm{u},\xi)=\frac{1}{2}\bm{u}^{\top}\bm{A}(\xi)\bm{u}+\bm{b}(\xi)^{\top}\bm{u}+c(\xi)$
for
$\bm{A}(\xi)$ $\coloneqq\nabla_{\bm{u}}^{2}{\mathcal{R}}(\bm{u},\xi)=2(N+1)^{-2}\bm{Q}(\xi)^{\top}\bm{Q}(\xi),$
$\bm{b}(\xi)$ $\coloneqq\nabla_{\bm{u}}{\mathcal{R}}(\bm{0},\xi)=-2(N+1)^{-2}\bm{Q}(\xi)^{\top}\bm{y}+\begin{bmatrix}\lambda^{\rm sy}\bm{D}_{m^{\rm sy}+1}^{\top}{\bm{g}^{\rm sy}}\\ \lambda^{\rm id}\bm{D}_{m^{\rm id}}^{\top}{\bm{g}^{\rm id}}\end{bmatrix},$
$c(\xi)$ $\coloneqq{\mathcal{R}}(\bm{0},\xi)=(N+1)^{-2}\|\bm{y}\|_{2}^{2},$
where we define the $(N+1)^{2}\times M$-matrix
$\bm{Q}(\xi)\coloneqq\begin{bmatrix}\bm{\Psi}^{\rm sy}(\bm{z})\otimes\bm{\Psi}^{\rm sy}(\bm{z})&\bm{R}_{N+1}\bm{R}_{N+1}^{\top}\bm{\Psi}^{\rm id}(\bm{z})\otimes\bm{\Psi}^{\rm id}(\bm{z})\end{bmatrix}\begin{bmatrix}\bm{D}_{m^{\rm sy}+1}&\bm{0}\\ \bm{0}&\bm{D}_{m^{\rm id}}\end{bmatrix}.$
The Hessian matrix $\bm{A}(\xi)$ is positive semidefinite, and hence ${\mathcal{R}}(\bm{u},\xi)$ is convex in $\bm{u}$ in ${\mathbb{R}}^{M}$. It is strictly convex if and only if $\bm{A}(\xi)$ is non-singular, which again holds if and only if $\bm{Q}(\xi)$ is injective. As the duplication matrices $\bm{D}_{m^{\rm sy}+1}$ and $\bm{D}_{m^{\rm id}}$ are injective, a sufficient (but not necessary) condition for $\bm{Q}(\xi)$ to be injective is that the $(m^{\rm sy}+1)^{2}+(m^{\rm id})^{2}$ column vectors of $\bm{\Psi}^{\rm sy}(\bm{z})\otimes\bm{\Psi}^{\rm sy}(\bm{z})$ and $\bm{R}_{N+1}\bm{R}_{N+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z})\otimes\bm{\Psi}^{\rm id}(\bm{z}))$ are jointly linearly independent. Necessary (but not sufficient) for the latter to hold is that $(N+1)^{2}\geq(m^{\rm sy}+1)^{2}+(m^{\rm id})^{2}$ and that both $\bm{\Psi}^{\rm sy}(\bm{z})$ and $\bm{\Psi}^{\rm id}(\bm{z})$ are injective.
We qualify this further in the following lemma. Recall that a function $f(\bm{u})$ is $\alpha$-strongly convex if $f(\bm{u})-(\alpha/2)\|\bm{u}\|_{2}^{2}$ is convex. Denote by $\sigma_{\rm min}(\bm{B})$ the smallest singular value of a matrix $\bm{B}$.
Lemma 4.1.
Define $\bm{P}(\xi)\coloneqq\begin{bmatrix}\bm{\Psi}^{\rm sy}(\bm{z})\otimes\bm{\Psi}^{\rm sy}(\bm{z})&\bm{R}_{N+1}\bm{R}_{N+1}^{\top}\bm{\Psi}^{\rm id}(\bm{z})\otimes\bm{\Psi}^{\rm id}(\bm{z})\end{bmatrix}$. Then
$\sigma_{\rm min}(\bm{A}(\xi))\geq 2(N+1)^{-2}\sigma_{\rm min}(\bm{P}(\xi))^{2}.$ (16)
Assume that
$4(N+1)^{-2}\sigma_{\rm min}(\bm{P}(\xi))^{2}\geq\alpha,\quad\text{for ${\mathbb{P}}$-a.e.\ $\xi\in\Xi$,}$ (17)
for some $\alpha>0$. Then ${\mathcal{R}}(\bm{u},\xi)$ is $\alpha$-strongly convex in $\bm{u}$, for ${\mathbb{P}}$-a.e. $\xi\in\Xi$.
In general, we cannot give a-priori lower bounds on $\sigma_{\rm min}(\bm{P}(\xi))$ in terms of the singular values of $\bm{\Psi}^{\rm sy}(\bm{z})$ and $\bm{\Psi}^{\rm id}(\bm{z})$ alone, as the former depends on the interaction between these two blocks. On the other hand, from the Rayleigh–Ritz Theorem, it follows that
$\sigma_{\rm min}(\bm{P}(\xi))$ $\leq\min\{\sigma_{\rm min}(\bm{\Psi}^{\rm sy}(\bm{z})\otimes\bm{\Psi}^{\rm sy}(\bm{z})),\sigma_{\rm min}(\bm{R}_{N+1}\bm{R}_{N+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z})\otimes\bm{\Psi}^{\rm id}(\bm{z})))\}$
$\leq\min\{\sigma_{\rm min}(\bm{\Psi}^{\rm sy}(\bm{z}))^{2},\sigma_{\rm min}(\bm{\Psi}^{\rm id}(\bm{z}))^{2}\}$
where we used that $(\bm{B}\otimes\bm{B})^{\top}(\bm{B}\otimes\bm{B})=(\bm{B}^{\top}\bm{B})\otimes(\bm{B}^{\top}\bm{B})$ and $\sigma_{\rm min}(\bm{B}\otimes\bm{B})=\sigma_{\rm min}(\bm{B})^{2}$ for any matrix $\bm{B}$, and that $\bm{R}_{N+1}\bm{R}_{N+1}^{\top}$ is an orthogonal projection. Hence in order that (17) holds, it is necessary (but not sufficient) that $\sigma_{\rm min}(\bm{\Psi}^{\rm sy}(\bm{z}))$ and $\sigma_{\rm min}(\bm{\Psi}^{\rm id}(\bm{z}))$ are properly bounded away from zero. [^15] For any matrices $\bm{A}$, $\bm{B}$ with same number of rows, the Rayleigh–Ritz Theorem implies that $\sigma_{\rm min}([\bm{A},\bm{B}])\leq\min\{\sigma_{\rm min}(\bm{A}),\sigma_{\rm min}(\bm{B})\}$. But while the right hand side can be strictly positive, the left hand side may be zero. For example if $\bm{A}=\bm{B}=1$.
We define the sample averages $\bm{A}_{T}\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\bm{A}(\xi_{t})$, $\bm{b}_{T}\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\bm{b}(\xi_{t})$, and $c_{T}\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}c(\xi_{t})$, so that the sample average (empirical) loss is given by
${\mathcal{R}}_{T}(\bm{u})\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{R}}(\bm{u},\xi_{t})=\frac{1}{2}\bm{u}^{\top}\bm{A}_{T}\bm{u}+\bm{b}_{T}^{\top}\bm{u}+c_{T}.$
We next provide conditions under which the population loss is well defined and the law of large numbers applies.
Lemma 4.2.
Assume that the following moments are finite,
${\mathbb{E}}[(N+1)^{-2}\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{4}],\quad{\mathbb{E}}[(N+1)^{-2}\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}^{4}],\quad{\mathbb{E}}[(N+1)^{-2}\|\bm{x}\|_{2}^{4}]<\infty.$ (18)
Then $\|\bm{A}(\xi)\|_{F}$, $\|\bm{b}(\xi)\|_{2}$, and $|c(\xi)|$ have finite expectation, and thus we can define the population loss, along with its gradient and Hessian,
${\mathcal{E}}(\bm{u})$ $\coloneqq{\mathbb{E}}[{\mathcal{R}}(\bm{u},\xi)]=\frac{1}{2}\bm{u}^{\top}{\bm{A}}\bm{u}+{\bm{b}}^{\top}\bm{u}+c,$
$\nabla_{\bm{u}}{\mathcal{E}}(\bm{u})$ $={\mathbb{E}}[\nabla_{\bm{u}}{\mathcal{R}}(\bm{u},\xi)]={\bm{A}}\bm{u}+{\bm{b}},$
$\nabla_{\bm{u}}^{2}{\mathcal{E}}(\bm{u})$ $={\mathbb{E}}[\nabla_{\bm{u}}^{2}{\mathcal{R}}(\bm{u},\xi)]={\bm{A}},$
for ${\bm{A}}\coloneqq{\mathbb{E}}[\bm{A}(\xi)]$, ${\bm{b}}\coloneqq{\mathbb{E}}[\bm{b}(\xi)]$, and $c\coloneqq{\mathbb{E}}[c(\xi)]$. Moreover, the law of large numbers applies and ${\mathcal{R}}_{T}(\cdot)\to{\mathcal{E}}(\cdot)$, $\nabla_{\bm{u}}{\mathcal{R}}_{T}(\cdot)\to\nabla_{\bm{u}}{\mathcal{E}}(\cdot)$, and $\nabla_{\bm{u}}^{2}{\mathcal{R}}_{T}(\cdot)\to\nabla_{\bm{u}}^{2}{\mathcal{E}}(\cdot)$ as $T\to\infty$ uniformly in $\bm{u}$ on compacts in ${\mathbb{R}}^{M}$ with probability 1.
The following is our main result in this section. It is non-standard in statistical learning, as it applies to an estimator that is subject to convex constraints.
Theorem 4.3.
  1. (i)
    Consistency: Assume that (18) holds and that $\bm{A}$ is non-singular, so that ${\mathcal{E}}$ is strictly convex and there exists a unique minimizer $\bm{u}^{\ast}\coloneqq\operatorname*{\arg\!\min}_{\bm{u}\in{\mathcal{U}}}{\mathcal{E}}(\bm{u})$. [^16] Given Jensen’s inequality, $\bm{u}^{\top}\bm{A}\bm{u}\geq\bm{u}^{\top}{\mathbb{E}}[(N+1)^{-1}\bm{Q}(\xi)]^{\top}{\mathbb{E}}[(N+1)^{-1}\bm{Q}(\xi)]\bm{u}$, so that non-singularity of $\bm{A}$ can be asserted by similar arguments as above Lemma 4.1. Then any sequence of minimizers $\bm{u}^{\ast}_{T}\in\operatorname*{\arg\!\min}_{\bm{u}\in{\mathcal{U}}}{\mathcal{R}}_{T}(\bm{u})$ converges, $\bm{u}^{\ast}_{T}\to\bm{u}^{\ast}$ as $T\to\infty$, with probability 1.
  2. (ii)
    Mean squared error bound: Assume further that ${\mathcal{R}}(\bm{u},\xi)$ is $\alpha$-strongly convex in $\bm{u}$ for ${\mathbb{P}}$-a.e. $\xi\in\Xi$, for some $\alpha>0$, see Lemma 4.1, and
    ${\mathbb{E}}[\|(\bm{A}(\xi)-\bm{A})\bm{u}^{\ast}+\bm{b}(\xi)-\bm{b}\|_{2}^{2}]\leq\sigma^{2},$ (19)
    for some $\sigma>0$. Then ${\mathcal{E}}$ and ${\mathcal{R}}_{T}$ are $\alpha$-strongly convex, so that the minimizers $\bm{u}^{\ast}_{T}=\operatorname*{\arg\!\min}_{\bm{u}\in{\mathcal{U}}}{\mathcal{R}}_{T}(\bm{u})$ are unique, and ${\mathbb{E}}[\|\bm{u}^{\ast}_{T}-\bm{u}^{\ast}\|_{2}^{2}]\leq\sigma^{2}/(\alpha^{2}T)$.
  3. (iii)
    Finite-sample guarantees: Assume further that
    ${\mathbb{E}}[\exp(\tau^{-2}\|(\bm{A}(\xi)-\bm{A})\bm{u}^{\ast}+\bm{b}(\xi)-\bm{b}\|_{2}^{2})]\leq\exp(1),$ (20)
    for some $\tau>0$. Then for all $\epsilon>0$, ${\mathbb{P}}[\|\bm{u}^{\ast}_{T}-\bm{u}^{\ast}\|_{2}\geq\epsilon]\leq 2\exp(-\tau^{-2}T\epsilon^{2}\alpha^{2}/3)$. This can equivalently be expressed as: for any $\delta\in(0,1)$, with sample probability of at least $1-\delta$, it holds that
    $\|\bm{u}^{\ast}_{T}-\bm{u}^{\ast}\|_{2}\leq\sqrt{\log(2/\delta)}\sqrt{3}\tau/(\alpha\sqrt{T}).$
  4. (iv)
    Consistency (i), mean squared error bound (ii), and finite-sample guarantees (iii) extend to the implied moment kernel functions (13), and thus the COCO estimates (15), using the fact that
    $\big{|}q_{\bm{u}_{T}^{\ast}}(z,z^{\prime})-q_{\bm{u}^{\ast}}(z,z^{\prime})\big{|}\leq C(z,z^{\prime})\|\bm{u}_{T}^{\ast}-\bm{u}^{\ast}\|_{2},\quad\text{for $z,z^{\prime}\in{\mathcal{Z}}_{\Delta}$,}$ (21)
    where $C(z,z^{\prime})\coloneqq\sqrt{2}\big{(}(1_{z=\Delta}+\|\bm{\phi}^{\rm sy}(z)\|_{2})(1_{z^{\prime}=\Delta}+\|\bm{\phi}^{\rm sy}(z^{\prime})\|_{2})+\|\bm{\phi}^{\rm id}(z)\|_{2}^{2}1_{z=z^{\prime}}\big{)}$.
  5. (v)
    Condition (20) implies (19) for $\sigma^{2}=\tau^{2}$. A sufficient condition for (20) to hold is that $\bm{\phi}^{\rm sy}$ and $\bm{\phi}^{\rm id}$ are uniformly bounded functions on ${\mathcal{Z}}$.
  6. (vi)
    All statements of this theorem hold verbatim if ${\mathcal{U}}$ is replaced by any closed convex subset of ${\mathcal{U}}$.
As an example of a closed convex subset of ${\mathcal{U}}$ mentioned in Theorem 4.3 (vi), consider the block parametrization
$\bm{U}^{{\rm sy}}_{\operatorname{diag}}\coloneq\begin{bmatrix}1&\bm{b}^{\top}\\ \bm{b}&\operatorname{diag}\bm{c}\end{bmatrix},$
for a ${\mathbb{R}}^{m^{\rm sy}}$-vector $\bm{c}$. This parametrization allows replacing the semidefinite constraint $\bm{U}\in{\mathcal{D}}$, which may restrict $m^{{\rm sy}}\leq 100$, due to the quadratic growth of the number of parameters. We show below that the number of quadratic constraints associated with the diagonal specification grows only linearly, which would allow to solve large problems, with essentially unrestricted $m^{{\rm sy}}$. The following result substantiates this claim and provides a constructive description of the quadratic constraints.
Lemma 4.4.
Parameter matrix $\bm{U}^{{\rm sy}}_{\operatorname{diag}}\in{\mathbb{S}}_{+}^{m^{{\rm sy}}+1}$ if and only if $c_{1},\ldots,c_{m^{{\rm sy}}}\geq 0$ and there are parameters $\tilde{c}_{1},\ldots,\tilde{c}_{m^{{\rm sy}}}\geq 0$ such that $\sum_{i=1}^{m^{{\rm sy}}}\tilde{c}_{i}\leq 1$ and $b_{i}^{2}\leq c_{i}\tilde{c}_{i}$. The set ${\mathcal{D}}^{{\rm sy}}_{\operatorname{diag}}\coloneq\{\bm{U}^{{\rm sy}}_{\operatorname{diag}}\in{\mathbb{S}}_{+}^{m^{{\rm sy}}+1}\}\subset{\mathcal{D}}^{{\rm sy}}$ is convex and closed.
5 Empirical study
This section empirically evaluates the COCO estimator. We first describe the data comprising US stock returns (1962–2021), outline the model specifications, and then assess both statistical performance and asset pricing implications.
5.1 Data and model specification
We use unbalanced stock data compiled by Gu et al. (2020a), covering March 1957 to December 2021. This dataset includes approximately 30,000 stocks, with an average of 6,200 stocks per month. It also contains Treasury bill data for calculating monthly excess returns. The dataset comprises 94 stock characteristics (61 updated annually, 13 quarterly, and 20 monthly), 74 industry dummies based on the first two digits of Standard Industrial Classification (SIC) codes, and eight macroeconomic predictors from Welch and Goyal (2008). We restrict the sample to data from 1962 onward, include only common stocks of corporations (sharecodes 10 and 11), and discard months where less than 30% of the characteristics are observed.
Figure 1 displays the number of stocks per month in our sample (in blue) alongside the running average (in red). Early in the sample period, several months have fewer than thirty stock observations. The cross-sectional sample size peaks in the years leading up to 2000, with the running average stabilizing around 4,000 stocks per month toward the end of the sample.
Figure 1: Size of cross section. The blue line shows the number of assets $N_{t}$ over time. The red line shows a running average. The sample consists of stock data compiled by Gu et al. (2020a), covering the period from 1962 to 2021.
We specify the systematic RKHS ${\mathcal{G}}^{\rm sy}$ using the cosine kernel (see Schölkopf and Smola, 2018) and the Gaussian kernel (see Rasmussen and Williams, 2005) given by, [^17] We also implemented Laplace and inverse multi-quadric kernels $k^{lap}(z,z^{\prime})\coloneqq e^{-\|z-z^{\prime}\|_{2}/{\rho_{lap}}}$, $k^{imq}(z,z^{\prime})\coloneqq 1/\sqrt{\|z-z^{\prime}\|_{2}^{2}+\rho_{imq}}$, which perform similarly. Results are available upon request.
$k^{cos}(z,z^{\prime})$ $\coloneqq\frac{\langle z,z^{\prime}\rangle_{2}}{\|z\|_{2}\,\|z^{\prime}\|_{2}},$ $\quad k^{gauss}(z,z^{\prime})$ $\coloneqq e^{-\frac{\|z-z^{\prime}\|^{2}_{2}}{2\,\rho_{gauss}}}.$
The cosine (or correlation) kernel is a finite-dimensional quasi-linear kernel with no hyperparameters, while the Gaussian kernel is non-linear, generating an infinite-dimensional space of smooth, rapidly decaying functions and includes a length-scale hyperparameter, $\rho_{gauss}>0$. We specify the idiosyncratic RKHS ${\mathcal{G}}^{\rm id}$ using the simplest configuration, with dimension $m^{\rm id}=1$, as outlined in Example 3.4.
[…]
To this end, we use the statistical scoring rule $\mathcal{R}_{c}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}\to\mathbb{R}$ proposed by Dawid and Sebastiani (1999),
$\mathcal{R}_{c}(\bm{x},\bm{\mu},\bm{\Sigma})\coloneqq\log\det\bm{\Sigma}+(\bm{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu}),$ (22)
which we evaluate for each validation month $t+1$ using observed returns $\bm{x}_{t+1}$ and our COCO estimates in (15) at time $t$. To compute $\bm{\Sigma}_{t}^{-1}$, we use the Woodbury formula,
$\bm{\Sigma}_{t}^{-1}=\frac{1}{u^{{\rm id}}}\left[\bm{I}_{N_{t}}-\bm{\phi}^{\rm sy}(\bm{z}_{t})\left((\bm{V}-\bm{b}\bm{b}^{\top})^{-1}+\frac{\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}\bm{\phi}^{\rm sy}(\bm{z}_{t})}{u^{{\rm id}}}\right)^{-1}\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}\frac{1}{u^{{\rm id}}}\right],$
exploiting the fact that $\bm{\Sigma}_{t}^{\rm id}$ is diagonal and full-rank for $u^{{\rm id}}>0$. Additionally, we efficiently compute the determinant in (22) using the formula by Sylvester (1851),
$\det\bm{\Sigma}_{t}=(u^{\rm id})^{N_{t}-m}\det\big{(}\bm{I}_{m}u^{\rm id}+\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}\bm{\phi}^{\rm sy}(\bm{z}_{t})(\bm{V}-\bm{b}\bm{b}^{\top})\big{)}.$
For model evaluation, we use an eight-year training window (96 months) and one month for validation, with all out-of-sample tests conducted on the first month following the validation month. Leveraging the high computational efficiency of the procedure in Section 2, we roll the training, validation, and test windows forward each month, iteratively repeating the training, validation, and testing steps.
5.2 Statistical performance
We assess the statistical performance of the COCO model by comparing it to a simple benchmark model, as no established benchmark exists for jointly estimating conditional first and second moments for unbalanced panels. Our benchmark is a purely idiosyncratic model with zero mean and constant covariance, defined as
$\bm{\mu}_{t}^{\text{const}}\coloneqq\mathbb{E}_{t}^{\text{const}}[\bm{x}_{t+1}]=\bm{0},\quad\bm{\Sigma}_{t}^{\text{const}}\coloneqq\sigma^{2}_{\text{const}}\bm{I}_{N_{t}},$ (23)
with a single parameter, $\sigma^{2}_{\text{const}}$, to be estimated. This model is nested within (3), with zero systematic component ($h^{\rm sy}=0$) and a one-dimensional idiosyncratic specification ($m^{{\rm id}}=1$). The minimizer of (8) for this idiosyncratic specification (23) has a closed-form solution,
$\sigma^{2}_{\text{const}}=\frac{\sum_{t=0}^{T-1}\left\{(N_{t}+1)^{-2}\|\bm{x}_{t+1}\|_{2}^{2}\right\}}{\sum_{t=0}^{T-1}\left\{(N_{t}+1)^{-2}N_{t}\right\}}.$
We compare out-of-sample realizations to conditional moments, evaluating first and second moments separately before jointly assessing them with the scoring rule (22). For first moments, we use the predictive out-of-sample $R$-squared measure,
$R^{2}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}(N_{s}+1)^{-2}\|\bm{x}_{s+1}-\bm{\phi}^{\rm sy}(\bm{z}_{s}){\bm{b}}\|_{2}^{2}}{\sum_{s=t}^{T-1}(N_{s}+1)^{-2}\|\bm{x}_{s+1}\|_{2}^{2}},$ (24)
where both numerator and denominator are weighted by $(N_{s}+1)^{-2}$ as in the first-moment error component of (7). For second moments, we compute an out-of-sample predictive $R$-squared measure as,
$R^{2,2}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}(N_{s}+1)^{-2}\|\bm{x}_{s+1}\bm{x}_{s+1}^{\top}-\bm{\phi}^{\rm sy}(\bm{z}_{s})\bm{V}\bm{\phi}^{\rm sy}(\bm{z}_{s})^{\top}-u^{{\rm id}}\bm{I}_{N_{s}}\|_{F}^{2}}{\sum_{s=t}^{T-1}(N_{s}+1)^{-2}\|\bm{x}_{s+1}\bm{x}_{s+1}^{\top}-\sigma^{2}_{\text{const}}\bm{I}_{N_{s}}\|_{F}^{2}},$ (25)
aligning with the second-moment error component in (7). Note that the parameters $\bm{b}$, $\bm{V}$, $u^{{\rm id}}$, $\sigma^{2}_{\text{const}}$, and the feature maps $\bm{\phi}^{\rm sy}$ in (24) and (25) in fact vary with $s$, as they are re-estimated and updated with each rolling training and validation window.
Figure 2 shows the out-of-sample $R^{2}$ over time. The top row displays a rolling average over $24$ months, while the bottom row shows expanding averages. Results indicate high persistence, with a slightly positive $R^{2}$ on average, as seen in the expanding averages. Higher-$m$ specifications tend to perform slightly worse than lower-$m$ ones. Figure 2 also highlights four major stock market crashes (defined by Adrian et al. (2023) from pre-crash peak to post-crash trough): the 1987 Crash (08/1987–12/1987), the Dot-Com Bubble (03/2000–10/2002), the Global Financial Crisis (10/2007–03/2009), and the COVID-19 Pandemic (02/2020–03/2020). No clear pattern is observed in $R^{2}$ across these crashes, with positive $R^{2}$ during the first two and negative during the last two.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
Figure 3 presents the corresponding out-of-sample $R^{2,2}$ over time, showing strong persistence with higher-$m$ specifications outperforming lower-$m$ ones across both kernels. The idiosyncratic specification performs better leading up to the Global Financial Crisis, after which the systematic specification shows marked improvement. Positive $R^{2,2}$ is observed during the 1987 Crash, the Dot-Com Bubble, and the COVID-19 Pandemic, with an overall positive $R^{2,2}$ across the sample period.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
The COCO model jointly estimates first and second moments. To evaluate the joint fit across both moments, we use the scoring rule ${\mathcal{R}}_{c}$ defined in (22) from Dawid and Sebastiani (1999), which is designed for this purpose. To assess the added value of the systematic specification over a purely idiosyncratic model, we define the scoring loss differential as
${\mathcal{R}}_{t,T,\text{OOS}}\coloneqq\frac{1}{T-t}\sum_{s=t}^{T-1}\big{(}{\mathcal{R}}_{c}(\bm{x}_{s+1},\bm{\mu}_{s},\bm{\Sigma}_{s})-{\mathcal{R}}_{c}(\bm{x}_{s+1},\bm{0},\sigma^{2}_{\text{const},s}\bm{I}_{N_{s}})\big{)}.$ (26)
Figure 4 shows that our model incorporating both systematic and idiosyncratic risk consistently outperforms the purely idiosyncratic benchmark (23) over most periods, with the exception of the sample’s initial phase. Differences between higher-$m$ and lower-$m$ specifications are minor. Although the idiosyncratic benchmark may outperform the full model for individual moments in certain periods, the joint assessment consistently favors the full specification, underscoring the statistical value in the COCO estimator. Expanding-window results further validate this, showing a uniform preference for the full model across time and kernel specifications. The scoring rule differential between the full and idiosyncratic models is especially pronounced during the four crisis periods.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss (expanding)
[…]
Following up on the factor representation discussed after (15), where the systematic risk factors $\bm{g}_{t+1}$ are typically latent, we apply Theorem 2.1 (iii) to derive the portfolio factor representation $\bm{x}_{t+1}=\bm{\phi}^{\rm sy}(\bm{z}_{t})\bm{f}_{t+1}+\bm{\epsilon}_{t+1}$. Here, the tradable systematic risk factors are given by $\bm{f}_{t+1}=\bm{\phi}^{\rm sy}(\bm{z}_{t})^{+}\bm{x}_{t+1}$, and the conditionally uncorrelated residuals are implicitly defined by $\bm{\epsilon}_{t+1}=\bm{x}_{t+1}-\bm{\phi}^{\rm sy}(\bm{z}_{t})\bm{f}_{t+1}$. We evaluate the explanatory power of these observable systematic factors through the explained cross-sectional variation, measured by the total $R^{2}$,
$R^{2,\bm{f}}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}\|\bm{x}_{s+1}-\bm{\phi}^{\rm sy}(\bm{z}_{s})\bm{f}_{s+1}\|^{2}_{2}}{\|\bm{x}_{s+1}\|_{2}^{2}}.$ (27)
For the same reasons outlined below (25), the feature maps $\bm{\phi}^{\rm sy}$ in (27) also vary with $s$. Figure 5 presents the out-of-sample total $R^{2}$ over time, which is significantly positive, reaching up to 40% in the early sample period and maintaining a running average of approximately 20%. The explained variation is particularly elevated during market crashes, aligning with the scoring rules and underscoring the significance of the systematic components during periods of market turbulence.
(a) Cosine kernel
(b) Gauss kernel
[…]
The high explained variation observed for the systematic factors motivates a closer examination of the contributions of systematic and idiosyncratic components to the conditional covariance estimates. To quantify this, we compute the ratio $\frac{\operatorname{tr}\bm{\Sigma}^{\rm sy}_{t}}{\operatorname{tr}\bm{\Sigma}_{t}}$, which represents the proportion of systematic variance relative to total variance in our model. This ratio serves as a measure of the predicted explained variation attributable to the systematic factors. Figure 6 reveals that the systematic component was most pronounced early in the sample period, initially exceeding 50% before settling to a running average of approximately 20%, consistent with Figure 5. Similar to the observed explained variation by systematic factors, the systematic component intensifies during market crashes, underscoring its importance during such periods.
(a) Cosine kernel
(b) Gauss kernel
[…]
5.3 Asset pricing implications
Given the strong statistical performance of the COCO estimator, we next assess its effectiveness in an out-of-sample portfolio setting. The literature has documented that mean-variance efficient portfolios often underperform out-of-sample, especially when compared to the naive $1/N_{t}$ portfolio (Basak et al., 2009). Such underperformance is frequently attributed to inaccuracies in estimated moments. Here, we revisit the conditional mean-variance efficient (cMVE) portfolio problem, utilizing conditional means and covariances from the COCO estimator, as detailed following (15). All Sharpe ratios reported below are annualized to facilitate comparison with existing literature.
Figure 7 displays time series of the predicted maximum Sharpe ratios, calculated as $\sqrt{12}\times\sqrt{\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{\mu}_{t}}$, for the two kernels considered in this study. For our model specification, which accounts for both systematic and idiosyncratic risk, the predicted maximum Sharpe ratio is generally positive, whereas the purely idiosyncratic benchmark model in (23) predicts it to be zero, as the conditional mean is identically zero in that case. Both panels in Figure 7 show a natural ordering, with higher values of $m$ yielding higher Sharpe ratios across all data points. The levels generated by the two kernels are comparable, with the cosine kernel achieving peaks of a predicted maximum Sharpe ratio of six for $m=40$. For $m=5$, the smallest predicted maximum Sharpe ratio averages around one. Notably, there are no distinct patterns observed during major market crashes. We next analyze how these predicted Sharpe ratios translate into realized Sharpe ratios.
(a) Cosine kernel
(b) Gauss kernel
[…]
In Figure 8, we plot rolling estimates of out-of-sample realized Sharpe ratios from the cMVE portfolios with monthly excess returns $\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{x}_{t+1}$ over time, alongside the Sharpe ratios from $1/N_{t}$ portfolios. Higher values of $m$ tend to generate higher realized Sharpe ratios, with peaks exceeding five for both kernels. Remarkably, these realized Sharpe ratios remain high despite only modest evidence of predictability in the first and second moments, underscoring the quality of the joint moment estimates, as also suggested by the scoring rule results in Figure 4. The $1/N_{t}$ portfolios generally exhibit lower Sharpe ratios, showing little correlation with those implied by the cMVE portfolios. The bottom row of Figure 8 shows Sharpe ratios estimated over expanding windows, consistently positive toward the end of the sample, with some values exceeding two for both kernels. While it remains challenging to pinpoint an optimal number of factors, both kernels perform least well with $m=5$, which nonetheless outperforms the $1/N_{t}$ portfolio. The largest Sharpe ratio declines occur during the Global Financial Crisis.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
m=5 m=10 m=20 m=40
(Intercept) $0.13^{***}$ $0.30^{***}$ $0.47^{***}$ $0.75^{***}$
$(0.02)$ $(0.03)$ $(0.04)$ $(0.05)$
Mkt $3.58^{***}$ $4.00^{***}$ $5.25^{***}$ $6.02^{***}$
$(0.47)$ $(0.69)$ $(0.90)$ $(1.23)$
SMB $1.42^{*}$ $3.49^{***}$ $4.80^{***}$ $3.44$
$(0.68)$ $(1.01)$ $(1.31)$ $(1.80)$
HML $2.61^{**}$ $0.91$ $0.13$ $0.30$
$(0.86)$ $(1.27)$ $(1.65)$ $(2.26)$
RMW $2.89^{**}$ $3.10^{*}$ $6.19^{***}$ $5.34^{*}$
$(0.91)$ $(1.35)$ $(1.76)$ $(2.40)$
CMA $5.91^{***}$ $8.89^{***}$ $11.55^{***}$ $11.16^{**}$
$(1.38)$ $(2.03)$ $(2.64)$ $(3.62)$
Adj. R2 $0.16$ $0.10$ $0.10$ $0.05$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
(a) Cosine kernel
m=5 m=10 m=20 m=40
(Intercept) $0.10^{***}$ $0.29^{***}$ $0.48^{***}$ $0.75^{***}$
$(0.02)$ $(0.03)$ $(0.04)$ $(0.07)$
Mkt $3.00^{***}$ $2.58^{***}$ $2.46^{**}$ $3.28^{*}$
$(0.37)$ $(0.62)$ $(0.90)$ $(1.54)$
SMB $1.72^{**}$ $3.94^{***}$ $3.57^{**}$ $1.44$
$(0.54)$ $(0.91)$ $(1.32)$ $(2.26)$
HML $2.13^{**}$ $0.37$ $-0.67$ $-0.71$
$(0.67)$ $(1.14)$ $(1.65)$ $(2.83)$
RMW $1.14$ $4.86^{***}$ $6.23^{***}$ $8.19^{**}$
$(0.72)$ $(1.21)$ $(1.76)$ $(3.01)$
CMA $5.79^{***}$ $10.67^{***}$ $10.88^{***}$ $8.71$
$(1.08)$ $(1.83)$ $(2.65)$ $(4.53)$
Adj. R2 $0.20$ $0.13$ $0.06$ $0.01$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
(b) Gaussian kernel
Table 1: Fama–French 5 factors and cMVE portfolio. The table shows results from a time-series regression of out-of-sample cMVE portfolio excess returns, $\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{x}_{t+1}$, on the five factors from Fama and French (2015). The analysis is based on unbalanced US common stock excess returns and associated characteristics from 1962 to 2021.
In summary, the higher-$m$ specifications that also generate higher out-of-sample Sharpe ratios, are largely unrelated to the Fama–French five factors. The findings in this section underscore the effectiveness of the COCO model in an asset pricing context, highlighting the substantial predictive value of the joint COCO estimates.
6 Conclusion
We introduce a nonparametric kernel-based estimator for jointly modeling conditional mean and covariance matrices in large unbalanced panels, addressing critical challenges in econometric analysis of high-dimensional data. Our proposed estimator, the joint conditional mean and covariance (COCO) estimator, is rigorously developed with consistency and finite-sample guarantees, ensuring reliability in both theory and practice. The COCO estimator constructs symmetric, positive semidefinite conditional covariance matrices in all states, leveraging infinite-dimensional hypothesis spaces to flexibly adapt to complex dependencies in data.
Empirically, we apply the COCO estimator to an extensive panel of US stock returns from 1962 to 2021, conditioning on macroeconomic and firm-specific covariates to obtain time-varying expected returns and covariances. Results demonstrate its strong statistical performance and relevance for asset pricing, generating conditional mean-variance efficient portfolios with substantial out-of-sample Sharpe ratios that significantly exceed those of equal-weighted benchmarks. The COCO estimator’s computational efficiency and adaptability underscore its potential for large-scale, reproducible empirical studies, providing a valuable tool for econometricians tackling complex data structures in finance and beyond.
Appendix A Proofs
This appendix contains all proofs.
A.1 Proof of Theorem 2.1
(i): Without loss of generality, we can assume that $b=0$; if not, we simply replace $g_{t+1}$ by $g_{t+1}-b$ and $\alpha(z)$ by $\alpha(z)+\langle\beta(z),b\rangle_{\mathcal{C}}$. We then incorporate $\alpha(z_{t,i})$ into the scalar product in (5) by extending $g_{t+1}$ with an orthogonal unit vector $q\in{\mathcal{C}}$, such that $\langle g_{t+1},q\rangle_{\mathcal{C}}=0$, $\langle\beta(z),q\rangle_{\mathcal{C}}=0$ for all $z\in{\mathcal{Z}}$, and $\|q\|_{\mathcal{C}}=1$. Such a vector $q$ always exists by assumption. Consequently, we can express $\alpha(z)+\langle\beta(z),g_{t+1}\rangle_{\mathcal{C}}=\langle\alpha(z)q+\beta(z),q+g_{t+1}\rangle_{\mathcal{C}}$. With regard to the extension (2), we extend $\alpha$, $\beta$ and $\gamma$ to ${\mathcal{Z}}_{\Delta}$ by setting them to zero for $z=\Delta$. Additionally, we introduce the auxiliary index $i=0$ by defining $x_{t+1,0}\coloneqq 1$ and $z_{t,0}\coloneqq\Delta$, and include the indicator function $1_{z=\Delta}$. This leads to the consistent extension of (5) given by
$x_{t+1,i}=\langle\alpha(z_{t,i})q+\beta(z_{t,i})+q1_{z_{t,i}=\Delta},q+g_{t+1}\rangle_{\mathcal{C}}+\gamma(z_{t,i})\xi_{t+1}(z_{t,i}).$ (28)
As a result, the conditional first and second moments are given by
${\mathbb{E}}_{t}[x_{t+1,i}\,x_{t+1,j}]=\Big{\langle}\big{(}Q+q\otimes q\big{)}\big{(}\alpha(z_{t,j})q+\beta(z_{t,j})+q1_{z_{t,j}=\Delta}\big{)},\\ \alpha(z_{t,i})q+\beta(z_{t,i})+q1_{z_{t,i}=\Delta}\Big{\rangle}_{\mathcal{C}}+\gamma(z_{t,i})^{2}1_{z_{t,i}=z_{t,j}},$ (29)
for all $i,j=0,\dots,N_{t}$. A simple check shows that (29) is perfectly captured by (3) or, equivalently, by (4), where we set $p\coloneqq(Q+q\otimes q)^{1/2}q$, $h^{\rm sy}(z)\coloneqq(Q+q\otimes q)^{1/2}(\alpha(z)q+\beta(z))$, and let $h^{\rm id}$ be such that $\|h^{\rm id}(z)\|_{\mathcal{C}}=\gamma(z)$. Note that $p$ is a unit vector, as $Qq=0$.
(ii): Conversely, let the moment kernel function $q_{h}(z,z^{\prime})$ be given in terms of a unit vector $p\in{\mathcal{C}}$ and feature maps $h^{\rm sy},h^{\rm id}:{\mathcal{Z}}\to{\mathcal{C}}$ as in (3). Define $\alpha(z)\coloneqq\langle h^{\rm sy}(z),p\rangle_{\mathcal{C}}$, $\beta(z)\coloneqq h^{\rm sy}(z)-\alpha(z)p$, $\gamma(z)\coloneqq\|h^{\rm id}(z)\|_{\mathcal{C}}$, and let $\{\eta_{t+1,i}:i=1,2,\dots\}$ and $\{\xi_{t+1}(z):z\in{\mathcal{Z}}\}$ be conditionally uncorrelated white noise processes with conditional mean zero and conditional variance one. Let $e_{0}\coloneqq p,e_{1},e_{2},\dots$ be an orthonormal basis of ${\mathcal{C}}$, and define $g_{t+1}\coloneqq\sum_{i\geq 1}e_{i}\eta_{t+1,i}$ and $q\coloneqq p$. Then $g_{t+1}$ has a constant conditional covariance operator given by $Qp=0$ and $Qe_{i}=e_{i}$ for $i=1,2,\dots$. It can now be easily verified that the right hand side of (29) equals $q_{h}(z_{t,i},z_{t,j})$, as desired.
(iii): This follows from (28) as proved in (Filipović and Schneider, 2024, Proposition 6.3), where also the formal expressions are given for the conditional mean and covariance of $f_{t+1}$ and $\epsilon_{t+1}$. Note that, in contrast to $g_{t+1}$ and the idiosyncratic risk in (5), the factors $f_{t+1}$ are not stationary and the conditional covariance matrix of $\epsilon_{t+1}$ is not diagonal and does not have full rank.
A.2 Proof of Theorem 3.1
Define the linear sample operator $S^{\tau}:{\mathcal{H}}^{\tau}\to{\mathcal{C}}^{N_{\rm tot}}$ by
$S^{\tau}h^{\tau}\coloneqq[h^{\tau}(z_{t,i}):i=1,\dots,N_{t},\,t=0,\dots,T-1].$
We claim that its adjoint ${S^{\tau}}^{\ast}\bm{\gamma}^{\tau}$ is given by the right hand side of (9), for $\bm{\gamma}^{\tau}=[\gamma_{t,i}^{\tau}:i=1,\dots,N_{t},\,t=0,\dots,T-1]$. Indeed, let $f\in{\mathcal{G}}^{\tau}$ and $v\in{\mathcal{C}}$, then
$\langle{S^{\tau}}^{\ast}\bm{\gamma}^{\tau},f\otimes v\rangle_{{\mathcal{H}}^{\tau}}$ $=\langle\bm{\gamma}^{\tau},{S^{\tau}}(f\otimes v)\rangle_{{\mathcal{C}}^{N_{\rm tot}}}=\sum_{t=0}^{T-1}\sum_{i=1}^{N_{t}}\langle k^{\tau}(\cdot,z_{t,i}),f\rangle_{{\mathcal{G}}^{\tau}}\langle\gamma^{\tau}_{t,i},v\rangle_{\mathcal{C}}$
$=\Big{\langle}\sum_{t=0}^{T-1}\sum_{i=1}^{N_{t}}k^{\tau}(\cdot,z_{t,i})\otimes\gamma^{\tau}_{t,i},f\otimes v\Big{\rangle}_{{\mathcal{H}}^{\tau}},$
which proves the claim. We define by ${\mathcal{G}}_{1}^{\tau}$ the subspace in ${\mathcal{G}}^{\tau}$ spanned by $\{k^{\tau}(\cdot,z_{t,i}):i=1,\dots,N_{t},\,t=0,\dots,T-1\}$. It has finite dimension, $\dim({\mathcal{G}}_{1}^{\tau})\leq N_{\rm tot}$, and thus is closed in ${\mathcal{G}}^{\tau}$. Hence $\operatorname{Im}({S^{\tau}}^{\ast})={\mathcal{G}}_{1}^{\tau}\otimes{\mathcal{C}}$ is a closed subspace in ${\mathcal{H}}^{\tau}$. Consequently, ${\mathcal{H}}^{\tau}=\ker(S^{\tau})\oplus\operatorname{Im}({S^{\tau}}^{\ast})$. Now let $h=(h^{\rm sy},h^{\rm id})$ be any minimizer of (8), and decompose $h^{\tau}=h^{\tau}_{0}+h^{\tau}_{1}$ with $h^{\tau}_{0}\in\ker(S^{\tau})$ and $h^{\tau}_{1}\in\operatorname{Im}({S^{\tau}}^{\ast})$. Clearly, the loss function ${\mathcal{L}}(h,\xi_{t})={\mathcal{L}}(Sh,\xi_{t})$ is a function of $Sh=(S^{\rm sy}h^{\rm sy},S^{\rm id}h^{\rm id})=h_{1}$ only. On the other hand, the norm $\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}\geq\|h^{\tau}_{1}\|_{{\mathcal{H}}^{\tau}}$ is greater than or equal for $h^{\tau}$ than for $h^{\tau}_{1}$, with equality if and only if $h^{\tau}_{0}=0$. This completes the proof.
A.3 Proof of Proposition 3.2
We have ${\mathcal{H}}^{\tau}_{0}\cong{\mathcal{G}}^{\tau}_{0}\otimes{\mathcal{C}}$, where ${\mathcal{G}}^{\tau}_{0}$ denotes the subspace of ${\mathcal{G}}^{\tau}$ spanned by $\bm{\phi}^{\tau}$. In the following, without loss of generality, we assume that the functions $\bm{\phi}^{\tau}$ are orthonormal in ${\mathcal{G}}^{\tau}$, otherwise we simply replace them by $\bm{\phi}^{\tau}\langle{\bm{\phi}^{\tau}}^{\top},\bm{\phi}^{\tau}\rangle_{{\mathcal{G}}^{\tau}}^{-1/2}$. We extend $\bm{\phi}^{\tau}$ to an orthonormal basis $\bm{\psi}^{\tau}=[\psi^{\tau}_{1}\coloneqq\phi^{\tau}_{1},\dots,\psi^{\tau}_{m^{\tau}}\coloneqq\phi^{\tau}_{m^{\tau}},\psi^{\tau}_{m^{\tau}+1},\dots,\psi^{\tau}_{M^{\tau}}]$ of ${\mathcal{G}}^{\tau}_{1}$, the subspace of ${\mathcal{G}}^{\tau}$ spanned by $k^{\tau}(\cdot,\bm{Z})$ with $m^{\tau}\leq{M^{\tau}}\coloneqq\dim({\mathcal{G}}^{\tau})\leq N_{\rm tot}$, as in the proof of Theorem 3.1. Accordingly, we have $k^{\tau}(\bm{Z},\bm{Z}^{\top})=\bm{\psi}^{\tau}(\bm{Z})\bm{\psi}^{\tau}(\bm{Z})^{\top}$, and by the same token $k_{0}(z,z^{\prime})=\bm{\phi}^{\tau}(z)\bm{\phi}^{\tau}(z^{\prime})^{\top}$, see Paulsen and Raghupathi (2016, Theorem 2.10). Any candidate function of the form (9) can thus be written as $h^{\tau}(z)=\bm{\psi}^{\tau}(z)\bm{\gamma}^{\tau}$, and its projection on ${\mathcal{H}}^{\tau}_{0}$ is given by $h_{0}^{\tau}(z)=\begin{bmatrix}\bm{\phi}^{\tau}(z)&\bm{0}^{\top}\end{bmatrix}\bm{\gamma}^{\tau}$, for a coefficient array $\bm{\gamma}^{\tau}\in{\mathcal{C}}^{M^{\tau}}$. Consequently, $q_{h}(z,z^{\prime})-q_{h_{0}}(z,z^{\prime})$ is a kernel function. As $\|\bm{A}\|_{F}\leq\operatorname{tr}(\bm{A})$ for any positive semidefinite matrix $\bm{A}$, the cross-sectional approximation errors of the implied conditional moment matrices $q_{h}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})$ can therefore be bounded by the respective trace errors. Concretely, let $E_{h}$ denote the left hand side of (12) and define $\bm{V}^{\tau}\coloneqq\langle\bm{\gamma}^{\tau},{\bm{\gamma}^{\tau}}^{\top}\rangle_{\mathcal{C}}$. Then
$E_{h}$ $\leq\sum_{t=0}^{T-1}\operatorname{tr}\big{(}q_{h}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})-q_{h_{0}}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})\big{)}=\operatorname{tr}(q_{h}(\bm{Z},\bm{Z}^{\top}))-\operatorname{tr}(q_{h_{0}}(\bm{Z},\bm{Z}^{\top}))$ (30)
$=\sum_{\tau\in\{{\rm sy},{\rm id}\}}\operatorname{tr}(\bm{\psi}^{\tau}(\bm{Z})\bm{V}^{\tau}\bm{\psi}^{\tau}(\bm{Z})^{\top})-\operatorname{tr}\bigg{(}\begin{bmatrix}\bm{\phi}^{\tau}(\bm{Z})&\bm{0}^{\top}\end{bmatrix}\bm{V}^{\tau}\begin{bmatrix}\bm{\phi}^{\tau}(\bm{Z})^{\top}\\ \bm{0}\end{bmatrix}\bigg{)}$
$\leq\sum_{\tau\in\{{\rm sy},{\rm id}\}}\|\bm{V}^{\tau}\|_{2}\underbrace{\big{(}\operatorname{tr}(\bm{\psi}^{\tau}(\bm{Z})\bm{\psi}^{\tau}(\bm{Z})^{\top})-\operatorname{tr}(\bm{\phi}^{\tau}(\bm{Z})\bm{\phi}^{\tau}(\bm{Z})^{\top})\big{)}}_{=\epsilon^{\tau}_{\rm approx}}$
where we used that $\operatorname{tr}(\bm{B}\bm{A}\bm{B}^{\top})=\operatorname{tr}(\bm{A}\bm{B}^{\top}\bm{B})\leq\|\bm{A}\|_{2}\operatorname{tr}(\bm{B}^{\top}\bm{B})=\|\bm{A}\|_{2}\operatorname{tr}(\bm{B}\bm{B}^{\top})$ for any positive semidefinite matrix $\bm{A}$ and conformal matrix $\bm{B}$. The bound (12) now follows because $\|\bm{V}^{\tau}\|_{2}\leq\operatorname{tr}(\bm{V}^{\tau})=\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}^{2}$, which completes the proof.
Note that the last inequality in (30) is tight, with equality for, e.g., $\bm{V}^{\tau}=\bm{I}_{M^{\tau}}$. This shows, as a side result, that $\epsilon^{\rm sy}_{\rm approx}+\epsilon^{\rm id}_{\rm approx}$ equals the worst case approximation error, when we take the maximum over all coefficients $\bm{\gamma}^{\tau}$ with $\|\langle\bm{\gamma}^{\tau},{\bm{\gamma}^{\tau}}^{\top}\rangle_{\mathcal{C}}\|_{2}\leq 1$.
A.4 Proof of Lemma 3.3
(i): Without loss of generality we can assume that $\dim{\mathcal{C}}<\infty$, otherwise we replace ${\mathcal{C}}$ by a finite-dimensional subspace. Define $\nu\coloneq\dim{\mathcal{C}}-1$, and consider an orthonormal basis $\xi_{0}\coloneqq p,\,\xi_{1},\dots,\xi_{\nu}$ of ${\mathcal{C}}$. Then there is a bijection between ${\mathcal{C}}^{m^{\rm sy}}$ and ${\mathbb{R}}^{m^{\rm sy}}\times{\mathbb{R}}^{{m^{\rm sy}}\times\nu}$: every $\bm{\gamma}^{\rm sy}\in{\mathcal{C}}^{m^{\rm sy}}$ can be expressed in unqiue coordinates as $\gamma^{\rm sy}_{i}=b_{i}p+\sum_{j=1}^{\nu}c_{ij}\xi_{j}$ for some vector ${\bm{b}}=[b_{i}:1\leq i\leq m^{\rm sy}]\in{\mathbb{R}}^{m^{\rm sy}}$ and matrix $\bm{C}=[c_{ij}:1\leq i\leq m^{\rm sy},1\leq j\leq\nu]\in{\mathbb{R}}^{m^{\rm sy}\times\nu}$, and vice versa. Expressed in these coordinates, we can write ${\bm{U}}^{\rm sy}(\bm{\gamma}^{\rm sy})=\begin{bmatrix}1&{\bm{b}}^{\top}\\ {\bm{b}}&{\bm{b}}{\bm{b}}^{\top}+\bm{C}\bm{C}^{\top}\end{bmatrix}$. It follows that $\bm{C}\bm{C}^{\top}$ is the Schur complement of the upper left block 1 of the matrix ${\bm{U}}^{\rm sy}(\bm{\gamma}^{\rm sy})$. Hence ${\bm{U}}^{\rm sy}:{\mathcal{C}}^{m^{\rm sy}}\to{\mathcal{D}}^{\rm sy}$ is surjective if and only if every matrix $\bm{\Sigma}\in{\mathbb{S}}^{m^{\rm sy}}_{+}$ can be expressed as $\bm{\Sigma}=\bm{C}\bm{C}^{\top}$ for some $\bm{C}\in{\mathbb{R}}^{m^{\rm sy}\times\nu}$. This holds if and only if $\nu\geq m^{\rm sy}$, see (Paulsen and Raghupathi, 2016, Theorem 4.7), which proves the first statement. For the second statement, let $\bm{A}\neq\bm{I}_{\nu}$ be any orthogonal $\nu\times\nu$-matrix, and define $\tilde{\bm{C}}=\bm{C}\bm{A}$ and $\tilde{\bm{\gamma}}$ accordingly as above. It follows that $\tilde{\bm{\gamma}}\neq\bm{\gamma}$ and $\bm{U}^{\rm sy}(\tilde{\bm{\gamma}})=\bm{U}^{\rm sy}(\bm{\gamma})$. If $\nu\geq 2$ then there exists infinitely many such matrices $\bm{A}$, which proves the claim.
(ii): This follows similarly as part (i), but without constraint on $\bm{U}^{\rm id}_{11}$.
A.5 Proof of Lemma 4.1
The bound (16) follows from the Rayleigh–Ritz Theorem (Horn and Johnson, 1990, Theorem 4.2.2) and because $\|\bm{D}_{n}\bm{v}\|_{2}\geq\|\bm{v}\|_{2}$. The second statement follows from elementary matrix algebra.
A.6 Proof of Lemma 4.2
We use the elementary facts $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{F}\|\bm{B}\|_{F}$ and $\|\bm{A}\otimes\bm{B}\|_{F}=\|\bm{A}\|_{F}\|\bm{B}\|_{F}$ for matrices $\bm{A}$ and $\bm{B}$, for the Frobenius norm $\|\cdot\|_{F}$. By construction, it follows that $\|\bm{y}(\bm{x})\|_{2}\leq 1+\|\bm{x}\|_{2}^{2}$, $\|\bm{\Psi}^{\rm sy}(\bm{z})\|_{F}^{2}=1+\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{2}$, $\|\bm{\Psi}^{\rm id}(\bm{z})\|_{F}=\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}$, $\|\bm{D}_{n}\|_{F}=n$, $\|\bm{R}_{n}\bm{B}\|_{F}=\|\bm{B}\|_{F}$, $\|\bm{R}_{n}^{\top}\bm{B}\|_{F}\leq\|\bm{B}\|_{F}$, for any conformal matrix $\bm{B}$. Hence
$\|\bm{A}(\xi)\|_{F}$ $\leq 2(N+1)^{-2}\|\bm{Q}(\xi)\|_{F}^{2}$ (31)
$\leq 2(N+1)^{-2}\big{(}(1+\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{2})^{2}(m^{\rm sy}+1)^{2}+\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}^{4}(m^{\rm id})^{2}\big{)},$
$\|\bm{b}(\xi)\|_{2}$ $\leq 2(N+1)^{-2}\big{(}(1+\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{2})(m^{\rm sy}+1)+\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}^{2}m^{\rm id}\big{)}$
$\quad+\lambda^{\rm sy}(m^{\rm sy}+1)\|\bm{g}^{\rm sy}\|_{2}+\lambda^{\rm id}m^{\rm id}\|\bm{g}^{\rm id}\|_{2},$
$|c(\xi)|$ $\leq(N+1)^{-2}\big{(}1+\|\bm{x}\|_{2}^{2}\big{)}^{2}.$
Combining (18) and (31) proves the lemma.
A.7 Proof of Theorem 4.3
Clearly, ${\mathcal{R}}(\bm{u},\xi)$ is a Carathéodory function, i.e., measurable in $\xi$ and continuous in $\bm{u}$, and therefore random lower semicontinuous (Shapiro et al., 2021, Section 9.2.4). The set ${\mathcal{U}}$ is closed and convex in ${\mathbb{R}}^{M}$. Now claim (i) follows from (Shapiro et al., 2021, Theorem 5.4).
Claims (ii) and (iii) follow from (Milz, 2023, Theorem 3), setting “$\bm{\Psi}(\bm{u})$” in Milz (2023) equal to the convex characteristic function of the feasible set ${\mathcal{U}}$ in ${\mathbb{R}}^{M}$, taking value 00 for $\bm{u}\in{\mathcal{U}}$ and +$\infty$+ ∞ otherwise.
Claim (iv) follows as (13) elementary implies the bound (21), using that the operator norm of the half-vectorization operator is given by $\sup_{\|\bm{u}\|_{2}\leq 1}\|\operatorname{vech}(\bm{u})\|_{F}=\sqrt{2}$.
Claim (v) follows from Jensen’s inequality and the bounds in (31).
Claim (vi) follows as the above proof applies to any closed convex subset of ${\mathcal{U}}$.
A.8 Proof of Lemma 4.4
$\bm{U}_{\operatorname{diag}}^{{\rm sy}}$ is clearly symmetric. Furthermore, all (non-leading) principal minors are diagonal matrices with entries along the diagonal that are combinations of $c_{1},\ldots,c_{m^{{\rm sy}}}\geq 0$ from the premises of the statements, and thus positive semidefinite. The top-left corner is equal to one, and therefore positive. To consider the remaining $l=1,\ldots,m^{{\rm sy}}$ leading principal minors, we apply the block determinant formula to obtain for the determinant of the $l$-th leading principal minor,
$\left(1-\sum_{j=1}^{l}{\frac{b_{j}^{2}}{c_{j}}}\right)\prod_{i=1}^{l}c_{i}\geq(1-\sum_{j=1}^{l}\tilde{c}_{j})\prod_{i=1}^{l}c_{i}\geq 0,$
from the premise of the statement. With all principal minors positive semidefinite, Sylvester’s criterion (Horn and Johnson, 1990, Theorem 7.2.5) applies, and yields that $\bm{U}^{{\rm sy}}_{\operatorname{diag}}$ is symmetric positive semidefinite. Conversely, the matrix $\bm{U}^{{\rm sy}}_{\operatorname{diag}}$ being positive semidefinite implies the leading principal minors to be non-negative, such that in turn $c_{1},\ldots,c_{m^{{\rm sy}}}\geq 0$, $\tilde{c}_{1},\ldots,\tilde{c}_{m^{{\rm sy}}}\geq 0$, $\sum_{i=1}^{m^{{\rm sy}}}\tilde{c}_{i}\leq 1$ and $b_{i}^{2}\leq c_{i}\tilde{c}_{i}$ for $l=1,\ldots,m^{{\rm sy}}$ (Horn and Johnson, 1990, Corollary 7.1.5).
From the block-diagonal specification clearly ${\mathcal{D}}^{{\rm sy}}_{\operatorname{diag}}\subset{\mathcal{D}}^{{\rm sy}}$. Finally, all constraints in the premise of the statement describe closed convex sets (the constraints $b_{i}^{2}\leq c_{i}\tilde{c}_{i},\,i=1,\ldots,m^{{\rm sy}}$ are commonly referred to as rotated quadratic cones, and jointly convex in $b_{i},c_{i}$ and $\tilde{c}_{i}$) and their intersection thus describes a closed convex set.