### v5

Abstract
We develop a nonparametric, kernel-based joint estimator for conditional mean and covariance matrices in large and unbalanced panels. The estimator is supported by rigorous consistency results and finite-sample guarantees, ensuring its reliability for empirical applications. We apply it to an extensive panel of monthly US stock excess returns from 1962 to 2021, using macroeconomic and firm-specific covariates as conditioning variables. The estimator effectively captures time-varying cross-sectional dependencies, demonstrating robust statistical and economic performance. We find that idiosyncratic risk explains, on average, more than 75% of the cross-sectional variance.
Keywords: nonparametric estimation, conditional mean, conditional covariance matrix, unbalanced panels, mean-variance efficient portfolio
JEL classification: C14, C58, G11
1 Introduction
The relationship between conditional expected returns, conditional risk, and both asset-level and macroeconomic covariates has been a central topic in financial economics for decades. Yet, inference in this domain remains constrained by the unbalanced and high-dimensional nature of real-world data. In this paper, we address these challenges by introducing a nonparametric, kernel-based framework for the joint estimation of conditional mean and covariance matrices, providing a powerful and tractable solution to the econometric inference problem highlighted by Cochrane (2011). Our framework is specifically designed to deliver positive semidefinite covariance matrices across any state and for cross sections of varying sizes, filling a significant gap in the literature. [^4] Many papers in finance build on conditional means and covariances, which are typically assumed to be exogenously given. For example, Goyenko et al. (2024) take the first and second moments of security returns as given when constructing mean–variance efficient portfolios net of trading costs.
Since Fama and MacBeth (1973), empirical researchers studying unbalanced panels have primarily relied on tools such as portfolio sorts (Fama and French, 1993, 2019; Kozak et al., 2020), models for expected returns applicable to both balanced and unbalanced panels (Connor et al., 2012; Fan et al., 2016; Freyberger et al., 2020; Gu et al., 2020b; Kelly et al., 2019; Kozak and Nagel, 2024), and econometric inference methods for linear factor models (Zaffaroni, 2019; Fortin et al., 2023a, b). While recent econometric literature has introduced conditional covariance estimators tailored for high-dimensional settings (Fan et al., 2013; Fan and and, 2018; Engle et al., 2019; Fan et al., 2019), there remains a critical gap: a scalable, nonparametric framework capable of jointly and consistently estimating conditional means and covariances in a large-scale, unbalanced context. [^5] We are not the first to model conditional means and covariances for asset returns. The approach most closely related to ours is Gao (2011), who estimates these quantities sequentially using a local nonparametric smoothing method. Kirby (2018) proposes a model based on a parametric GARCH volatility specification. In a different direction, Clarke and Linn (2024) represent covariances as a superposition of indicator functions and likewise estimate means and covariances sequentially. None of these studies provide finite-sample guarantees for their estimators, as we do in this paper. Moreover, our framework is formulated in a general conditional modeling setting and applies beyond financial data.
We address this problem by proposing a novel, nonparametric, kernel-based model for jointly estimating conditional first and second moments for unbalanced panels of arbitrary size, requiring only that these conditional moments can be represented within a large and flexible hypothesis space. Our model uniquely ensures that, at any point in time and across any cross-sectional dimension, conditional return covariances incorporate both systematic and idiosyncratic components, and remain symmetric and positive semidefinite, despite their nonparametric construction. We refer to this model as the joint conditional mean and covariance (COCO) estimator. Our approach tackles a significant limitation in the literature, which typically focuses on either covariances or first moments independently. Moreover, our model’s functional form is optimal with respect to the mean squared loss and, while it is broadly applicable, aligns precisely with the characterization of economies that can be spanned by factor portfolios, as discussed in Kozak and Nagel (2024).
The COCO estimator is computationally efficient and scalable, capable of handling large datasets on standard desktop computing hardware. Its parsimonious structure means that finite-dimensional specifications may not even require validation, enhancing practical applicability. Furthermore, the model provides a natural low-rank representation with controlled approximation error, leading to a Chamberlain and Rothschild (1983)-type conditional factor structure, where the rank of the conditional covariance matrix corresponds to the number of systematic factors. Crucially, the estimator emerges from a convex optimization problem, ensuring reproducibility—a distinct advantage over non-convex models prevalent in deep learning and other econometric frameworks.
We empirically test the COCO estimator on an extensive unbalanced panel of monthly US stock returns from 1962 to 2021, incorporating both asset-level and macroeconomic covariates. Our results indicate that the COCO estimator offers moderate predictability for realized excess returns, with stronger and more reliable predictability for their squares and mixed products, which correspond to conditional second moments, especially in the early sample period. By jointly assessing both moments, the COCO estimator significantly outperforms a baseline model that accounts only for idiosyncratic risk. On the other hand, we find that idiosyncratic risk explains, on average, more than 75% of the cross-sectional variance. The conditional mean-variance efficient (cMVE) portfolio constructed from the COCO estimates achieves substantial annualized out-of-sample Sharpe ratios, markedly outperforming equal-weighted portfolios over the entire sample period. Furthermore, cMVE returns exhibit weak correlations with the Fama–-French five factors (Fama and French, 2015). As the number of systematic factors in our model increases, the connection to the Fama–French factors diminishes, ultimately rendering the variation in cMVE portfolio returns largely unrelated to the traditional five-factor model. The empirical findings are complemented by a simulation study that further supports the robustness and reliability of our method.
The remainder of this paper is structured as follows. Section 2 introduces the nonparametric model for conditional moments and establishes its connection to data-generating linear factor models. Section 3 defines the joint estimator, deriving a representation theorem for the optimal conditional moment function and a corresponding low-rank approximation. Section 4 establishes consistency and finite-sample guarantees for the estimator. Section 5 presents a large-scale empirical analysis using a panel of US stock returns, highlighting the estimator’s statistical performance and its implications for asset pricing. Section 6 concludes. The appendix contains additional theoretical results, all proofs, and a simulation study.
2 Conditional mean and covariance model
We begin by introducing the econometric framework and notation used throughout the paper. We consider discrete time periods $(t,t+1]$, $t=0,1,2,\dots$, e.g., months. For each period $(t,t+1]$, there are $N_{t}$ assets $i=1,\dots,N_{t}$ with observable covariates $z_{t,i}$ at time $t$. These covariates take values in a fixed covariate space ${\mathcal{Z}}$, common across all periods. The assets yield returns $x_{t+1,i}$ over $(t,t+1]$. We remain agnostic about the type of “return” that could be gross, simple, logarithmic, excess, or forward gross. More generally, our estimator is applicable to a wide range of real-valued variables $x_{t+1,i}$, beyond returns, provided they are accompanied by observed covariates $z_{t,i}$. In the empirical study, we will work with simple excess returns, as is customary in the literature and convenient for asset pricing.
Our goal is to define a model for conditional first and second moments, ${\mathbb{E}}_{t}[x_{t+1,i}]$ and ${\mathbb{E}}_{t}[x_{t+1,i}x_{t+1,j}]$, of the returns, given the information set at time $t$. To this end, we assume that these conditional moments are given by functions $\mu:{\mathcal{Z}}\to{\mathbb{R}}$ and $q:{\mathcal{Z}}\times{\mathcal{Z}}\to{\mathbb{R}}$ of the respective covariates such that
${\mathbb{E}}_{t}[x_{t+1,i}]$ $=\mu(z_{t,i}),$
${\mathbb{E}}_{t}[x_{t+1,i}x_{t+1,j}]$ $=q(z_{t,i},z_{t,j}),$
which implies that the conditional covariance is given by
$\operatorname{Cov}_{t}[x_{t+1,i},x_{t+1,j}]=q(z_{t,i},z_{t,j})-\mu(z_{t,i})\mu(z_{t,j}).$
For notational convenience and to facilitate the matrix-based representations used in this paper, we stack the asset-level data into arrays: $\bm{x}_{t+1}\coloneqq[x_{t+1,i}:1\leq i\leq N_{t}]\in\mathbb{R}^{N_{t}}$ denotes the vector of asset returns, and $\bm{z}_{t}\coloneqq[z_{t,i}:1\leq i\leq N_{t}]\in{\mathcal{Z}}^{N_{t}}$ denotes the corresponding array of covariates. In a similar vein, we write $\mu(\bm{z}_{t})\coloneqq[\mu(z_{t,i}):1\leq i\leq N_{t}]$ and $q(\bm{z}_{t},\bm{z}_{t}^{\top})\coloneqq[q(z_{t,i},z_{t,j}):1\leq i,j\leq N_{t}]$ for the corresponding arrays of function values.
The central modeling challenge is to specify functions $\mu$ and $q$ such that, for any $t$, the $N_{t}\times N_{t}$-matrix
$q(\bm{z}_{t},\bm{z}_{t}^{\top})-\mu(\bm{z}_{t})\mu(\bm{z}_{t})^{\top}$ is symmetric and positive semidefinite. (1)
Meeting this condition ensures that the matrix qualifies as a valid and consistent model for a conditional covariance matrix. Since the rank-one matrix $\mu(\bm{z}_{t})\mu(\bm{z}_{t})^{\top}$ is positive semidefinite, Condition (1) implies that $q(\bm{z}_{t},\bm{z}_{t}^{\top})$ is also symmetric and positive semidefinite. This is precisely the defining property of a real-valued positive-type, or kernel function, see, e.g., Paulsen and Raghupathi (2016, Section 2.2). We thus assume that $q$ is a kernel function on ${\mathcal{Z}}\times{\mathcal{Z}}$. To assert condition (1), we extend the covariate space ${\mathcal{Z}}_{\Delta}\coloneqq{\mathcal{Z}}\cup\{\Delta\}$, for some auxiliary point ${\Delta}\notin{\mathcal{Z}}$. We then extend $q$ to be a kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ such that
$q({\Delta},{\Delta})=1,$ (2)
and set $\mu(z)\coloneqq q(z,{\Delta})$. This implies that the implied covariance function $c(z,z^{\prime})\coloneq q(z,z^{\prime})-\mu(z)\mu(z^{\prime})=q(z,z^{\prime})-q(z,{\Delta})q(z^{\prime},{\Delta})$ is the Schur complement of $q$ with respect to ${\Delta}$. It is therefore itself a kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ (see Paulsen and Raghupathi, 2016, Theorem 4.5), and thus (1) holds.
Our goal, therefore, boils down to specifying an appropriate kernel function $q$ on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ that satisfies (2). To achieve this, we introduce a novel nonparametric approach for directly learning such a kernel function, grounded in principles of finance and specifically tailored to fit the data optimally.
Specifically, we adopt the common assumption that the conditional covariance can be decomposed into a systematic and an idiosyncratic component. The former captures the conditional dependence between returns, and their risk premiums, explained by common underlying risk factors. The latter captures the conditional uncorrelated individual return risks, which asymptotically have a conditional mean of zero under the absence of arbitrage in large cross sections (see Ross, 1976; Chamberlain, 1983; Chamberlain and Rothschild, 1983; Reisman, 1988). We take this into account and decompose $q(z,z^{\prime})=q^{\rm sy}(z,z^{\prime})+q^{\rm id}(z,z^{\prime})$ into the sum of two corresponding kernel functions, where the idiosyncratic component $q^{\rm id}(z,z^{\prime})=q^{\rm id}(z,z^{\prime})1_{z=z^{\prime}}$ is supported on the diagonal of the product space ${\mathcal{Z}}\times{\mathcal{Z}}$. Accordingly, we assume that $q^{\rm sy}(\Delta,\Delta)=1$ such that the systematic component captures the structural condition (2).
We denote by ${\mathcal{C}}$ an auxiliary separable Hilbert space and select an arbitrary unit vector $p\in{\mathcal{C}}$, so that $\langle p,p\rangle_{\mathcal{C}}=1$. For concreteness, we assume $\mathcal{C}$ to be $\ell^{2}$, the space of square-summable sequences, which is standard in this context; however, other choices are possible. [^6] For example, one may take ${\mathcal{C}}=L^{2}(\Omega,{\mathcal{F}},{\mathbb{M}})$, the space of square-integrable random variables on an auxiliary probability space $(\Omega,{\mathcal{F}},{\mathbb{M}})$. A natural choice for the unit vector in this case is the constant function $p=1$. For the minimal dimensional requirements of ${\mathcal{C}}$, see Lemma B.1 in the appendix. For any pair of feature maps $h=(h^{{\rm sy}},h^{{\rm id}})$, where $h^{\tau}:{\mathcal{Z}}\to{\mathcal{C}}$, we extend these maps to ${\mathcal{Z}}_{\Delta}$ by defining their values at $\Delta$ to be the zero element in ${\mathcal{C}}$,
$h^{\tau}(\Delta)\coloneqq 0,\quad\text{for }\tau\in\{{\rm sy},{\rm id}\}.$ (3)
This extension enables the definition of a moment kernel function on ${\mathcal{Z}}_{\Delta}\times{\mathcal{Z}}_{\Delta}$ as follows: [^7] This construction leverages the fact that inner products are kernel functions, and that sums, and products of kernel functions are also valid kernel functions, see Paulsen and Raghupathi (2016, Section 2.3.4 and Chapter 5).
$q_{h}(z,z^{\prime})\coloneqq\underbrace{\langle h^{\rm sy}(z)+p1_{z=\Delta},h^{\rm sy}(z^{\prime})+p1_{z^{\prime}=\Delta}\rangle_{\mathcal{C}}}_{\text{systematic component }q^{{\rm sy}}_{h}(z,z^{\prime})}+\underbrace{\|h^{\rm id}(z)\|_{\mathcal{C}}^{2}\,1_{z=z^{\prime}}}_{\text{idiosyncratic component }q^{{\rm id}}_{h}(z,z^{\prime})}.$ (4)
From (3), it immediately follows that (2) is satisfied. This implies the conditional mean and covariance functions read
$\mu_{h}(z)$ $=\langle{h^{\rm sy}}(z),p\rangle_{\mathcal{C}},$ (5)
[…]
We henceforth assume that $z_{t,i}=z_{t,j}$ if and only if $i=j$, for each cross section $t$. This assumption is made without loss of generality, as otherwise we could simply assume that the index $i$ is part of the covariates $z_{t,i}$. Consequently, this ensures a diagonal idiosyncratic matrix component in the expressions presented below.
We now demonstrate that our framework (4) for the moment kernel function is universal in the sense that it encompasses all data-generating conditional factor models of the form
$x_{t+1,i}=\underbrace{\alpha(z_{t,i})}_{\text{intercept}}+\underbrace{\langle\beta(z_{t,i}),g_{t+1}\rangle_{\mathcal{C}}}_{\text{systematic risk}}+\underbrace{\gamma(z_{t,i})w_{t+1}(z_{t,i})}_{\text{idiosyncratic risk}},$ (6)
where $\alpha:{\mathcal{Z}}\to{\mathbb{R}}$ is a conditional intercept function, $\beta:{\mathcal{Z}}\to{\mathcal{C}}$ a factor loadings map, and $\gamma:{\mathcal{Z}}\to[0,\infty)$ an idiosyncratic volatility function. The term $g_{t+1}$ is a ${\mathcal{C}}$-valued stationary risk factor process with constant conditional mean $b\coloneqq{\mathbb{E}}_{t}[g_{t+1}]$ and covariance operator $Q\coloneqq\operatorname{Cov}_{t}[g_{t+1}]$. We assume that $g_{t+1}$ and $\beta(z)$ take values in a subspace of ${\mathcal{C}}$ of codimension 1.888This assumption is without loss of generality, as we show in the proof of Theorem 2.1. The collection of real-valued random variables $\{w_{t+1}(z):z\in{\mathcal{Z}}\}$ is a white noise process, such that ${\mathbb{E}}_{t}[w_{t+1}(z)]=0$ and ${\mathbb{E}}_{t}[w_{t+1}(z)w_{t+1}(z^{\prime})]=1_{z=z^{\prime}}$. It is conditionally uncorrelated with $g_{t+1}$, in the sense that $\operatorname{Cov}_{t}[g_{t+1},w_{t+1}(z)]={\mathbb{E}}_{t}[g_{t+1}w_{t+1}(z)]=0\in{\mathcal{C}}$ for all $z\in{\mathcal{Z}}$.
The following theorem formalizes our claim. The third part gives a representation of the data-generating conditional factor model (6) in terms of factors that are linear in $\bm{x}_{t+1}$ and therefore observable, in contrast to $g_{t+1}$, which may be latent. These observable factors can be interpreted as portfolio returns, with significant implications for asset pricing, as examined in detail in Filipović and Schneider (2024). Here and below, we write $A^{+}$ to denote the Moore–Penrose pseudoinverse of a bounded linear operator $A:{\mathcal{C}}\to\mathbb{R}^{N_{t}}$, defined pointwise by $A^{+}\bm{v}\coloneqq\lim_{\lambda\downarrow 0}(A^{*}A+\lambda I_{\mathcal{C}})^{-1}A^{*}\bm{v}$, where $A^{*}$ denotes the adjoint of $A$.
Theorem 2.1.
The proposed framework is universal in the following sense:
  1. (i)
    Every data-generating conditional factor model (6) has conditional mean and covariance functions of the form (5).
  2. (ii)
    Conversely, for every moment kernel function (4) there exists a data-generating conditional factor model of the form (6) with conditional mean and covariance functions given by (5).
  3. (iii)
    If $\alpha(z)=0$, the data-generating conditional factor model (6) can be represented as
    $x_{t+1,i}=\langle\beta(z_{t,i}),f_{t+1}\rangle_{\mathcal{C}}+\epsilon_{t+1,i}$ (7)
    in terms of the linear ${\mathcal{C}}$-valued factors $f_{t+1}\coloneqq(\bm{S}_{t}\beta(\bm{z}_{t}))^{+}\bm{S}_{t}\bm{x}_{t+1}$, where $\bm{S}_{t}$ is the $N_{t}\times N_{t}$-diagonal matrix with diagonal elements $S_{t,ii}\coloneqq\gamma(z_{t,i})^{-1}$ if $\gamma(z_{t,i})>0$ and $S_{t,ii}\coloneqq 1$ otherwise. The residuals given by $\epsilon_{t+1,i}\coloneqq x_{t+1,i}-\langle\beta(z_{t,i}),f_{t+1}\rangle_{\mathcal{C}}$ have zero conditional mean ${\mathbb{E}}_{t}[\epsilon_{t+1,i}]=0$ and are conditionally uncorrelated with $f_{t+1}$.
3 Joint estimation
To estimate $h=(h^{\rm sy},h^{\rm id})$, we leverage the law of iterated expectations for conditional moments and cast the estimation problem as a matrix-valued regression,
$\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}=\begin{bmatrix}1&\langle p,h^{\rm sy}(\bm{z}_{t})\rangle_{\mathcal{C}}\\ \langle h^{\rm sy}(\bm{z}_{t}),p\rangle_{\mathcal{C}}&\langle h^{\rm sy}(\bm{z}_{t}),h^{\rm sy}(\bm{z}_{t})\rangle_{\mathcal{C}}\end{bmatrix}+\begin{bmatrix}0&\bm{0}\\ \bm{0}&\operatorname{diag}(\|h^{\rm id}(\bm{z}_{t})\|_{\mathcal{C}}^{2})\end{bmatrix}+\bm{E}_{t+1},$
where $\bm{E}_{t+1}$ denotes a matrix of errors satisfying ${\mathbb{E}}_{t}[\bm{E}_{t+1}]=\bm{0}$. For notational convenience, we define a data point as $\xi_{t}\coloneq(N_{t},\bm{x}_{t+1},\bm{z}_{t})$, which summarizes the relevant information from the cross section. We also introduce a weight function $w(N_{t})\coloneqq 1/N_{t}$, which accounts for variation in cross-sectional sample sizes $N_{t}$. [^9] We can easily generalize the weighting in the loss function (8) by any exogenous weights $\nu_{t,i}\in(0,1)$, $0\leq i\leq N_{t}$, such that $\sum_{i}\nu_{t,i}=1$ and set ${\mathcal{L}}(h,\xi_{t})=w(N_{t})\sum_{0\leq i,j\leq N_{t}}\nu_{t,i}\nu_{t,j}(x_{t+1,i}x_{t+1,j}-q_{h}(z_{t,i},z_{t,j}))^{2}.$ This is captured by (8) simply by replacing the data $x_{t,i}$ by $\nu_{t,i}^{1/2}x_{t,i}$ and $q_{h}(z_{t,i},z_{t,j})$ by $\nu_{t,i}^{1/2}q_{h}(z_{t,i},z_{t,j})\nu_{t,j}^{1/2}$. For example, choosing $\nu_{t,0}\in(0,1)$ and setting $\nu_{t,i}=(1-\nu_{t,0})/N_{t}$ for all $i\geq 1$, allows to balance the weights given to the first and second moment error terms in (8). Alternative choices for $w(N_{t})$ are also possible. Our choice of $w(N_{t})=1/N_{t}$ is motivated by the scaled Frobenius norm used in (Ledoit and Wolf, 2004, Definition 1); see also (Ledoit and Wolf, 2020, Equation (1.1)). In addition, (Bodnar et al., 2014, Theorem 3.1) provides evidence that the squared Frobenius norm of the sample covariance matrix scales linearly with the dimension $p$, provided the sample size $n$ grows proportionally with $p$. Importantly, all theoretical results and analysis in this paper are derived in terms of a general weight function $w(N_{t})$, and our findings do not depend on the specific choice $w(N_{t})=1/N_{t}$. This yields the following weighted squared loss function, which reflects the regression structure implied by the conditional moments:
${\mathcal{L}}(h,\xi_{t})$ :-$\coloneq w(N_{t})\left\|\bm{E}_{t+1}\right\|_{F}^{2}:$ (8)
$=2\underbrace{w(N_{t})\left\|\bm{x}_{t+1}-\langle h^{\rm sy}(\bm{z}_{t}),p\rangle_{\mathcal{C}}\right\|_{2}^{2}}_{\text{first moment error}}$
$\quad+\underbrace{w(N_{t})\left\|\bm{x}_{t+1}\bm{x}_{t+1}^{\top}-\langle h^{\rm sy}(\bm{z}_{t}),h^{\rm sy}(\bm{z}_{t})^{\top}\rangle_{\mathcal{C}}-\operatorname{diag}(\|h^{\rm id}(\bm{z}_{t})\|_{\mathcal{C}}^{2})\right\|_{F}^{2}}_{\text{second moment error}},$
where $\|\cdot\|_{F}$ and $\|\cdot\|_{2}$ denote the Frobenius and Euclidean norm, respectively.
The flexibility and empirical success of our approach crucially depends on the specification of the feature map $h=(h^{\rm sy},h^{\rm id})$ as an element in a potentially infinite-dimensional hypothesis space ${\mathcal{H}}$. Specifically, we assume that ${\mathcal{H}}={\mathcal{H}}^{\rm sy}\times{\mathcal{H}}^{\rm id}$ is the product space of separable ${\mathcal{C}}$-valued reproducing kernel Hilbert spaces (RKHS) ${\mathcal{H}}^{\rm sy}$, ${\mathcal{H}}^{\rm id}$, consisting of functions $h^{\rm sy},h^{\rm id}:{\mathcal{Z}}\to{\mathcal{C}}$, and with operator-valued reproducing kernels $K^{\rm sy},K^{\rm id}$ on ${\mathcal{Z}}$. We refer the reader to Paulsen and Raghupathi (2016, Chapter 6) for the definition and basic properties of these RKHSs. For tractability we further assume that the kernels are separable, $K^{\rm sy}(z,z^{\prime})=k^{\rm sy}(z,z^{\prime})I_{\mathcal{C}}$, $K^{\rm id}(z,z^{\prime})=k^{\rm id}(z,z^{\prime})I_{\mathcal{C}}$, for some given scalar reproducing kernels $k^{\rm sy},k^{\rm id}$ of separable RKHS ${\mathcal{G}}^{\rm sy},{\mathcal{G}}^{\rm id}$ on ${\mathcal{Z}}$, so that ${\mathcal{H}}^{\rm sy}\cong{\mathcal{G}}^{\rm sy}\otimes{\mathcal{C}}$, ${\mathcal{H}}^{\rm id}\cong{\mathcal{G}}^{\rm id}\otimes{\mathcal{C}}$ can be identified with tensor product spaces. To control model complexity and mitigate overfitting, we add penalty terms with regularization parameters $\lambda^{\rm sy},\lambda^{\rm id}>0$ to the objective (8), resulting in the regularized loss function,
${\mathcal{R}}(h,\xi_{t})\coloneqq{\mathcal{L}}(h,\xi_{t})+\underbrace{\lambda^{\rm sy}\|h^{\rm sy}\|_{{\mathcal{H}}^{\rm sy}}^{2}+\lambda^{\rm id}\|h^{\rm id}\|_{{\mathcal{H}}^{\rm id}}^{2}}_{\text{regularization}}.$ (9)
Finally, taking the sample average, we arrive at the non-standard kernel ridge regression problem,
minimize $\underset{h\in{\mathcal{H}}}{\text{minimize }}\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{R}}(h,\xi_{t}).$ (10)
Notably, problem (10) is not convex in $h\in{\mathcal{H}}$, due to the inner product appearing in the loss function (8). [^10] In fact, for any given $z,z^{\prime}\in{\mathcal{Z}}$, the function $Q:{\mathcal{H}}^{\rm sy}\to{\mathbb{R}}$, $h^{\rm sy}\mapsto Q(h^{\rm sy})=\langle h^{\rm sy}(z),h^{\rm sy}(z^{\prime})\rangle_{\mathcal{C}}$ is neither convex nor concave in $h^{\rm sy}$ in general. We see this by means of the following example. Let $h^{\rm sy}_{1},h^{\rm sy}_{2}\in{\mathcal{H}}^{\rm sy}$ such that $h^{\rm sy}_{1}(z)=0$ and $h^{\rm sy}_{2}(z^{\prime})=0$. Then $Q(h^{{\rm sy}}_{1})=Q(h^{\rm sy}_{2})=0$. On the other hand, for any $s\in(0,1)$, $Q(sh^{\rm sy}_{1}+(1-s)h^{\rm sy}_{2})=(1-s)s\big{\langle}h^{\rm sy}_{2}(z),h^{\rm sy}_{1}(z^{\prime})\big{\rangle}_{\mathcal{C}}$, which could be either positive or negative. It can therefore neither be bounded below nor above by $sQ(h^{\rm sy}_{1})+(1-s)Q(h^{\rm sy}_{2})=0$. It follows that, in general, there are infinitely many solutions $h$ of (10), although they all imply the same optimal moment kernel function $q_{h}$. [^11] This follows from Lemma B.1 in the appendix.
As a first step towards solving (10), we establish a representer theorem for this non-standard problem, which generalizes Micchelli and Pontil (2005, Theorem 4.1). For further use, we denote the total sample size by $N_{\rm tot}\coloneqq\sum_{t=0}^{T-1}N_{t}$.
Theorem 3.1 (Representer Theorem).
Any minimizer $h=(h^{\rm sy},h^{\rm id})$ of (10) is of the form
$h^{\tau}(\cdot)=\sum_{t=0}^{T-1}\sum_{i=1}^{N_{t}}k^{\tau}(\cdot,z_{t,i})\gamma^{\tau}_{t,i},\quad\text{for coefficients $\gamma^{\tau}_{t,i}\in{\mathcal{C}}$,}$ (11)
for both components $\tau\in\{{\rm sy},{\rm id}\}$.
Inserting the optimal functional form (11), problem (10) can be equivalently expressed in terms of $N_{\rm tot}$ pairs of coefficients $(\gamma^{{\rm sy}}_{t,i},\gamma^{{\rm id}}_{t,i})\in{\mathcal{C}}\times{\mathcal{C}}$. Although the optimal form (11) grants a considerable simplification of the full infinite-dimensional problem, it is generally still computationally infeasible for large $N_{\rm tot}$. In the following we therefore propose a low-rank approximation, along with a reparametrization, of problem (10). This will result in a low-dimensional convex optimization problem, which approximates the original problem.
To this end, we consider the Nyström method (Drineas and Mahoney, 2005), and denote by $\bm{Z}\coloneq[\bm{z}_{t}:0\leq t\leq T-1]\in{\mathcal{Z}}^{N_{\rm tot}}$ the full sample array of covariates. For each component $\tau\in\{{\rm sy},{\rm id}\}$, we consider a subsample $\Pi^{\tau}\subset\{1,\dots,N_{\rm tot}\}$ of size $m^{\tau}\leq N_{\rm tot}$ that approximates the full kernel matrix such that the trace error
$\epsilon^{\tau}_{\rm approx}\coloneqq\operatorname{tr}\left(k^{\tau}(\bm{Z},\bm{Z}^{\top})-k^{\tau}(\bm{Z},\bm{Z}_{\Pi^{\tau}}^{\top})k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}_{\Pi^{\tau}}^{\top})^{-1}k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}^{\top})\right)$ (12)
is small. This subsample selection is facilitated by a pivoted Cholesky decomposition (see Harbrecht et al., 2012; Chen et al., 2023). It yields $m^{\tau}$ linearly independent functions $\phi^{\tau}_{i}(\cdot)$ in ${\mathcal{G}}^{\tau}$, forming an ${\mathbb{R}}^{m^{\tau}}$-valued feature map defined as $\bm{\phi}^{\tau}(\cdot)\coloneqq[\phi^{\tau}_{1}(\cdot),\dots,\phi^{\tau}_{m^{\tau}}(\cdot)]\coloneq k^{\tau}(\cdot,\bm{Z}_{\Pi^{\tau}}^{\top})\bm{B}^{\tau}$, where $\bm{B}^{\tau}$ is an arbitrarily chosen invertible square matrix. [^12] The functions $\bm{\phi}^{\tau}$ are orthonormal in ${\mathcal{G}}^{\tau}$ if and only if $\bm{B}^{\tau}{\bm{B}^{\tau}}^{\top}=k^{\tau}(\bm{Z}_{\Pi^{\tau}},\bm{Z}_{\Pi^{\tau}}^{\top})^{-1}$. However, this assumption is not imposed here, allowing for the use of flexible, user-defined feature maps and thereby enhancing the modularity of our framework. We restrict problem (10) to the subspace ${\mathcal{H}}_{0}={\mathcal{H}}^{\rm sy}_{0}\times{\mathcal{H}}^{\rm id}_{0}$ of ${\mathcal{H}}$ consisting of functions of the form
$h^{\tau}_{0}(\cdot)=\sum_{i=1}^{m^{\tau}}\phi^{\tau}_{i}(\cdot)\gamma^{\tau}_{i},\quad\text{for coefficients $\gamma^{\tau}_{i}\in{\mathcal{C}}$,}$ (13)
for both components $\tau\in\{{\rm sy},{\rm id}\}$. The following proposition provides a heuristic for assessing the quality of this low-rank approximation. [^13] However, note that the optimizer of problem (10) restricted to $h_{0}\in{\mathcal{H}}_{0}={\mathcal{H}}^{\rm sy}_{0}\times{\mathcal{H}}^{\rm id}_{0}$ is generally not given as orthogonal projection on ${\mathcal{H}}_{0}$ of any optimizer of the unrestricted problem.
Proposition 3.2.
Let $h^{\tau}\in{\mathcal{H}}^{\tau}$ be an arbitrary candidate function of the form (11), and denote by $h^{\tau}_{0}$ its projection on ${\mathcal{H}}^{\tau}_{0}$, which is given by the expression on the right hand side of (11) with the kernel function $k^{\tau}(z,z^{\prime})$ replaced by its projection $k^{\tau}_{0}(z,z^{\prime})=\bm{\phi}^{\tau}(z)\langle{\bm{\phi}^{\tau}}^{\top},\bm{\phi}^{\tau}\rangle_{{\mathcal{G}}^{\tau}}^{-1}\bm{\phi}^{\tau}(z^{\prime})^{\top}$. Then the difference $q_{h}(z,z^{\prime})-q_{h_{0}}(z,z^{\prime})$ is a kernel function, and the aggregated cross-sectional approximation error of the implied conditional moment matrices is bounded by
$\sum_{t=0}^{T-1}\big{\|}q_{h}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})-q_{h_{0}}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})\big{\|}_{F}\leq\sum_{\tau\in\{{\rm sy},{\rm id}\}}\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}^{2}\epsilon_{\rm approx}^{\tau},$ (14)
where we denote the extended covariate array $\bar{\bm{z}_{t}}^{\top}\coloneqq\begin{bmatrix}\Delta&\bm{z}_{t}^{\top}\end{bmatrix}\in\{\Delta\}\times{\mathcal{Z}}^{N_{t}}$.
The following theorem provides a reparametrization of the moment kernel and regularized loss function when restricted to feature maps in the subspace ${\mathcal{H}}_{0}$. This reparametrization reduces the estimation problem (10) to a convex optimization over the convex feasible set of pairs of matrices $\bm{U}=(\bm{U}^{\rm sy},\bm{U}^{\rm id})$, defined as
[…]
Existence and uniqueness of this convex problem are established in Section 4. We denote by $\operatorname{Diag}(\bm{A})\coloneqq\operatorname{diag}(\operatorname{diag}(\bm{A}))$ the matrix-to-diagonal matrix operator, which extracts the diagonal of a square matrix $\bm{A}$ and converts that vector to a conformal diagonal matrix. [^14] We follow the convention of overloading the $\operatorname{diag}(\cdot)$ operator, such that $\operatorname{diag}(\bm{v})$ returns a square diagonal matrix with the elements of vector $\bm{v}$ on the main diagonal, and $\operatorname{diag}(\bm{A})$ returns a column vector of the main diagonal elements of a square matrix $\bm{A}$. In a similar vein, we overload notation for functions such as ${\mathcal{R}}$, using the same symbol to denote functions defined on different domains, such as ${\mathcal{H}}$ or ${\mathcal{D}}$, depending on the argument.
Theorem 3.3.
For every feature map $h_{0}\in{\mathcal{H}}_{0}$ there exists a unique pair of matrices $\bm{U}=(\bm{U}^{\rm sy},\bm{U}^{\rm id})\in{\mathcal{D}}$ such that the moment kernel function (4) can be represented as $q_{h_{0}}(z,z^{\prime})=q_{\bm{U}}(z,z^{\prime})$ where
$q_{\bm{U}}(z,z^{\prime})\coloneqq\begin{bmatrix}1_{z=\Delta}&\bm{\phi}^{\rm sy}(z)\end{bmatrix}\bm{U}^{\rm sy}\begin{bmatrix}1_{z^{\prime}=\Delta}&\bm{\phi}^{\rm sy}(z^{\prime})\end{bmatrix}^{\top}+\bm{\phi}^{\rm id}(z)\bm{U}^{\rm id}\bm{\phi}^{\rm id}(z^{\prime})^{\top}1_{z=z^{\prime}}.$ (15)
The regularized loss function (9) it turn can be represented as ${\mathcal{R}}(h_{0},\xi_{t})={\mathcal{R}}(\bm{U},\xi_{t})$ where
${\mathcal{R}}(\bm{U},\xi_{t})\coloneqq{\mathcal{L}}(\bm{U},\xi_{t})+\lambda^{\rm sy}\operatorname{tr}(\bm{G}^{\rm sy}\bm{U}^{\rm sy})+\lambda^{\rm id}\operatorname{tr}(\bm{G}^{\rm id}\bm{U}^{\rm id}),$ (16)
with weighted squared loss
${\mathcal{L}}(\bm{U},\xi_{t})\coloneqq w(N_{t})\bigg{\|}\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}-\bm{\Psi}^{\rm sy}(\bm{z}_{t})\bm{U}^{\rm sy}\bm{\Psi}^{\rm sy}(\bm{z}_{t})^{\top}-\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})\bigg{\|}_{F}^{2},$
for the matrix-valued mappings
$\bm{\Psi}^{\rm sy}(\bm{z}_{t})\coloneqq\begin{bmatrix}1&\bm{0}^{\top}\\ \bm{0}&\bm{\phi}^{\rm sy}(\bm{z}_{t})\end{bmatrix}\in{\mathbb{R}}^{(N_{t}+1)\times(m^{\rm sy}+1)},\quad\bm{\Psi}^{\rm id}(\bm{z}_{t})\coloneqq\begin{bmatrix}\bm{0}^{\top}\\ \bm{\phi}^{\rm id}(\bm{z}_{t})\end{bmatrix}\in{\mathbb{R}}^{(N_{t}+1)\times m^{\rm id}},$
and Gram matrices
$\bm{G}^{\rm sy}\coloneqq\begin{bmatrix}0&\bm{0}^{\top}\\ \bm{0}&\langle{\bm{\phi}^{\rm sy}}^{\top},\bm{\phi}^{\rm sy}\rangle_{{\mathcal{G}}^{\rm sy}}\end{bmatrix}\in{\mathbb{S}}_{+}^{m^{\rm sy}+1},\quad\bm{G}^{\rm id}\coloneqq\langle{\bm{\phi}^{\rm id}}^{\top},\bm{\phi}^{\rm id}\rangle_{{\mathcal{G}}^{\rm id}}\in{\mathbb{S}}_{+}^{m^{\rm id}}.$ (17)
Hence ${\mathcal{R}}(\bm{U},\xi_{t})$ is linear-quadratic, and problem (10) becomes convex in $\bm{U}\in{\mathcal{D}}$,
$\underset{\bm{U}\in{\mathcal{D}}}{\text{minimize }}\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{R}}(\bm{U},\xi_{t}).$ (18)
The representation of the moment kernel (15) induces the conditional mean and covariance (COCO) functions (5) expressed in terms of $\bm{U}=(\bm{U}^{\rm sy},\bm{U}^{\rm id})\in{\mathcal{D}}$ as
$\mu_{\bm{U}}(z)$ $=\bm{\phi}^{\rm sy}(z){\bm{b}},$ (19)
$c_{\bm{U}}(z,z^{\prime})$ $=\bm{\phi}^{\rm sy}(z)\big{(}\bm{V}-{\bm{b}}{\bm{b}}^{\top}\big{)}\bm{\phi}^{\rm sy}(z^{\prime})^{\top}+\bm{\phi}^{\rm id}(z)\bm{U}^{\rm id}\bm{\phi}^{\rm id}(z^{\prime})^{\top}1_{z=z^{\prime}},\quad\text{for $\begin{bmatrix}1&{\bm{b}}^{\top}\\ {\bm{b}}&\bm{V}\end{bmatrix}\coloneqq{\bm{U}^{\rm sy}}$.}$
Given a cross section $\xi_{t}=(N_{t},\bm{x}_{t+1},\bm{z}_{t})$ we obtain the corresponding COCO estimates
$\bm{\mu}_{t}$ $=\bm{\phi}^{\rm sy}(\bm{z}_{t}){\bm{b}},$ (20)
$\bm{\Sigma}_{t}$ $=\underbrace{\bm{\phi}^{\rm sy}(\bm{z}_{t})\big{(}\bm{V}-{\bm{b}}{\bm{b}}^{\top}\big{)}\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}}_{\eqqcolon\bm{\Sigma}^{\rm sy}_{t}}+\underbrace{\operatorname{Diag}(\bm{\phi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\phi}^{\rm id}(\bm{z}_{t})^{\top})}_{\eqqcolon\bm{\Sigma}^{\rm id}_{t}}.$
with systematic and idiosyncratic components $\bm{\Sigma}_{t}^{\rm sy}$ and $\bm{\Sigma}_{t}^{\rm id}$.
It follows from (20) that $\bm{\mu}_{t}\in\operatorname{Im}(\bm{\Sigma}_{t})$ if $\bm{\Sigma}_{t}^{\rm id}$ (and hence $\bm{\Sigma}_{t}$) is invertible, an assumption we adopt henceforth. [^15] It is always satisfied in the empirical study below. This implies that the conditional mean-variance efficient (cMVE) portfolio, with weights $\bm{w}_{t}=\bm{\Sigma}_{t}^{+}\bm{\mu}_{t}$, is well-defined and attains the maximum Sharpe ratio, which is given by $\sqrt{\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{\mu}_{t}}$.161616(Filipović and Schneider, 2024, Proposition 6.3) demonstrate that the cMVE portfolio can be replicated by trading exclusively in the $m^{\rm sy}$ factor portfolios $\bm{f}_{t+1}$ defined in Theorem 2.1(iii).
From the COCO estimates (20), we can also deduce the linear factor representation
$\bm{x}_{t+1}=\bm{\phi}^{\rm sy}(\bm{z}_{t})\bm{g}_{t+1}+(\bm{\Sigma}_{t}^{\rm id})^{1/2}w_{t+1}(\bm{z}_{t}),$ (21)
which holds in terms of conditional first and second moments. Here, $\bm{g}_{t+1}$ represents an $m^{\rm sy}$-dimensional systematic risk factor process with constant conditional mean ${\mathbb{E}}_{t}[\bm{g}_{t+1}]=\bm{b}$ and covariance matrix $\operatorname{Cov}_{t}[\bm{g}_{t+1}]=\bm{V}-\bm{b}\bm{b}^{\top}$, and $w_{t+1}(\bm{z}_{t})$ is a conditionally uncorrelated white noise process, as specified after (6). This result aligns with and constitutes a special case of Theorem 2.1(ii).
In the empirical study below we specify the idiosyncratic component as follows.
Example 3.4.
Arguably, the simplest idiosyncratic specification is in dimension $m^{\rm id}=1$, with constant feature map $\bm{\phi}^{\rm id}(\cdot)=\phi^{\rm id}_{1}(\cdot)\coloneqq 1$, and $\bm{U}^{\rm id}=u^{\rm id}\in[0,\infty)$. The idiosyncratic component of the covariance function in (19) becomes $u^{\rm id}1_{z=z^{\prime}}$, and the estimate in (20) reads $\bm{\Sigma}_{t}^{\rm id}=u^{\rm id}\bm{I}_{N_{t}}$.
4 Properties of the COCO estimator
In this section, we establish the uniqueness, consistency, and finite-sample guarantees of the COCO estimator. To facilitate the analysis and subsequent implementation, we first express the regularized loss function in vectorized form. All theoretical results are then stated in terms of these vectorized parameters.
4.1 Vectorization of the loss function
We use the (half-)vectorization of (symmetric) matrices $\bm{A}\in{\mathbb{R}}^{n\times n}$ defined as
$\operatorname{vec}(\bm{A})$ $\coloneqq[A_{11},A_{21},\dots,A_{n1},A_{21},\dots,A_{nn}]^{\top}\in{\mathbb{R}}^{n^{2}},$
$\operatorname{vech}(\bm{A})$ $\coloneqq[A_{11},A_{21},\dots,A_{n1},A_{22},A_{23},\dots,A_{nn}]^{\top}\in{\mathbb{R}}^{n(n+1)/2},$
as well as the duplication matrix $\bm{D}_{n}\in{\mathbb{R}}^{n^{2}\times n(n+1)/2}$, defined such that $\operatorname{vec}(\bm{A})=\bm{D}_{n}\operatorname{vech}(\bm{A})$ for all $\bm{A}\in{\mathbb{S}}^{n}$. The composition of $\operatorname{vec}$ and $\operatorname{diag}$ can be expressed as $\operatorname{vec}\operatorname{diag}(\bm{v})=\bm{R}_{n}\bm{v}$ for the $n^{2}\times n$-matrix $\bm{R}_{n}$ whose $i$th column is the standard basis vector $\bm{e}_{(i-1)n+i}$ in ${\mathbb{R}}^{n^{2}}$. In turn, $\bm{R}_{n}^{\top}(\bm{A}\otimes\bm{A})$ is the $n\times m^{2}$-matrix whose $i$th row is $\bm{A}_{i,\cdot}\otimes\bm{A}_{i,\cdot}$, for a $n\times m$-matrix $\bm{A}$. Note that $\bm{R}_{n}^{\top}\bm{R}_{n}=\bm{I}_{n}$ and $\bm{R}_{n}\bm{R}_{n}^{\top}$ is the orthogonal projection in ${\mathbb{R}}^{n^{2}}$ on the $n$-dimensional subspace spanned by $\bm{e}_{(i-1)n+i}$, $i=1,\dots,n$.
The data points $\xi_{t}=(N_{t},\bm{x}_{t+1},\bm{z}_{t})$ take values in the set $\Xi\coloneqq\bigcup_{n\in{\mathbb{N}}}\{\{n\}\times{\mathbb{R}}^{n}\times{\mathcal{Z}}^{n}\}$, which represents the union over all possible cross-sectional sizes. We write $\xi=(N,\bm{x},\bm{z})$ for a generic point in $\Xi$ and denote the vectorized return product matrix as
$\bm{y}(\bm{x})\coloneqq\operatorname{vec}\bigg{(}\begin{bmatrix}1&\bm{x}^{\top}\\ \bm{x}&\bm{x}\bm{x}^{\top}\end{bmatrix}\bigg{)}\in{\mathbb{R}}^{(N+1)^{2}}.$
We define the vectorized Gram matrices (17)
$\bm{g}^{\rm sy}\coloneqq\operatorname{vec}(\bm{G}^{\rm sy})\in{\mathbb{R}}^{(m^{\rm sy}+1)^{2}},\quad\bm{g}^{\rm id}\coloneqq\operatorname{vec}(\bm{G}^{\rm id})\in{\mathbb{R}}^{(m^{\rm id})^{2}},$
and vectorized parameters $\bm{u}^{\rm sy}\coloneqq\operatorname{vech}(\bm{U}^{\rm sy})$, $\bm{u}^{\rm id}\coloneqq\operatorname{vech}(\bm{U}^{\rm id})$, taking values in the vectorized feasible set
${\mathcal{U}}\coloneqq\operatorname{vech}({\mathcal{D}})=\operatorname{vech}({\mathcal{D}}^{\rm sy})\times\operatorname{vech}({\mathbb{S}}_{+}^{m^{\rm id}})\subset{\mathbb{R}}^{M},$
for the total parameter dimension $M\coloneqq(m^{\rm sy}+1)(m^{\rm sy}+2)/2+m^{\rm id}(m^{\rm id}+1)/2$.
Using the above notation, we can then express the regularized loss function in (16) as a quadratic polynomial in the vectorized parameters as stated in the following lemma.
Lemma 4.1.
The regularized loss function (9) can be represented in terms of the vectorized parameter $\bm{u}=\begin{bmatrix}\bm{u}^{\rm sy}\\ \bm{u}^{\rm id}\end{bmatrix}\in{\mathbb{R}}^{M}$ as ${\mathcal{R}}(\bm{U},\xi)={\mathcal{R}}(\bm{u},\xi)$ where
${\mathcal{R}}(\bm{u},\xi)\coloneqq\frac{1}{2}\bm{u}^{\top}\bm{A}(\xi)\bm{u}+\bm{b}(\xi)^{\top}\bm{u}+c(\xi),$ (22)
for the coefficients
$\bm{A}(\xi)$ $\coloneqq\nabla_{\bm{u}}^{2}{\mathcal{R}}(\bm{u},\xi)=2w(N)\bm{Q}(\xi)^{\top}\bm{Q}(\xi),$
$\bm{b}(\xi)$ $\coloneqq\nabla_{\bm{u}}{\mathcal{R}}(\bm{0},\xi)=-2w(N)\bm{Q}(\xi)^{\top}\bm{y}(\bm{x})+\begin{bmatrix}\lambda^{\rm sy}\bm{D}_{m^{\rm sy}+1}^{\top}{\bm{g}^{\rm sy}}\\ \lambda^{\rm id}\bm{D}_{m^{\rm id}}^{\top}{\bm{g}^{\rm id}}\end{bmatrix},$
$c(\xi)$ $\coloneqq{\mathcal{R}}(\bm{0},\xi)=w(N)\|\bm{y}(\bm{x})\|_{2}^{2},$
and where we define the matrix-valued mappings
$\bm{P}(\xi)$ $\coloneqq\begin{bmatrix}\bm{\Psi}^{\rm sy}(\bm{z})\otimes\bm{\Psi}^{\rm sy}(\bm{z})&\bm{R}_{N+1}\bm{R}_{N+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z})\otimes\bm{\Psi}^{\rm id}(\bm{z}))\end{bmatrix}\in{\mathbb{R}}^{(N+1)^{2}\times((m^{\rm sy}+1)^{2}+(m^{\rm id})^{2})},$
$\bm{Q}(\xi)$ $\coloneqq\bm{P}(\xi)\begin{bmatrix}\bm{D}_{m^{\rm sy}+1}&\bm{0}\\ \bm{0}&\bm{D}_{m^{\rm id}}\end{bmatrix}\in{\mathbb{R}}^{(N+1)^{2}\times M}.$
Strict and strong convexity of ${\mathcal{R}}(\bm{u},\xi)$ in $\bm{u}\in{\mathbb{R}}^{M}$ are discussed in detail in Appendix A.
4.2 Consistency and finite-sample guarantees
We assume that the data points $\xi_{t}$, $t=0,\dots,T-1$, are i.i.d. drawn from a distribution ${\mathbb{P}}$ with support in $\Xi$. We define the sample averages $\bm{A}_{T}\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\bm{A}(\xi_{t})$, $\bm{b}_{T}\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\bm{b}(\xi_{t})$, and $c_{T}\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}c(\xi_{t})$, so that the sample average (empirical) regularized loss in (18) is given by
${\mathcal{R}}_{T}(\bm{u})\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}{\mathcal{R}}(\bm{u},\xi_{t})=\frac{1}{2}\bm{u}^{\top}\bm{A}_{T}\bm{u}+\bm{b}_{T}^{\top}\bm{u}+c_{T}.$
We next provide conditions under which the population loss is well defined and the law of large numbers applies.
Lemma 4.2.
Assume that the following moments are finite,
${\mathbb{E}}[w(N)\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{4}]<\infty,\quad{\mathbb{E}}[w(N)\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}^{4}]<\infty,\quad{\mathbb{E}}[w(N)\|\bm{x}\|_{2}^{4}]<\infty.$ (23)
Then $\|\bm{A}(\xi)\|_{F}$, $\|\bm{b}(\xi)\|_{2}$, and $|c(\xi)|$ have finite expectation, and thus we can define the population loss, along with its gradient and Hessian,
${\mathcal{E}}(\bm{u})$ $\coloneqq{\mathbb{E}}[{\mathcal{R}}(\bm{u},\xi)]=\frac{1}{2}\bm{u}^{\top}{\bm{A}}\bm{u}+{\bm{b}}^{\top}\bm{u}+c,$
$\nabla_{\bm{u}}{\mathcal{E}}(\bm{u})$ $={\mathbb{E}}[\nabla_{\bm{u}}{\mathcal{R}}(\bm{u},\xi)]={\bm{A}}\bm{u}+{\bm{b}},$
$\nabla_{\bm{u}}^{2}{\mathcal{E}}(\bm{u})$ $={\mathbb{E}}[\nabla_{\bm{u}}^{2}{\mathcal{R}}(\bm{u},\xi)]={\bm{A}},$
for ${\bm{A}}\coloneqq{\mathbb{E}}[\bm{A}(\xi)]$, ${\bm{b}}\coloneqq{\mathbb{E}}[\bm{b}(\xi)]$, and $c\coloneqq{\mathbb{E}}[c(\xi)]$. Moreover, the law of large numbers applies and ${\mathcal{R}}_{T}(\cdot)\to{\mathcal{E}}(\cdot)$, $\nabla_{\bm{u}}{\mathcal{R}}_{T}(\cdot)\to\nabla_{\bm{u}}{\mathcal{E}}(\cdot)$, and $\nabla_{\bm{u}}^{2}{\mathcal{R}}_{T}(\cdot)\to\nabla_{\bm{u}}^{2}{\mathcal{E}}(\cdot)$ as $T\to\infty$ uniformly in $\bm{u}$ on compacts in ${\mathbb{R}}^{M}$ with probability 1.
The main result of this section is stated below. Unlike standard results in statistical learning, it applies to an estimator constrained by a convex parameter space.
Theorem 4.3.
  1. (i)
    Consistency: Assume that (23) holds and that $\bm{A}$ is non-singular, so that ${\mathcal{E}}$ is strictly convex and there exists a unique minimizer $\bm{u}^{\ast}\coloneqq\operatorname*{\arg\!\min}_{\bm{u}\in{\mathcal{U}}}{\mathcal{E}}(\bm{u})$. [^17] Given Jensen’s inequality, $\bm{u}^{\top}\bm{A}\bm{u}\geq\bm{u}^{\top}{\mathbb{E}}[(N+1)^{-1}\bm{Q}(\xi)]^{\top}{\mathbb{E}}[(N+1)^{-1}\bm{Q}(\xi)]\bm{u}$, so that non-singularity of $\bm{A}$ can be asserted by similar arguments as above Lemma A.1. Then any sequence of minimizers $\bm{u}^{\ast}_{T}\in\operatorname*{\arg\!\min}_{\bm{u}\in{\mathcal{U}}}{\mathcal{R}}_{T}(\bm{u})$ converges, $\bm{u}^{\ast}_{T}\to\bm{u}^{\ast}$ as $T\to\infty$, with probability 1.
  2. (ii)
    Mean squared error bound: Assume further that ${\mathcal{R}}(\bm{u},\xi)$ is $\alpha$-strongly convex in $\bm{u}$ for ${\mathbb{P}}$-a.e. $\xi\in\Xi$, for some $\alpha>0$, see Lemma A.1, and
    ${\mathbb{E}}[\|(\bm{A}(\xi)-\bm{A})\bm{u}^{\ast}+\bm{b}(\xi)-\bm{b}\|_{2}^{2}]\leq\sigma^{2},$ (24)
    for some $\sigma>0$. Then ${\mathcal{E}}$ and ${\mathcal{R}}_{T}$ are $\alpha$-strongly convex, so that the minimizers $\bm{u}^{\ast}_{T}=\operatorname*{\arg\!\min}_{\bm{u}\in{\mathcal{U}}}{\mathcal{R}}_{T}(\bm{u})$ are unique, and
    ${\mathbb{E}}[\|\bm{u}^{\ast}_{T}-\bm{u}^{\ast}\|_{2}^{2}]\leq\frac{\sigma^{2}}{\alpha^{2}T}.$
  3. (iii)
    Finite-sample guarantees: Assume further that
    ${\mathbb{E}}[\exp(\tau^{-2}\|(\bm{A}(\xi)-\bm{A})\bm{u}^{\ast}+\bm{b}(\xi)-\bm{b}\|_{2}^{2})]\leq\exp(1),$ (25)
    for some $\tau>0$. Then for all $\epsilon>0$, ${\mathbb{P}}[\|\bm{u}^{\ast}_{T}-\bm{u}^{\ast}\|_{2}\geq\epsilon]\leq 2\exp(-\tau^{-2}T\epsilon^{2}\alpha^{2}/3)$. This can equivalently be expressed as: for any $\delta\in(0,1)$, with sample probability of at least $1-\delta$, it holds that
    $\|\bm{u}^{\ast}_{T}-\bm{u}^{\ast}\|_{2}\leq\frac{\sqrt{\log(2/\delta)}\sqrt{3}\tau}{\alpha\sqrt{T}}.$
  4. (iv)
    Consistency (i), mean squared error bound (ii), and finite-sample guarantees (iii) extend to the implied moment kernel functions (15), and thus the COCO estimates (20), using the fact that
    $\big{|}q_{\bm{u}_{T}^{\ast}}(z,z^{\prime})-q_{\bm{u}^{\ast}}(z,z^{\prime})\big{|}\leq C(z,z^{\prime})\|\bm{u}_{T}^{\ast}-\bm{u}^{\ast}\|_{2},\quad\text{for $z,z^{\prime}\in{\mathcal{Z}}_{\Delta}$,}$ (26)
    where $C(z,z^{\prime})\coloneqq\sqrt{2}\big{(}(1_{z=\Delta}+\|\bm{\phi}^{\rm sy}(z)\|_{2})(1_{z^{\prime}=\Delta}+\|\bm{\phi}^{\rm sy}(z^{\prime})\|_{2})+\|\bm{\phi}^{\rm id}(z)\|_{2}^{2}1_{z=z^{\prime}}\big{)}$.
  5. (v)
    Condition (25) implies (24) for $\sigma^{2}=\tau^{2}$. A sufficient condition for (25) to hold is that $\bm{\phi}^{\rm sy}$ and $\bm{\phi}^{\rm id}$ are uniformly bounded functions on ${\mathcal{Z}}$, the individual returns $x_{t+1,i}$ are uniformly bounded, and $N_{t}^{2}w(N_{t})$ is uniformly bounded, ${\mathbb{P}}$-almost surely.
  6. (vi)
    All statements of this theorem hold verbatim if ${\mathcal{U}}$ is replaced by any closed convex subset of ${\mathcal{U}}$.
As an example of a closed convex subset of ${\mathcal{U}}$ mentioned in Theorem 4.3(vi), consider the block parametrization
$\bm{U}^{{\rm sy}}_{\operatorname{diag}}\coloneq\begin{bmatrix}1&\bm{b}^{\top}\\ \bm{b}&\operatorname{diag}\bm{c}\end{bmatrix},$
for a ${\mathbb{R}}^{m^{\rm sy}}$-vector $\bm{c}$. This parametrization allows replacing the semidefinite constraint $\bm{U}\in{\mathcal{D}}$, which may restrict $m^{{\rm sy}}\leq 100$, due to the quadratic growth of the number of parameters. We show below that the number of quadratic constraints associated with the diagonal specification grows only linearly, which would allow to solve large problems, with essentially unrestricted $m^{{\rm sy}}$. The following result substantiates this claim and provides a constructive description of the quadratic constraints.
Lemma 4.4.
Parameter matrix $\bm{U}^{{\rm sy}}_{\operatorname{diag}}\in{\mathbb{S}}_{+}^{m^{{\rm sy}}+1}$ if and only if $c_{1},\ldots,c_{m^{{\rm sy}}}\geq 0$ and there are parameters $\tilde{c}_{1},\ldots,\tilde{c}_{m^{{\rm sy}}}\geq 0$ such that $\sum_{i=1}^{m^{{\rm sy}}}\tilde{c}_{i}\leq 1$ and $b_{i}^{2}\leq c_{i}\tilde{c}_{i}$. The set ${\mathcal{D}}^{{\rm sy}}_{\operatorname{diag}}\coloneq\{\bm{U}^{{\rm sy}}_{\operatorname{diag}}\in{\mathbb{S}}_{+}^{m^{{\rm sy}}+1}\}\subset{\mathcal{D}}^{{\rm sy}}$ is convex and closed.
The next section presents a large-scale implementation of the COCO estimator.
5 Empirical study
This section empirically evaluates the COCO estimator. We first describe the data comprising US stock returns (1962–2021), outline the model specifications, and then assess both statistical performance and asset pricing implications.
5.1 Data and model specification
We use unbalanced monthly stock data compiled by Gu et al. (2020a), covering March 1957 to December 2021. This dataset includes approximately 30,000 stocks, with an average of 6,200 stocks per month. It also contains Treasury bill data for calculating monthly excess returns. The dataset comprises 94 stock-level characteristics (61 updated annually, 13 quarterly, and 20 monthly), 74 industry dummies based on the first two digits of Standard Industrial Classification (SIC) codes, and eight macroeconomic predictors from Welch and Goyal (2008). We restrict the sample to data from 1962 onward, include only common stocks of corporations (sharecodes 10 and 11), and discard months where less than 30% of the covariates are observed.
Figure 1 displays the number of stocks per month (in blue) alongside the running average (in red). Early in the sample period, several months have fewer than thirty stock observations. The cross-sectional sample size peaks in the years leading up to 2000, with the running average stabilizing around 4,000 stocks per month toward the end of the sample.
Figure 1: Size of cross section. The blue line shows the number of assets $N_{t}$ over time. The red line shows a running average. The sample consists of stock data compiled by Gu et al. (2020a), covering the period from 1962 to 2021.
We specify the systematic RKHS ${\mathcal{G}}^{\rm sy}$ using the cosine kernel (see Schölkopf and Smola, 2018) and the Gaussian kernel (see Rasmussen and Williams, 2005) given by, [^18] We also implemented Laplace and inverse multi-quadric kernels $k^{lap}(z,z^{\prime})\coloneqq e^{-\|z-z^{\prime}\|_{2}/{\rho_{lap}}}$, $k^{imq}(z,z^{\prime})\coloneqq 1/\sqrt{\|z-z^{\prime}\|_{2}^{2}+\rho_{imq}}$, which perform similarly. Results are available upon request.
$k^{cos}(z,z^{\prime})\coloneqq\frac{\langle z,z^{\prime}\rangle_{2}}{\|z\|_{2}\,\|z^{\prime}\|_{2}},\quad k^{gauss}(z,z^{\prime})\coloneqq e^{-\frac{\|z-z^{\prime}\|^{2}_{2}}{2\,\rho_{gauss}}}.$
The cosine (or correlation) kernel is a finite-dimensional quasi-linear kernel with no hyperparameters, while the Gaussian kernel is non-linear, generating an infinite-dimensional space of smooth, rapidly decaying functions and includes a length-scale hyperparameter, $\rho_{gauss}>0$. We specify the idiosyncratic RKHS ${\mathcal{G}}^{\rm id}$ using the simplest configuration, with dimension $m^{\rm id}=1$, as outlined in Example 3.4.
[…]
To this end, we use the statistical scoring rule ${\mathcal{S}}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}\to\mathbb{R}$ proposed by Dawid and Sebastiani (1999),
${\mathcal{S}}(\bm{x},\bm{\mu},\bm{\Sigma})\coloneqq\log\det\bm{\Sigma}+(\bm{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu}),$ (27)
which we evaluate for each validation month $t+1$ using observed returns $\bm{x}_{t+1}$ and the COCO estimates in (20) at time $t$. To compute $\bm{\Sigma}_{t}^{-1}$, we use the Woodbury formula,
$\bm{\Sigma}_{t}^{-1}=\frac{1}{u^{{\rm id}}}\bigg{[}\bm{I}_{N_{t}}-\bm{\phi}^{\rm sy}(\bm{z}_{t})\bigg{(}(\bm{V}-\bm{b}\bm{b}^{\top})^{-1}+\frac{\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}\bm{\phi}^{\rm sy}(\bm{z}_{t})}{u^{{\rm id}}}\bigg{)}^{-1}\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}\frac{1}{u^{{\rm id}}}\bigg{]},$
exploiting the fact that $\bm{\Sigma}_{t}^{\rm id}$ is diagonal and full-rank for $u^{{\rm id}}>0$. Additionally, we efficiently compute the determinant in (27) using the formula by Sylvester (1851),
$\det\bm{\Sigma}_{t}=(u^{\rm id})^{N_{t}-m}\det\big{(}\bm{I}_{m}u^{\rm id}+\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top}\bm{\phi}^{\rm sy}(\bm{z}_{t})(\bm{V}-\bm{b}\bm{b}^{\top})\big{)}.$
We solve the semidefinite convex problem (18) using ApS (2025), constraining the smallest eigenvalue of $\bm{U}^{{\rm sy}}$ to be greater than $31e-$ 3, and $u^{{\rm id}}\geq 1e-6$, respectively.
For model evaluation, we use an eight-year training window (96 months) and one month for validation, with all out-of-sample tests conducted on the first month following the validation month. Leveraging the high computational efficiency of the procedure in Section 2, we roll the training, validation, and test windows forward each month, iteratively repeating the training, validation, and testing steps.
5.2 Statistical performance
We assess the statistical performance of the COCO model by comparing it to a simple benchmark model, as no established benchmark exists for jointly estimating conditional first and second moments for unbalanced panels. Our benchmark is a purely idiosyncratic model with zero mean and constant covariance, defined through
$\bm{\mu}_{t}^{\text{bm}}\coloneqq\bm{0},\quad\bm{\Sigma}_{t}^{\text{bm}}\coloneqq\sigma^{2}_{\text{bm}}\bm{I}_{N_{t}},$ (28)
with a single parameter, $\sigma^{2}_{\text{bm}}$, to be estimated. This model is nested within (4), with zero systematic component ($h^{\rm sy}=0$) and a one-dimensional idiosyncratic specification ($m^{{\rm id}}=1$). The minimizer of (10) for this idiosyncratic specification (28) has closed-form solution,
$\sigma^{2}_{\text{bm}}=\frac{\sum_{t=0}^{T-1}w(N_{t})\|\bm{x}_{t+1}\|_{2}^{2}}{\sum_{t=0}^{T-1}w(N_{t})N_{t}}.$
We compare out-of-sample realizations to conditional moments, evaluating first and second moments separately before jointly assessing them with the scoring rule (27). For first moments, we use the predictive out-of-sample $R$-squared measure,
$R^{2}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}-\bm{\phi}^{\rm sy}(\bm{z}_{s}){\bm{b}}\|_{2}^{2}}{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}\|_{2}^{2}},$ (29)
where both numerator and denominator are weighted by $w(N_{s})$ as in the first-moment error component of (8). For second moments, we compute an out-of-sample predictive $R$-squared measure as,
$R^{2,2}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}\bm{x}_{s+1}^{\top}-\bm{\phi}^{\rm sy}(\bm{z}_{s})\bm{V}\bm{\phi}^{\rm sy}(\bm{z}_{s})^{\top}-u^{{\rm id}}\bm{I}_{N_{s}}\|_{F}^{2}}{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}\bm{x}_{s+1}^{\top}-\sigma^{2}_{\text{bm}}\bm{I}_{N_{s}}\|_{F}^{2}},$ (30)
aligning with the second-moment error component in (8). Note that the parameters $\bm{b}$, $\bm{V}$, $u^{{\rm id}}$, $\sigma^{2}_{\text{bm}}$, and the feature maps $\bm{\phi}^{\rm sy}$ in (29) and (30) in fact vary with $s$, as they are re-estimated and updated with each rolling training and validation window.
Figure 2 shows out-of-sample $R^{2}$ over time. The top row displays a rolling estimates over $24$ months, while the bottom row shows expanding estimates. Results indicate high persistence, with slightly positive $R^{2}$ on average, as seen in the expanding averages. Higher-$m$ specifications tend to perform slightly worse than lower-$m$ ones. Figure 2 also highlights four major stock market crashes (defined by Adrian et al. (2023) from pre-crash peak to post-crash trough): the 1987 Crash (08/1987–12/1987), the Dot-Com Bubble (03/2000–10/2002), the Global Financial Crisis (10/2007–03/2009), and the COVID-19 Pandemic (02/2020–03/2020). No clear pattern is observed in $R^{2}$ across these crashes, with positive $R^{2}$ during the first two and negative during the last two.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
Figure 3 presents the corresponding out-of-sample $R^{2,2}$ over time, showing strong persistence with higher-$m$ specifications outperforming lower-$m$ ones across both kernels. The idiosyncratic specification performs better leading up to the Global Financial Crisis, after which the systematic specification shows marked improvement. Positive $R^{2,2}$ is observed during the 1987 Crash, the Dot-Com Bubble, and the COVID-19 Pandemic, with an overall positive $R^{2,2}$ across the sample period. Specifications with higher $m$ perform better on average.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
The COCO model jointly estimates first and second moments. To evaluate the joint fit across both moments, we use the scoring rule ${\mathcal{S}}$ defined in (27) from Dawid and Sebastiani (1999), which is designed for this purpose. To assess the added value of the systematic specification over a purely idiosyncratic model, we define the scoring loss differential as
${\mathcal{S}}_{t,T,\text{OOS}}\coloneqq\frac{1}{T-t}\sum_{s=t}^{T-1}\big{(}{\mathcal{S}}(\bm{x}_{s+1},\bm{0},\sigma^{2}_{\text{bm}}\bm{I}_{N_{s}})-{\mathcal{S}}(\bm{x}_{s+1},\bm{\mu}_{s},\bm{\Sigma}_{s})\big{)}.$ (31)
Figure 4 shows that our model incorporating both systematic and idiosyncratic risk consistently outperforms the purely idiosyncratic benchmark (28) over most periods (a higher score differential is better). Differences between higher-$m$ and lower-$m$ specifications are substantial, with the higher $m$ specifications performing better. The COCO estimator outperforms the idiosyncratic model uniformly across all data points, underscoring that there is statistical value in the specification. Expanding-window results further validate this, confirming the uniform preference for the full model across time and kernel specifications. Notably, the scoring rule differential between the full and idiosyncratic models is especially pronounced during the four crisis periods.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss (expanding)
[…]
Following up on the factor representation discussed after (20), where the systematic risk factors $\bm{g}_{t+1}$ are typically latent, we apply Theorem 2.1(iii) to derive the portfolio factor representation $\bm{x}_{t+1}=\bm{\phi}^{\rm sy}(\bm{z}_{t})\bm{f}_{t+1}+\bm{\epsilon}_{t+1}$. The portfolio factors, defined as $\bm{f}_{t+1}=\bm{\phi}^{\rm sy}(\bm{z}_{t})^{+}\bm{x}_{t+1}$, serve as proxies for the systematic risk factors. The conditionally uncorrelated residuals are implicitly defined by $\bm{\epsilon}_{t+1}=\bm{x}_{t+1}-\bm{\phi}^{\rm sy}(\bm{z}_{t})\bm{f}_{t+1}$. We evaluate the explanatory power of these observable factors through the explained cross-sectional variation, measured by the total $R^{2}$,
$R^{2,\bm{f}}_{t,T,\text{OOS}}\coloneqq 1-\frac{1}{T-t}\sum_{s=t}^{T-1}\frac{\|\bm{x}_{s+1}-\bm{\phi}^{\rm sy}(\bm{z}_{s})\bm{f}_{s+1}\|^{2}_{2}}{\|\bm{x}_{s+1}\|_{2}^{2}}.$ (32)
For the same reasons outlined below (30), the feature maps $\bm{\phi}^{\rm sy}$ in (32) also vary with $s$. Figure 5 presents the out-of-sample total $R^{2}$ over time, which is significantly positive, maintaining a running average of up to 15% for the $m=40$ specifications, and 10% for $m=5$. Total $R^{2}$ is monotonically increasing in the number of factors. The explained variation is particularly elevated during market crashes, aligning with the scoring rules and underscoring the significance of the systematic components during periods of market turbulence.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
To contrast total $R^{2}$ with the contributions of systematic and idiosyncratic components to the conditional covariance estimates, we compute the ratio
$\rho_{t}^{\bm{f}}\coloneqq\frac{\operatorname{tr}(\bm{\phi}^{\rm sy}(\bm{z}_{t})\operatorname{Cov}_{t}[\bm{f}_{t+1}]\bm{\phi}^{\rm sy}(\bm{z}_{t})^{\top})}{\operatorname{tr}\bm{\Sigma}_{t}},$ (33)
which quantifies the proportion of factor-related variance relative to total variance in our model. This ratio provides a direct measure of the predicted explained variation attributable to the systematic factors. Figure 6 reveals that the systematic component was most pronounced early in the sample period, initially exceeding 40% and then decreasing to a running average below 25%, which is consistent with Figure 5. Hence the idiosyncratic risk explains, on average, more than 75% of the cross-sectional variance. [^19] Strictly speaking, the contribution of the systematic component to the total cross-sectional conditional variance is given by the ratio $\frac{\operatorname{tr}\bm{\Sigma}_{t}^{\rm sy}}{\operatorname{tr}\bm{\Sigma}_{t}}$, which is smaller than $\rho_{t}^{\bm{f}}$. However, the difference is negligible in this empirical study. Indeed, based on the general result in (Filipović and Schneider, 2024, Lemma 6.2), one can infer the bounds $\frac{\operatorname{tr}\bm{\Sigma}_{t}^{\rm sy}}{\operatorname{tr}\bm{\Sigma}_{t}}\leq\rho_{t}^{\bm{f}}\leq\frac{\operatorname{tr}\bm{\Sigma}_{t}^{\rm sy}}{\operatorname{tr}\bm{\Sigma}_{t}}+\frac{m}{N_{t}}$. Similar to the observed explained variation by the portfolio factors, the systematic component is monotonically increasing in the number of factors, and intensifies during market crashes, underscoring its importance during such periods.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
5.3 Asset pricing implications
Given the strong statistical performance of the COCO estimator, we next assess its effectiveness in an out-of-sample portfolio setting. The literature has documented that mean-variance efficient portfolios often underperform out-of-sample, especially when compared to the naive $1/N_{t}$ portfolio rule (Basak et al., 2009). Such underperformance is frequently attributed to inaccuracies in estimated moments. Here, we revisit the conditional cMVE portfolio problem, utilizing conditional means and covariances from the COCO estimator, as detailed following (20). All Sharpe ratios reported below are annualized to facilitate comparison with existing literature.
Figure 7 displays time series of the predicted maximum Sharpe ratios, calculated in annualized terms as $\sqrt{12}\times\sqrt{\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{\mu}_{t}}$, for the two kernels considered in this study. For our model specification, which accounts for both systematic and idiosyncratic risk, the predicted maximum Sharpe ratio is generally positive, whereas the purely idiosyncratic benchmark model in (28) predicts it to be zero, as the conditional mean is identically zero in that case. Both panels in Figure 7 show a natural ordering, with higher values of $m$ yielding higher Sharpe ratios across all data points. The levels generated by the two kernels are substantially different, with the Gauss kernel achieving peaks of a predicted maximum Sharpe ratio of up to eight for $m=40$. For $m=5$, the smallest predicted maximum Sharpe ratio averages above one for both kernel specifications. Notably, there are no distinct patterns observed during major market crashes. We next analyze how these predicted Sharpe ratios translate into realized Sharpe ratios.
(a) Cosine kernel
(b) Gauss kernel
[…]
In Figure 8, we plot rolling estimates of out-of-sample realized Sharpe ratios from the cMVE portfolios with monthly excess returns $\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{x}_{t+1}$ over time, alongside the Sharpe ratios from $1/N_{t}$ portfolios. Higher values of $m$ tend to generate higher realized Sharpe ratios, with peaks exceeding five for both kernels. Remarkably, these realized Sharpe ratios remain high despite only modest evidence of predictability in the first and second moments, underscoring the quality of the joint moment estimates, as also suggested by the scoring rule results in Figure 4. The $1/N_{t}$ portfolios generally exhibit lower Sharpe ratios, showing little correlation with those implied by the cMVE portfolios. The bottom row of Figure 8 shows Sharpe ratios estimated over expanding windows, consistently positive toward the end of the sample, with some values exceeding two for both kernels. While it remains challenging to pinpoint an optimal number of factors, both kernels perform least well with $m=5$, which nonetheless outperforms the $1/N_{t}$ portfolio. The largest Sharpe ratio declines occur during the Global Financial Crisis.
(a) Cosine kernel
(b) Gauss kernel
(c) Cosine kernel (expanding)
(d) Gauss kernel (expanding)
[…]
m=5 m=10 m=20 m=40
(Intercept) $0.13^{***}$ $0.30^{***}$ $0.47^{***}$ $0.75^{***}$
$(0.02)$ $(0.03)$ $(0.04)$ $(0.05)$
Mkt $3.58^{***}$ $4.00^{***}$ $5.25^{***}$ $6.02^{***}$
$(0.47)$ $(0.69)$ $(0.90)$ $(1.23)$
SMB $1.42^{*}$ $3.49^{***}$ $4.80^{***}$ $3.44$
$(0.68)$ $(1.01)$ $(1.31)$ $(1.80)$
HML $2.61^{**}$ $0.91$ $0.13$ $0.30$
$(0.86)$ $(1.27)$ $(1.65)$ $(2.26)$
RMW $2.89^{**}$ $3.10^{*}$ $6.19^{***}$ $5.34^{*}$
$(0.91)$ $(1.35)$ $(1.76)$ $(2.40)$
CMA $5.91^{***}$ $8.89^{***}$ $11.55^{***}$ $11.16^{**}$
$(1.38)$ $(2.03)$ $(2.64)$ $(3.62)$
Adj. R2 $0.16$ $0.10$ $0.10$ $0.05$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
(a) Cosine kernel
m=5 m=10 m=20 m=40
(Intercept) $0.10^{***}$ $0.29^{***}$ $0.48^{***}$ $0.75^{***}$
$(0.02)$ $(0.03)$ $(0.04)$ $(0.07)$
Mkt $3.00^{***}$ $2.58^{***}$ $2.46^{**}$ $3.28^{*}$
$(0.37)$ $(0.62)$ $(0.90)$ $(1.54)$
SMB $1.72^{**}$ $3.94^{***}$ $3.57^{**}$ $1.44$
$(0.54)$ $(0.91)$ $(1.32)$ $(2.26)$
HML $2.13^{**}$ $0.37$ $-0.67$ $-0.71$
$(0.67)$ $(1.14)$ $(1.65)$ $(2.83)$
RMW $1.14$ $4.86^{***}$ $6.23^{***}$ $8.19^{**}$
$(0.72)$ $(1.21)$ $(1.76)$ $(3.01)$
CMA $5.79^{***}$ $10.67^{***}$ $10.88^{***}$ $8.71$
$(1.08)$ $(1.83)$ $(2.65)$ $(4.53)$
Adj. R2 $0.20$ $0.13$ $0.06$ $0.01$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
(b) Gaussian kernel
Table 1: Fama–French 5 factors and cMVE portfolio. The table shows results from a time-series regression of out-of-sample cMVE portfolio excess returns, $\bm{\mu}_{t}^{\top}\bm{\Sigma}_{t}^{+}\bm{x}_{t+1}$, on the five factors from Fama and French (2015). The analysis is based on unbalanced US common stock excess returns and associated covariates from 1962 to 2021.
In summary, the higher-$m$ specifications that also generate higher out-of-sample Sharpe ratios, are largely unrelated to the Fama–French five factors. The findings in this section underscore the effectiveness of the COCO model in an asset pricing context, highlighting the substantial predictive value of the joint COCO estimates. To complement the empirical findings, Appendix C presents a simulation study that further demonstrates the robustness and reliability of our method.
6 Conclusion
We introduce a nonparametric, kernel-based estimator for jointly modeling conditional means and covariance matrices in large, unbalanced panels. We term it the joint conditional mean and covariance (COCO) estimator. COCO is rigorously developed and supported by both consistency and finite-sample guarantees, ensuring strong performance in theory and practice. By construction, it produces symmetric, positive semidefinite conditional covariance matrices in all states and leverages infinite-dimensional hypothesis spaces to flexibly capture complex, nonlinear dependencies in the data.
Empirically, we apply the COCO estimator to a large panel of US stock returns from 1962 to 2021, conditioning on both macroeconomic and firm-specific covariates to obtain time-varying estimates of expected returns and covariances. The results highlight COCO’s strong statistical performance and practical relevance for asset pricing: it delivers conditional mean–variance efficient portfolios with substantial out-of-sample Sharpe ratios that significantly outperform equal-weighted benchmarks. A simulation study further confirms the robustness and reliability of these findings.
Its computational efficiency and flexibility make COCO well suited for large-scale, reproducible empirical analysis, offering a powerful tool for econometricians working with complex data structures in finance and related fields.
Appendix A Convexity of the regularized loss function
In this appendix, we discuss the convexity properties of the regularized loss function ${\mathcal{R}}(\bm{u},\xi)$ in (22). The Hessian matrix $\bm{A}(\xi)$ on the right hand side of (22) is positive semidefinite, and hence ${\mathcal{R}}(\bm{u},\xi)$ is convex in $\bm{u}$ in ${\mathbb{R}}^{M}$. It is strictly convex if and only if $\bm{A}(\xi)$ is non-singular, which again holds if and only if $\bm{Q}(\xi)$ is injective. As the duplication matrices $\bm{D}_{m^{\rm sy}+1}$ and $\bm{D}_{m^{\rm id}}$ are injective, a sufficient (but not necessary) condition for $\bm{Q}(\xi)$ to be injective is that the $(m^{\rm sy}+1)^{2}+(m^{\rm id})^{2}$ column vectors of $\bm{P}(\xi)$ are jointly linearly independent. Necessary (but not sufficient) for the latter to hold is that $(N+1)^{2}\geq(m^{\rm sy}+1)^{2}+(m^{\rm id})^{2}$ and that both $\bm{\Psi}^{\rm sy}(\bm{z})$ and $\bm{\Psi}^{\rm id}(\bm{z})$ are injective.
We qualify this further in the following lemma. Recall that a function $f(\bm{u})$ is $\alpha$-strongly convex if $f(\bm{u})-(\alpha/2)\|\bm{u}\|_{2}^{2}$ is convex. Denote by $\sigma_{\rm min}(\bm{B})$ the smallest singular value of a matrix $\bm{B}$.
Lemma A.1.
Assume that there exists some $\alpha>0$ such that
$2w(N)\sigma_{\rm min}(\bm{P}(\xi))^{2}\geq\alpha,\quad\text{for ${\mathbb{P}}$-a.e.\ $\xi\in\Xi$.}$ (34)
Then ${\mathcal{R}}(\bm{u},\xi)$ is $\alpha$-strongly convex in $\bm{u}$, for ${\mathbb{P}}$-a.e. $\xi\in\Xi$.
In general, we cannot give a-priori lower bounds on $\sigma_{\rm min}(\bm{P}(\xi))$ in terms of the singular values of $\bm{\Psi}^{\rm sy}(\bm{z})$ and $\bm{\Psi}^{\rm id}(\bm{z})$ alone, as the former depends on the interaction between these two blocks. On the other hand, from the Rayleigh–Ritz Theorem (Horn and Johnson, 1990, Theorem 4.2.2), it follows that
$\sigma_{\rm min}(\bm{P}(\xi))$ $\leq\min\{\sigma_{\rm min}(\bm{\Psi}^{\rm sy}(\bm{z})\otimes\bm{\Psi}^{\rm sy}(\bm{z})),\sigma_{\rm min}(\bm{R}_{N+1}\bm{R}_{N+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z})\otimes\bm{\Psi}^{\rm id}(\bm{z})))\}$
$\leq\min\{\sigma_{\rm min}(\bm{\Psi}^{\rm sy}(\bm{z}))^{2},\sigma_{\rm min}(\bm{\Psi}^{\rm id}(\bm{z}))^{2}\}$
where we used that $(\bm{B}\otimes\bm{B})^{\top}(\bm{B}\otimes\bm{B})=(\bm{B}^{\top}\bm{B})\otimes(\bm{B}^{\top}\bm{B})$ and $\sigma_{\rm min}(\bm{B}\otimes\bm{B})=\sigma_{\rm min}(\bm{B})^{2}$ for any matrix $\bm{B}$, and that $\bm{R}_{N+1}\bm{R}_{N+1}^{\top}$ is an orthogonal projection. Hence in order that (34) holds, it is necessary (but not sufficient) that $\sigma_{\rm min}(\bm{\Psi}^{\rm sy}(\bm{z}))$ and $\sigma_{\rm min}(\bm{\Psi}^{\rm id}(\bm{z}))$ are properly bounded away from zero. [^20] For any matrices $\bm{A}$, $\bm{B}$ with same number of rows, the Rayleigh–Ritz Theorem implies that $\sigma_{\rm min}([\bm{A},\bm{B}])\leq\min\{\sigma_{\rm min}(\bm{A}),\sigma_{\rm min}(\bm{B})\}$. But while the right hand side can be strictly positive, the left hand side may be zero. For example if $\bm{A}=\bm{B}=1$.
Appendix B Proofs
This appendix contains all proofs.
B.1 Proof of Theorem 2.1
As stated below (6), we assume that $g_{t+1}$ and $\beta(z)$ take values in a subspace of ${\mathcal{C}}$ of codimension one. This assumption is consistent with Lemma B.1 and can be made without loss of generality, since we may simply extend ${\mathcal{C}}$ to $\mathbb{R}\oplus{\mathcal{C}}$ if needed. We also note that the representation (6) of $x_{t+1,i}$ is not unique. For instance, we can demean the factors, replacing $g_{t+1}$ by $g_{t+1}-b$ and $\alpha(z)$ by $\alpha(z)+\langle\beta(z),b\rangle_{\mathcal{C}}$. Further, we can incorporate the intercept in the systematic component, replacing $\beta(z)$ by $\alpha(z){u}+\beta(z)$ and $g_{t+1}$ by $g_{t+1}+{u}$, for some unit vector ${u}\in{\mathcal{C}}$ that is orthogonal to $g_{t+1}$ and $\beta(z)$, which by assumption exists. Moreover, we can rotate the factors, replacing $g_{t+1}$ by $Ag_{t+1}$ and $\beta(z)$ by $B\beta(z)$ for any linear operators $A,B$ on ${\mathcal{C}}$ such that $B^{\ast}A=I_{\mathcal{C}}$.
We now proceed with the proof of Theorem 2.1(i): Without loss of generality, we can assume that $b=0$; if not, we simply replace $g_{t+1}$ by $g_{t+1}-b$ and $\alpha(z)$ by $\alpha(z)+\langle\beta(z),b\rangle_{\mathcal{C}}$. We then incorporate $\alpha(z_{t,i})$ into the scalar product in (6) by extending $g_{t+1}$ with an orthogonal unit vector ${u}\in{\mathcal{C}}$, such that $\langle g_{t+1},{u}\rangle_{\mathcal{C}}=0$, $\langle\beta(z),{u}\rangle_{\mathcal{C}}=0$ for all $z\in{\mathcal{Z}}$, and $\|{u}\|_{\mathcal{C}}=1$. Such a vector ${u}$ always exists by assumption. Consequently, we can express $\alpha(z)+\langle\beta(z),g_{t+1}\rangle_{\mathcal{C}}=\langle\alpha(z){u}+\beta(z),{u}+g_{t+1}\rangle_{\mathcal{C}}$. With regard to the extension (3), we extend $\alpha$, $\beta$ and $\gamma$ to ${\mathcal{Z}}_{\Delta}$ by setting them to zero for $z=\Delta$. Additionally, we introduce the auxiliary index $i=0$ by defining $x_{t+1,0}\coloneqq 1$ and $z_{t,0}\coloneqq\Delta$, and include the indicator function $1_{z=\Delta}$. This leads to the consistent extension of (6) given by
$x_{t+1,i}=\langle\alpha(z_{t,i}){u}+\beta(z_{t,i})+{u}1_{z_{t,i}=\Delta},{u}+g_{t+1}\rangle_{\mathcal{C}}+\gamma(z_{t,i})w_{t+1}(z_{t,i}).$ (35)
As a result, the conditional first and second moments are given by
${\mathbb{E}}_{t}[x_{t+1,i}\,x_{t+1,j}]=\Big{\langle}\big{(}Q+{u}\otimes{u}\big{)}\big{(}\alpha(z_{t,j}){u}+\beta(z_{t,j})+{u}1_{z_{t,j}=\Delta}\big{)},\\ \alpha(z_{t,i}){u}+\beta(z_{t,i})+{u}1_{z_{t,i}=\Delta}\Big{\rangle}_{\mathcal{C}}+\gamma(z_{t,i})^{2}1_{z_{t,i}=z_{t,j}},$ (36)
for all $i,j=0,\dots,N_{t}$. A simple check shows that (36) is perfectly captured by (4) or, equivalently, by (5), where we set $p\coloneqq(Q+{u}\otimes{u})^{1/2}{u}$, $h^{\rm sy}(z)\coloneqq(Q+{u}\otimes{u})^{1/2}(\alpha(z){u}+\beta(z))$, and let $h^{\rm id}$ be such that $\|h^{\rm id}(z)\|_{\mathcal{C}}=\gamma(z)$. Note that $p$ is a unit vector, as $Q{u}=0$.
(ii): Conversely, let the moment kernel function $q_{h}(z,z^{\prime})$ be given in terms of a unit vector $p\in{\mathcal{C}}$ and feature maps $h^{\rm sy},h^{\rm id}:{\mathcal{Z}}\to{\mathcal{C}}$ as in (4). Define $\alpha(z)\coloneqq\langle h^{\rm sy}(z),p\rangle_{\mathcal{C}}$, $\beta(z)\coloneqq h^{\rm sy}(z)-\alpha(z)p$, $\gamma(z)\coloneqq\|h^{\rm id}(z)\|_{\mathcal{C}}$, and let $\{\zeta_{t+1,i}:i=1,2,\dots\}$ and $\{w_{t+1}(z):z\in{\mathcal{Z}}\}$ be conditionally uncorrelated white noise processes with conditional mean zero and conditional variance one. Let $e_{0}\coloneqq p,e_{1},e_{2},\dots$ be an orthonormal basis of ${\mathcal{C}}$, and define $g_{t+1}\coloneqq\sum_{i\geq 1}e_{i}\zeta_{t+1,i}$ and ${u}\coloneqq p$. Then $g_{t+1}$ has a constant conditional covariance operator given by $Qp=0$ and $Qe_{i}=e_{i}$ for $i=1,2,\dots$. It can now be easily verified that the right hand side of (36) equals $q_{h}(z_{t,i},z_{t,j})$, as desired.
(iii): This follows from (35) as proved in (Filipović and Schneider, 2024, Lemma 6.2), where also the formal expressions are given for the conditional mean and covariance of $f_{t+1}$ and $\epsilon_{t+1}$. Note that, in contrast to $g_{t+1}$ and the idiosyncratic risk in (6), the factors $f_{t+1}$ are not stationary and the conditional covariance matrix of $\epsilon_{t+1}$ is not diagonal and does not have full rank.
B.2 Proof of Theorem 3.1
Define the linear sample operator $S^{\tau}:{\mathcal{H}}^{\tau}\to{\mathcal{C}}^{N_{\rm tot}}$ by
$S^{\tau}h^{\tau}\coloneqq[h^{\tau}(z_{t,i}):i=1,\dots,N_{t},\,t=0,\dots,T-1].$
We claim that its adjoint ${S^{\tau}}^{\ast}\bm{\gamma}^{\tau}$ is given by the right hand side of (11), for $\bm{\gamma}^{\tau}=[\gamma_{t,i}^{\tau}:i=1,\dots,N_{t},\,t=0,\dots,T-1]$. Indeed, let $f\in{\mathcal{G}}^{\tau}$ and $v\in{\mathcal{C}}$, then
$\langle{S^{\tau}}^{\ast}\bm{\gamma}^{\tau},f\otimes v\rangle_{{\mathcal{H}}^{\tau}}$ $=\langle\bm{\gamma}^{\tau},{S^{\tau}}(f\otimes v)\rangle_{{\mathcal{C}}^{N_{\rm tot}}}=\sum_{t=0}^{T-1}\sum_{i=1}^{N_{t}}\langle k^{\tau}(\cdot,z_{t,i}),f\rangle_{{\mathcal{G}}^{\tau}}\langle\gamma^{\tau}_{t,i},v\rangle_{\mathcal{C}}$
$=\Big{\langle}\sum_{t=0}^{T-1}\sum_{i=1}^{N_{t}}k^{\tau}(\cdot,z_{t,i})\otimes\gamma^{\tau}_{t,i},f\otimes v\Big{\rangle}_{{\mathcal{H}}^{\tau}},$
which proves the claim. We define by ${\mathcal{G}}_{1}^{\tau}$ the subspace in ${\mathcal{G}}^{\tau}$ spanned by $\{k^{\tau}(\cdot,z_{t,i}):i=1,\dots,N_{t},\,t=0,\dots,T-1\}$. It has finite dimension, $\dim({\mathcal{G}}_{1}^{\tau})\leq N_{\rm tot}$, and thus is closed in ${\mathcal{G}}^{\tau}$. Hence $\operatorname{Im}({S^{\tau}}^{\ast})={\mathcal{G}}_{1}^{\tau}\otimes{\mathcal{C}}$ is a closed subspace in ${\mathcal{H}}^{\tau}$. Consequently, ${\mathcal{H}}^{\tau}=\ker(S^{\tau})\oplus\operatorname{Im}({S^{\tau}}^{\ast})$. Now let $h=(h^{\rm sy},h^{\rm id})$ be any minimizer of (10), and decompose $h^{\tau}=h^{\tau}_{0}+h^{\tau}_{1}$ with $h^{\tau}_{0}\in\ker(S^{\tau})$ and $h^{\tau}_{1}\in\operatorname{Im}({S^{\tau}}^{\ast})$. Clearly, the loss function ${\mathcal{L}}(h,\xi_{t})={\mathcal{L}}(Sh,\xi_{t})$ is a function of $Sh=(S^{\rm sy}h^{\rm sy},S^{\rm id}h^{\rm id})=h_{1}$ only. On the other hand, the norm $\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}\geq\|h^{\tau}_{1}\|_{{\mathcal{H}}^{\tau}}$ is greater than or equal for $h^{\tau}$ than for $h^{\tau}_{1}$, with equality if and only if $h^{\tau}_{0}=0$. As the regularization parameters in (9) are assumed to be positive, $\lambda^{\rm sy},\lambda^{\rm id}>0$, this completes the proof.
B.3 Proof of Proposition 3.2
We have ${\mathcal{H}}^{\tau}_{0}\cong{\mathcal{G}}^{\tau}_{0}\otimes{\mathcal{C}}$, where ${\mathcal{G}}^{\tau}_{0}$ denotes the subspace of ${\mathcal{G}}^{\tau}$ spanned by $\bm{\phi}^{\tau}$. In the following, without loss of generality, we assume that the functions $\bm{\phi}^{\tau}$ are orthonormal in ${\mathcal{G}}^{\tau}$, otherwise we simply replace them by $\bm{\phi}^{\tau}\langle{\bm{\phi}^{\tau}}^{\top},\bm{\phi}^{\tau}\rangle_{{\mathcal{G}}^{\tau}}^{-1/2}$. We extend $\bm{\phi}^{\tau}$ to an orthonormal basis $\bm{\psi}^{\tau}=[\psi^{\tau}_{1}\coloneqq\phi^{\tau}_{1},\dots,\psi^{\tau}_{m^{\tau}}\coloneqq\phi^{\tau}_{m^{\tau}},\psi^{\tau}_{m^{\tau}+1},\dots,\psi^{\tau}_{M^{\tau}}]$ of ${\mathcal{G}}^{\tau}_{1}$, the subspace of ${\mathcal{G}}^{\tau}$ spanned by $k^{\tau}(\cdot,\bm{Z})$ with $m^{\tau}\leq{M^{\tau}}\coloneqq\dim({\mathcal{G}}^{\tau})\leq N_{\rm tot}$, as in the proof of Theorem 3.1. Accordingly, we have $k^{\tau}(\bm{Z},\bm{Z}^{\top})=\bm{\psi}^{\tau}(\bm{Z})\bm{\psi}^{\tau}(\bm{Z})^{\top}$, and by the same token $k_{0}(z,z^{\prime})=\bm{\phi}^{\tau}(z)\bm{\phi}^{\tau}(z^{\prime})^{\top}$, see Paulsen and Raghupathi (2016, Theorem 2.10). Any candidate function of the form (11) can thus be written as $h^{\tau}(z)=\bm{\psi}^{\tau}(z)\bm{\gamma}^{\tau}$, and its projection on ${\mathcal{H}}^{\tau}_{0}$ is given by $h_{0}^{\tau}(z)=\begin{bmatrix}\bm{\phi}^{\tau}(z)&\bm{0}^{\top}\end{bmatrix}\bm{\gamma}^{\tau}$, for a coefficient array $\bm{\gamma}^{\tau}\in{\mathcal{C}}^{M^{\tau}}$. Consequently, $q_{h}(z,z^{\prime})-q_{h_{0}}(z,z^{\prime})$ is a kernel function. As $\|\bm{A}\|_{F}\leq\operatorname{tr}(\bm{A})$ for any positive semidefinite matrix $\bm{A}$, the cross-sectional approximation errors of the implied conditional moment matrices $q_{h}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})$ can therefore be bounded by the respective trace errors. Concretely, let $E_{h}$ denote the left hand side of (14) and define $\bm{V}^{\tau}\coloneqq\langle\bm{\gamma}^{\tau},{\bm{\gamma}^{\tau}}^{\top}\rangle_{\mathcal{C}}$. Then
$E_{h}$ $\leq\sum_{t=0}^{T-1}\operatorname{tr}\big{(}q_{h}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})-q_{h_{0}}(\bar{\bm{z}_{t}},\bar{\bm{z}_{t}}^{\top})\big{)}=\operatorname{tr}(q_{h}(\bm{Z},\bm{Z}^{\top}))-\operatorname{tr}(q_{h_{0}}(\bm{Z},\bm{Z}^{\top}))$ (37)
$=\sum_{\tau\in\{{\rm sy},{\rm id}\}}\operatorname{tr}(\bm{\psi}^{\tau}(\bm{Z})\bm{V}^{\tau}\bm{\psi}^{\tau}(\bm{Z})^{\top})-\operatorname{tr}\bigg{(}\begin{bmatrix}\bm{\phi}^{\tau}(\bm{Z})&\bm{0}^{\top}\end{bmatrix}\bm{V}^{\tau}\begin{bmatrix}\bm{\phi}^{\tau}(\bm{Z})^{\top}\\ \bm{0}\end{bmatrix}\bigg{)}$
$\leq\sum_{\tau\in\{{\rm sy},{\rm id}\}}\|\bm{V}^{\tau}\|_{2}\underbrace{\big{(}\operatorname{tr}(\bm{\psi}^{\tau}(\bm{Z})\bm{\psi}^{\tau}(\bm{Z})^{\top})-\operatorname{tr}(\bm{\phi}^{\tau}(\bm{Z})\bm{\phi}^{\tau}(\bm{Z})^{\top})\big{)}}_{=\epsilon^{\tau}_{\rm approx}}$
where we used that $\operatorname{tr}(\bm{B}\bm{A}\bm{B}^{\top})=\operatorname{tr}(\bm{A}\bm{B}^{\top}\bm{B})\leq\|\bm{A}\|_{2}\operatorname{tr}(\bm{B}^{\top}\bm{B})=\|\bm{A}\|_{2}\operatorname{tr}(\bm{B}\bm{B}^{\top})$ for any positive semidefinite matrix $\bm{A}$ and conformal matrix $\bm{B}$. The bound (14) now follows because $\|\bm{V}^{\tau}\|_{2}\leq\operatorname{tr}(\bm{V}^{\tau})=\|h^{\tau}\|_{{\mathcal{H}}^{\tau}}^{2}$, which completes the proof.
Note that the last inequality in (37) is tight, with equality for, e.g., $\bm{V}^{\tau}=\bm{I}_{M^{\tau}}$. This shows, as a side result, that $\epsilon^{\rm sy}_{\rm approx}+\epsilon^{\rm id}_{\rm approx}$ equals the worst case approximation error, when we take the maximum over all coefficients $\bm{\gamma}^{\tau}$ with $\|\langle\bm{\gamma}^{\tau},{\bm{\gamma}^{\tau}}^{\top}\rangle_{\mathcal{C}}\|_{2}\leq 1$.
B.4 Proof of Theorem 3.3
We express any feature maps $h_{0}(\cdot)=(h^{\rm sy}_{0}(\cdot),h^{\rm id}_{0}(\cdot))$ of the form (13) in vector notation as
$h^{\tau}_{0}(\cdot)=\bm{\phi}^{\tau}(\cdot)\bm{\gamma}^{\tau},$ (38)
for the corresponding arrays of coefficients $\bm{\gamma}^{\tau}\coloneqq[\gamma^{\tau}_{1},\dots,\gamma^{\tau}_{m^{\tau}}]^{\top}\in{\mathcal{C}}^{m^{\tau}}$. The regularized loss function (9) in turn can be represented in terms of the coefficients $\bm{\gamma}=(\bm{\gamma}^{\rm sy},\bm{\gamma}^{\rm id})\in{\mathcal{C}}^{m^{\rm sy}}\times{\mathcal{C}}^{m^{\rm id}}$ as ${\mathcal{R}}(h_{0},\xi_{t})={\mathcal{R}}(\bm{\gamma},\xi_{t})$ where
${\mathcal{R}}(\bm{\gamma},\xi_{t})\coloneqq w(N_{t})\bigg{\|}\begin{bmatrix}1&\bm{x}_{t+1}^{\top}\\ \bm{x}_{t+1}&\bm{x}_{t+1}\bm{x}_{t+1}^{\top}\end{bmatrix}-\bm{\Psi}^{\rm sy}(\bm{z}_{t})\bm{U}^{\rm sy}(\bm{\gamma}^{\rm sy})\bm{\Psi}^{\rm sy}(\bm{z}_{t})^{\top}\\ -\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}(\bm{\gamma}^{\rm id})\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})\bigg{\|}_{F}^{2}\\ +\lambda^{\rm sy}\operatorname{tr}(\bm{G}^{\rm sy}\bm{U}^{\rm sy}(\bm{\gamma}^{\rm sy}))+\lambda^{\rm id}\operatorname{tr}(\bm{G}^{\rm id}\bm{U}^{\rm id}(\bm{\gamma}^{\rm id})),$
for the matrix-valued mapping $\bm{U}(\cdot)=({\bm{U}}^{\rm sy}(\cdot),{\bm{U}}^{\rm id}(\cdot)):{\mathcal{C}}^{m^{\rm sy}}\times{\mathcal{C}}^{m^{\rm id}}\to{\mathcal{D}}$ given by
${\bm{U}}^{\rm sy}(\bm{\gamma}^{\rm sy})\coloneqq\begin{bmatrix}1&\langle p,{\bm{\gamma}^{\rm sy}}^{\top}\rangle_{\mathcal{C}}\\ \langle p,{\bm{\gamma}^{\rm sy}}\rangle_{\mathcal{C}}&\langle{\bm{\gamma}^{\rm sy}},{\bm{\gamma}^{\rm sy}}^{\top}\rangle_{\mathcal{C}}\end{bmatrix},\quad{\bm{U}}^{\rm id}(\bm{\gamma}^{\rm id})\coloneqq\langle{\bm{\gamma}^{\rm id}},{\bm{\gamma}^{\rm id}}^{\top}\rangle_{\mathcal{C}},$ (39)
and we used that the norm of $h_{0}^{\tau}$ becomes $\|h_{0}^{\tau}\|_{{\mathcal{H}}^{\tau}}^{2}=\sum_{i,j=1}^{m^{\tau}}\langle\phi^{\tau}_{i},\phi^{\tau}_{j}\rangle_{{\mathcal{G}}^{\tau}}\langle\gamma^{\tau}_{i},\gamma^{\tau}_{j}\rangle_{\mathcal{C}}=\operatorname{tr}(\bm{G}^{\tau}\bm{U}^{\tau}(\bm{\gamma}^{\tau}))$. By Lemma B.1 below, and as we assumed that ${\mathcal{C}}=\ell^{2}$, the mapping $\bm{U}(\cdot)$ is surjective and hence the regularized loss function can directly be reparametrized in terms of $\bm{U}=(\bm{U}^{\rm sy},\bm{U}^{\rm id})\in{\mathcal{D}}$, as stated in (16). The representation of the moment kernel function (15) follows by the same token. This completes the proof of Theorem 3.3.
The following lemma provides the basis for the proof of Theorem 3.3. Notably, it holds for any choice of the auxiliary Hilbert space ${\mathcal{C}}$, which may be finite-dimensional, distinct from $\ell^{2}$ as considered in the main text.
Lemma B.1.
For the mappings ${\bm{U}}^{\rm sy}:{\mathcal{C}}^{m^{\rm sy}}\to{\mathcal{D}}^{\rm sy}$ and ${\bm{U}}^{\rm id}:{\mathcal{C}}^{m^{\rm id}}\to{\mathbb{S}}^{m^{\rm id}}_{+}$ defined in (39) the following hold;
  1. (i)
    ${\bm{U}}^{\rm sy}$ is surjective if and only if $\dim{\mathcal{C}}\geq m^{\rm sy}+1$. If $\dim{\mathcal{C}}\geq 3$ then for any $\bm{\gamma}^{\rm sy}\in{\mathcal{C}}^{m^{\rm sy}}$ there exist infinitely many $\tilde{\bm{\gamma}}^{\rm sy}\neq\bm{\gamma}^{\rm sy}$ in ${\mathcal{C}}^{m^{\rm sy}}$ such that $\bm{U}^{\rm sy}(\tilde{\bm{\gamma}}^{\rm sy})=\bm{U}^{\rm sy}(\bm{\gamma}^{\rm sy})$.
  2. (ii)
    ${\bm{U}}^{\rm id}$ is surjective if and only if $\dim{\mathcal{C}}\geq m^{\rm id}$. If $\dim{\mathcal{C}}\geq 2$ then for any $\bm{\gamma}^{\rm id}\in{\mathcal{C}}^{m^{\rm id}}$ there exist infinitely many $\tilde{\bm{\gamma}}^{\rm id}\neq\bm{\gamma}^{\rm id}$ in ${\mathcal{C}}^{m^{\rm id}}$ such that $\bm{U}^{\rm id}(\tilde{\bm{\gamma}}^{\rm id})=\bm{U}^{\rm id}(\bm{\gamma}^{\rm id})$.
Hence the minimal dimensional requirements of ${\mathcal{C}}$ for Theorem 3.3 to apply are $\dim{\mathcal{C}}\geq\max\{m^{\rm sy}+1,m^{\rm id}\}$.
Proof of Lemma B.1.
(i): Without loss of generality we can assume that $\dim{\mathcal{C}}<\infty$, otherwise we replace ${\mathcal{C}}$ by a finite-dimensional subspace. Define $\nu\coloneq\dim{\mathcal{C}}-1$, and consider an orthonormal basis $\xi_{0}\coloneqq p,\,\xi_{1},\dots,\xi_{\nu}$ of ${\mathcal{C}}$. Then there is a bijection between ${\mathcal{C}}^{m^{\rm sy}}$ and ${\mathbb{R}}^{m^{\rm sy}}\times{\mathbb{R}}^{{m^{\rm sy}}\times\nu}$: every $\bm{\gamma}^{\rm sy}\in{\mathcal{C}}^{m^{\rm sy}}$ can be expressed in unique coordinates as $\gamma^{\rm sy}_{i}=b_{i}p+\sum_{j=1}^{\nu}c_{ij}\xi_{j}$ for some vector ${\bm{b}}=[b_{i}:1\leq i\leq m^{\rm sy}]\in{\mathbb{R}}^{m^{\rm sy}}$ and matrix $\bm{C}=[c_{ij}:1\leq i\leq m^{\rm sy},1\leq j\leq\nu]\in{\mathbb{R}}^{m^{\rm sy}\times\nu}$, and vice versa. Expressed in these coordinates, we can write ${\bm{U}}^{\rm sy}(\bm{\gamma}^{\rm sy})=\begin{bmatrix}1&{\bm{b}}^{\top}\\ {\bm{b}}&{\bm{b}}{\bm{b}}^{\top}+\bm{C}\bm{C}^{\top}\end{bmatrix}$. It follows that $\bm{C}\bm{C}^{\top}$ is the Schur complement of the upper left block 1 of the matrix ${\bm{U}}^{\rm sy}(\bm{\gamma}^{\rm sy})$. Hence ${\bm{U}}^{\rm sy}:{\mathcal{C}}^{m^{\rm sy}}\to{\mathcal{D}}^{\rm sy}$ is surjective if and only if every matrix $\bm{\Sigma}\in{\mathbb{S}}^{m^{\rm sy}}_{+}$ can be expressed as $\bm{\Sigma}=\bm{C}\bm{C}^{\top}$ for some $\bm{C}\in{\mathbb{R}}^{m^{\rm sy}\times\nu}$. This holds if and only if $\nu\geq m^{\rm sy}$, see (Paulsen and Raghupathi, 2016, Theorem 4.7), which proves the first statement. For the second statement, let $\bm{A}\neq\bm{I}_{\nu}$ be any orthogonal $\nu\times\nu$-matrix, and define $\tilde{\bm{C}}=\bm{C}\bm{A}$ and $\tilde{\bm{\gamma}}$ accordingly as above. It follows that $\tilde{\bm{\gamma}}\neq\bm{\gamma}$ and $\bm{U}^{\rm sy}(\tilde{\bm{\gamma}})=\bm{U}^{\rm sy}(\bm{\gamma})$. If $\nu\geq 2$ then there exists infinitely many such matrices $\bm{A}$, which proves the claim.
(ii): This follows similarly as part (i), but without the constraint $\bm{U}^{\rm id}_{11}=1$. ∎
B.5 Proof of Lemma 4.1
Using the introduced notation, we express the regularized loss function ${\mathcal{R}}(\bm{U},\xi_{t})$ in (16) in terms of the vectorized parameter $\bm{u}=\begin{bmatrix}\bm{u}^{\rm sy}\\ \bm{u}^{\rm id}\end{bmatrix}$ as
${\mathcal{R}}(\bm{u},\xi_{t})=w(N_{t})\Big{\|}\bm{y}(\bm{x}_{t+1})-(\bm{\Psi}^{\rm sy}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm sy}(\bm{z}_{t}))\bm{D}_{m^{\rm sy}+1}\bm{u}^{\rm sy}\\ -\bm{R}_{N_{t}+1}\bm{R}_{N_{t}+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm id}(\bm{z}_{t}))\bm{D}_{m^{\rm id}}\bm{u}^{\rm id}\Big{\|}_{2}^{2}\\ +\lambda^{\rm sy}{\bm{g}^{\rm sy}}^{\top}\bm{D}_{m^{\rm sy}+1}\bm{u}^{\rm sy}+\lambda^{\rm id}{\bm{g}^{\rm id}}^{\top}\bm{D}_{m^{\rm id}}\bm{u}^{\rm id},$
where we used that $\operatorname{vec}(\bm{\Psi}^{\rm sy}(\bm{z}_{t})\bm{U}^{\rm sy}\bm{\Psi}^{\rm sy}(\bm{z}_{t})^{\top})=(\bm{\Psi}^{\rm sy}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm sy}(\bm{z}_{t}))\operatorname{vec}(\bm{U}^{\rm sy})$ and
$\operatorname{vec}(\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top}))=\bm{R}_{N_{t}+1}\bm{R}_{N_{t}+1}^{\top}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\otimes\bm{\Psi}^{\rm id}(\bm{z}_{t}))\bm{D}_{m^{\rm id}}\bm{u}^{\rm id},$
given the $i$th diagonal element $(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top})_{ii}=(\bm{\Psi}_{i,\cdot}^{\rm id}(\bm{z}_{t})\otimes\bm{\Psi}_{i,\cdot}^{\rm id}(\bm{z}_{t}))\operatorname{vec}(\bm{U}^{\rm id})$. Expanding the squared norm and collecting terms then gives (22), which proves the lemma.
For the simple idiosyncratic specification in dimension $m^{\rm id}=1$ in Example 3.4, the above expression simplifies to
$\operatorname{vec}(\operatorname{Diag}(\bm{\Psi}^{\rm id}(\bm{z}_{t})\bm{U}^{\rm id}\bm{\Psi}^{\rm id}(\bm{z}_{t})^{\top}))=\underbrace{\bm{R}_{N_{t}+1}}_{(N_{t}+1)^{2}\times(N_{t}+1)}\underbrace{\begin{bmatrix}0\\ \bm{1}\end{bmatrix}}_{(N_{t}+1)\times 1}u^{\rm id},$
and the regularization penalty term reads $\lambda^{\rm id}{\bm{g}^{\rm id}}^{\top}\bm{D}_{m^{\rm id}}\bm{u}^{\rm id}=\lambda^{\rm id}u^{\rm id}$.
B.6 Proof of Lemma A.1
From the Rayleigh–Ritz Theorem (Horn and Johnson, 1990, Theorem 4.2.2), and using that $\|\bm{D}_{n}\bm{v}\|_{2}\geq\|\bm{v}\|_{2}$, it follows that
$\sigma_{\rm min}(\bm{A}(\xi))=2w(N)\sigma_{\rm min}(\bm{Q}(\xi))^{2}\geq 2w(N)\sigma_{\rm min}(\bm{P}(\xi))^{2}.$
This proves the lemma.
B.7 Proof of Lemma 4.2
We use the elementary facts $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|_{F}\|\bm{B}\|_{F}$ and $\|\bm{A}\otimes\bm{B}\|_{F}=\|\bm{A}\|_{F}\|\bm{B}\|_{F}$ for matrices $\bm{A}$ and $\bm{B}$, for the Frobenius norm $\|\cdot\|_{F}$. By construction, it follows that $\|\bm{y}(\bm{x})\|_{2}\leq 1+\|\bm{x}\|_{2}^{2}$, $\|\bm{\Psi}^{\rm sy}(\bm{z})\|_{F}^{2}=1+\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{2}$, $\|\bm{\Psi}^{\rm id}(\bm{z})\|_{F}=\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}$, $\|\bm{D}_{n}\|_{F}=n$, $\|\bm{R}_{n}\bm{B}\|_{F}=\|\bm{B}\|_{F}$, $\|\bm{R}_{n}^{\top}\bm{B}\|_{F}\leq\|\bm{B}\|_{F}$, for any conformal matrix $\bm{B}$. Hence
$\|\bm{A}(\xi)\|_{F}$ $\leq 2w(N)\|\bm{Q}(\xi)\|_{F}^{2}$ (40)
$\leq 2w(N)\big{(}(1+\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{2})^{2}(m^{\rm sy}+1)^{2}+\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}^{4}(m^{\rm id})^{2}\big{)},$
$\|\bm{b}(\xi)\|_{2}$ $\leq 2w(N)\big{(}(1+\|\bm{\phi}^{\rm sy}(\bm{z})\|_{F}^{2})(m^{\rm sy}+1)+\|\bm{\phi}^{\rm id}(\bm{z})\|_{F}^{2}m^{\rm id}\big{)}$
$\quad+\lambda^{\rm sy}(m^{\rm sy}+1)\|\bm{g}^{\rm sy}\|_{2}+\lambda^{\rm id}m^{\rm id}\|\bm{g}^{\rm id}\|_{2},$
$|c(\xi)|$ $\leq w(N)\big{(}1+\|\bm{x}\|_{2}^{2}\big{)}^{2}.$
Combining (23) and (40) proves the lemma.
B.8 Proof of Theorem 4.3
Clearly, ${\mathcal{R}}(\bm{u},\xi)$ is a Carathéodory function, i.e., measurable in $\xi$ and continuous in $\bm{u}$, and therefore random lower semicontinuous (Shapiro et al., 2021, Section 9.2.4). The set ${\mathcal{U}}$ is closed and convex in ${\mathbb{R}}^{M}$. Now claim (i) follows from (Shapiro et al., 2021, Theorem 5.4).
Claims (ii) and (iii) follow from (Milz, 2023, Theorem 3), setting “$\bm{\Psi}(\bm{u})$” in Milz (2023) equal to the convex characteristic function of the feasible set ${\mathcal{U}}$ in ${\mathbb{R}}^{M}$, taking value 00 for $\bm{u}\in{\mathcal{U}}$ and +$\infty$+ ∞ otherwise.
Claim (iv) follows as (15) elementary implies the bound (26), using that the operator norm of the half-vectorization operator is given by $\sup_{\|\bm{u}\|_{2}\leq 1}\|\operatorname{vech}(\bm{u})\|_{F}=\sqrt{2}$.
Claim (v) follows from Jensen’s inequality and the bounds in (40).
Claim (vi) follows as the above proof applies to any closed convex subset of ${\mathcal{U}}$.
B.9 Proof of Lemma 4.4
$\bm{U}_{\operatorname{diag}}^{{\rm sy}}$ is clearly symmetric. Furthermore, all (non-leading) principal minors are diagonal matrices with entries along the diagonal that are combinations of $c_{1},\ldots,c_{m^{{\rm sy}}}\geq 0$ from the premises of the statements, and thus positive semidefinite. The top-left corner is equal to one, and therefore positive. To consider the remaining $l=1,\ldots,m^{{\rm sy}}$ leading principal minors, we apply the block determinant formula to obtain for the determinant of the $l$-th leading principal minor,
$\left(1-\sum_{j=1}^{l}{\frac{b_{j}^{2}}{c_{j}}}\right)\prod_{i=1}^{l}c_{i}\geq(1-\sum_{j=1}^{l}\tilde{c}_{j})\prod_{i=1}^{l}c_{i}\geq 0,$
from the premise of the statement. With all principal minors positive semidefinite, Sylvester’s criterion (Horn and Johnson, 1990, Theorem 7.2.5) applies, and yields that $\bm{U}^{{\rm sy}}_{\operatorname{diag}}$ is symmetric positive semidefinite. Conversely, the matrix $\bm{U}^{{\rm sy}}_{\operatorname{diag}}$ being positive semidefinite implies the leading principal minors to be non-negative, such that in turn $c_{1},\ldots,c_{m^{{\rm sy}}}\geq 0$, $\tilde{c}_{1},\ldots,\tilde{c}_{m^{{\rm sy}}}\geq 0$, $\sum_{i=1}^{m^{{\rm sy}}}\tilde{c}_{i}\leq 1$ and $b_{i}^{2}\leq c_{i}\tilde{c}_{i}$ for $l=1,\ldots,m^{{\rm sy}}$ (Horn and Johnson, 1990, Corollary 7.1.5).
From the block-diagonal specification clearly ${\mathcal{D}}^{{\rm sy}}_{\operatorname{diag}}\subset{\mathcal{D}}^{{\rm sy}}$. Finally, all constraints in the premise of the statement describe closed convex sets (the constraints $b_{i}^{2}\leq c_{i}\tilde{c}_{i},\,i=1,\ldots,m^{{\rm sy}}$ are commonly referred to as rotated quadratic cones, and jointly convex in $b_{i},c_{i}$ and $\tilde{c}_{i}$) and their intersection thus describes a closed convex set.
Appendix C Simulation study
Simulations are essential for assessing the performance of the proposed method under controlled conditions. In this appendix, we investigate the COCO model in a controlled simulation environment. To this end, we employ the simple form (21) of the data-generating model from Theorem 2.1, and specify $\bm{g}_{t+1}$ as an $m^{\rm sy}=40$-dimensional normal random vector with constant conditional mean $\mathbb{E}_{t}[\bm{g}_{t+1}]=\bm{b}^{\text{pop}}$ and conditional covariance matrix $\mathrm{Cov}_{t}[\bm{g}_{t+1}]=\bm{V}^{\text{pop}}-\bm{b}^{\text{pop}}(\bm{b}^{\text{pop}})^{\top}$. The idiosyncratic component $w_{t+1}(\bm{z}_{t})$ is drawn from a normal distribution with mean vector $\bm{0}_{N_{t}}$ and covariance matrix $u^{\text{pop}\,{\rm id}}\bm{I}_{N_{t}}$. We set the population parameters $\bm{b}^{\text{pop}},\bm{V}^{\text{pop}},\bm{\phi}^{{\rm sy}\,\text{pop}}(\cdot)$, and $u^{\text{pop}\,{\rm id}}$ to their full-sample estimates based on the cosine kernel, thereby eliminating the need for validation. The observed covariates are used without modification to generate the simulated return data $\bm{x}_{t+1}^{\text{sim}}$. The resulting simulated dataset thus combines observed covariates with simulated returns.
Next, we replicate the out-of-sample estimation procedure described in Section 5, following exactly the same steps as in the empirical analysis and computing the same out-of-sample statistics. Denoting the (ground truth) population conditional mean and covariance by
$\begin{split}\bm{\mu}_{t}^{\text{pop}}&\coloneqq\bm{\phi}^{{\rm sy}\,\text{pop}}(\bm{z}_{t})\bm{b}^{\text{pop}},\\ \bm{\Sigma}_{t}^{\text{pop}}&\coloneqq\bm{\phi}^{{\rm sy}\,\text{pop}}(\bm{z}_{t})(\bm{V}^{\text{pop}}-\bm{b}^{\text{pop}}(\bm{b}^{\text{pop}})^{\top})\bm{\phi}^{{\rm sy}\,\text{pop}}(\bm{z}_{t})^{\top}+u^{\text{pop}\,{\rm id}}\bm{I}_{N_{t}},\end{split}$ (41)
we define the corresponding out-of-sample evaluation metrics analogously to (29), (30), and (31) as follows
$R^{2,\text{pop}}_{t,T,\text{OOS}}$ $\coloneqq 1-\frac{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}^{\text{sim}}-\bm{\phi}^{{\rm sy}\,\text{pop}}(\bm{z}_{s})\bm{b}^{\text{pop}}\|_{2}^{2}}{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}^{\text{sim}}\|_{2}^{2}},$ (42)
$R^{2,2,\text{pop}}_{t,T,\text{OOS}}$ $\coloneqq 1-\frac{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}^{\text{sim}}\bm{x}_{s+1}^{\text{sim}\top}-\bm{\phi}^{{\rm sy}\,\text{pop}}(\bm{z}_{s})\bm{V}^{\text{pop}}\bm{\phi}^{{\rm sy}\,\text{pop}}(\bm{z}_{s})^{\top}-u^{{\rm id}\,\text{pop}}\bm{I}_{N_{s}}\|_{F}^{2}}{\sum_{s=t}^{T-1}w(N_{s})\|\bm{x}_{s+1}^{\text{sim}}\bm{x}_{s+1}^{\text{sim}\top}-\sigma^{2}_{\text{bm}}\bm{I}_{N_{s}}\|_{F}^{2}},$ (43)
${\mathcal{S}}_{t,T,\text{OOS}}^{\text{pop}}$ $\coloneqq\frac{1}{T-t}\sum_{s=t}^{T-1}\big{(}{\mathcal{S}}(\bm{x}_{s+1}^{\text{sim}},\bm{0},\sigma^{2}_{\text{bm}}\bm{I}_{N_{s}})-{\mathcal{S}}(\bm{x}_{s+1}^{\text{sim}},\bm{\mu}_{s}^{\text{pop}},\bm{\Sigma}_{s}^{\text{pop}})\big{)}).$ (44)
(a) $R^{2}$ and $R^{2,\text{pop}}$
(b) $R^{2,2}$ and $R^{2,2,\text{pop}}$
(c) $R^{2}$ and $R^{2,\text{pop}}$ expanding
(d) $R^{2,2}$ and $R^{2,2,\text{pop}}$ expanding
[…]
Figure 9 shows $R^{2}_{t,T,\text{OOS}}$, $R^{2,2}_{t,T,\text{OOS}}$, $R^{2,\text{pop}}_{t,T,\text{OOS}}$, and $R^{2,2,\text{pop}}_{t,T,\text{OOS}}$ computed from simulated data. The population model accommodates $R^{2}_{t,T,\text{OOS}}$ of around 0.5% and $R^{2,\text{pop}}_{t,T,\text{OOS}}$ of around 7.5%. While the COCO model does not attain either, higher-factor specifications get quite close to $R^{2,2,\text{pop}}_{t,T,\text{OOS}}$, but less so to $R^{2,\text{pop}}_{t,T,\text{OOS}}$. The patterns observed in the simulated data are quite similar to those of the real data.
(a) Scoring loss (rolling)
(b) Scoring loss (expanding)
[…]
Next, we investigate the scoring loss differential of the population, and the COCO estimator with the purely idiosyncratic model in Figure 10. Here, slight differences to the real data become visible in that the performance of the COCO model is best when the number of stocks is the highest, while the real data exhibits some additional patterns at the beginning of the sample (cf. Figure 1). Simulated data yield higher-dimensional models performing better than lower-dimensional ones.
(a) Explained variation
(b) Ratio factor-explained to total variance
(c) Explained variation (expanding)
(d) Ratio factor-explained to total variance (expanding)
[…]
Figure 11 shows the amount of cross-sectional variation explained by the population, and the COCO model. The amount of variation is increasing monotonically with $m^{{\rm sy}}$. Differences in variation explained between higher-dimensional and lower-dimensional models are higher than with real data. While Figure 6 (real data) shows merely a few percentage points difference between $m^{\rm sy}=5$ and $m^{\rm sy}=40$, Figure 11(a) shows a two-fold increase. As far as the ratio of systematic to idiosyncratic risk is concerned, Figure 11(b) shows that the COCO model with $m^{{\rm sy}}=20,40$ gets close to population levels.
(a) Realized Sharpe ratio (rolling)
(b) Realized Sharpe ratio (expanding)
(c) Predicted Sharpe ratio
[…]
Figure 13: Sampling distribution of distance to population parameter. From simulation experiments $i=1,\ldots,2000$, this figure shows the boxplots of the sampling distribution of the log deviation $\log\left(\|\operatorname{vech}\bm{U}^{{\rm sy}}_{T,i}-\operatorname{vech}\bm{U}^{{\rm sy}\,\text{pop}}\|_{2}^{2}+(u^{{\rm id}}_{T,i}-u^{{\rm id}\,\text{pop}})^{2}\right)$ for $T=100,1000,10000$ and $m^{{\rm sy}}=40$.
Figure 12 shows realized Sharpe ratios in the simulated economy, with stark differences between the population and the COCO model. While the COCO-induced Sharpe ratios can be found in the vicinity of the observed, real data, knowledge of the population moments yield very high Sharpe ratios of three in annualized terms. Predicted Sharpe ratios are lower than realized ones for specifications with higher $m$, as shown by Figure 12(c), but agree with $m=5$.
Figure 13 shows box plots of the expected distance between the COCO estimator and the population parameter, based on 2000 simulated datasets of sizes $T=100,1000,10000$, shown on a logarithmic scale. Despite the fact that the COCO estimator arises from a non-standard, constrained optimization problem, the reduction in expected distance aligns closely with the rate predicted by the mean squared error bound in Theorem 4.3(ii), as well as the finite-sample guarantee in Theorem 4.3(iii). The figure reveals a nearly linear decay of the upper bound with the logarithm of the sample size $T$, indicating that the theoretical bounds are relatively tight.